### v1

# Energy-Efficient Channel Decoding for Wireless Federated Learning: Convergence Analysis and Adaptive Design

Abstract
One of the most critical challenges for deploying distributed learning, such as federated learning (FL), in wireless networks is the limited battery capacity of mobile devices. While it is a common belief that the major energy consumption of mobile devices comes from the uplink data transmission, this paper presents a novel finding, namely the channel decoding operation also contributes significantly to the overall energy consumption of mobile devices in FL. Motivated by this new observation, we propose an energy-efficient adaptive channel decoding scheme that leverages the intrinsic robustness of FL to model errors. In particular, the robustness is exploited to reduce the energy consumption of channel decoders at mobile devices by adaptively adjusting the number of decoding iterations. We theoretically prove that FL with communication errors can converge at the same rate as error-free communication as long as the bit error rate (BER) is properly constrained. An adaptive channel decoding scheme is then proposed to improve the energy efficiency of FL systems. Experimental results demonstrate that the proposed method maintains the same learning accuracy while reducing the channel decoding energy consumption by 20% when compared to existing approaches.
Index Terms:
Federated learning (FL), energy efficiency, channel decoding, low-density parity checking (LDPC), bit error rate (BER).
I Introduction
I-A Motivation
Deep learning has achieved remarkable success in numerous realms, such as computer vision and natural language processing [1]. However, the conventional centralized learning paradigm encounters the data isolation issue due to the rising concern about privacy [2]. As a result, uploading the distributed data, which may contain privacy-sensitive information, for centralized learning becomes prohibited. To address this issue, federated learning (FL) was proposed [3] to enable the training of deep learning models without sharing the data of users (e.g., IoT devices). However, to fully unlock the potential of FL, we need to tackle several critical challenges, such as data and device heterogeneity [4, 5, 6], as well as the communication bottleneck[7]. Another issue for wireless FL is the energy consumption of mobile devices due to heavy computation and communication workloads for training and transferring the deep neural network (DNN) parameters[8].
Many innovative approaches have been proposed to reduce the energy consumption of mobile devices in FL from different perspectives. For instance, reinforcement learning was utilized in [9] to identify a near-optimal clustering of participating devices to reduce the overall energy consumption of mobile devices. In [10], the trade-off between learning accuracy and users’ total energy consumption was optimized through computational and communication resource allocation. Besides system-level resource management, some works focused on the power consumption of specific hardware components such as transmitters, receivers, and artificial intelligence (AI) chips. Techniques such as model sparsification [11] and quantization [12, 13, 14, 15, 16, 17] have been proposed to minimize the communication overhead, thus reducing the energy consumption of the transmitters. The authors of [18] and [19] minimized the transmission energy while maintaining the learning performance. From computation perspective, model sparsification and quantization techniques are also beneficial for minimizing the processing energy consumption [20, 21, 22]. Additionally, some works, such as [10], [23], and [24], considered the joint optimization of the uplink transmission and local training energy consumption. However, very few works have paid attention to the energy consumption of channel decoder. As will be shown in Section II-B, channel decoders consume a substantial portion of energy because of the iterative process to decode massive DNN parameters.
To address the research gap, in this work, we first quantitatively analyze the energy consumption of different components in mobile clients and show that the channel decoder contributes significantly to the overall energy consumption. Building on this observation, an energy-efficient channel decoding strategy is designed for wireless FL systems, by leveraging the fact that FL is robust to model errors due to its statistical learning nature[25]. Furthermore, we conduct a theoretical analysis on the relation between the bit error rate (BER) and the convergence rate of FL, and determine the maximum tolerable BER that can maintain the same convergence rate as the case with error-free communication. Finally, we propose a practical approach to adjust the decoding iterations in each communication round to achieve the maximum tolerable BER, while reducing the energy consumption.
I-B Contributions
This paper investigates the impact of channel decoding on the energy consumption of wireless FL. The major contributions are summarized as follows:
  • We theoretically analyze the relation between BER and the convergence rate of FL. The maximum tolerable BER to achieve the same convergence rate as error-free FL is determined.
  • Based on the analysis, an adaptive low-density parity checking (LDPC) decoding mechanism is proposed, which controls the maximum number of decoding iterations to reduce the energy consumption of wireless FL while satisfying the BER requirement for a target convergence rate.
  • Experimental results show that the proposed adaptive scheme effectively reduces the energy consumption by 20% when compared with the traditional scheme, while guaranteeing the same learning performance.
I-C Organization
The remainder of this paper is organized as follows. Section II introduces the system model and presents a comprehensive energy consumption analysis for mobile devices. Section III develops an energy-efficient LDPC decoding strategy for FL systems by exploiting the relationship between the convergence rate of FL and the achieved BER of LDPC decoding. Experimental results are presented in Section IV. Finally, concluding remarks are drawn in Section V.
II System Model and Preliminaries
In this section, we first introduce the wireless FL system. An overview regarding the energy consumption of different components at a mobile unit is then presented, which highlights the significant contribution of channel decoding to the total energy consumption of wireless FL. For ease of reference, we list the key notations of the system model in Table I.
TABLE I: Key Notations
Notation Definition
$\mathbf{w}$ DNN model.
$\Delta\mathbf{w}$ The local model update.
$f$ The loss function of training.
$f^{*}$ The minimum value of the training loss.
${\nabla}f$ The gradient.
$r$ The index of the communication round.
$R$ The total communication rounds.
$e$ The index of the local training step.
$E$ The local training steps.
$\sigma_{L}^{2}$ The variance of local gradients estimator.
$\sigma_{G}^{2}$ The dissimilarity between local and global gradients.
$k$ The index of the mobile device.
$n$ The number of total mobile devices.
$s$ The number of selected clients in every round
$S_{r}$ The set of selected clients in the $r$-th round.
$d$ The dimensions of the DNN model.
$p_{k}$ The data size ratio of the $k$-th client.
$T$ The total training iterations of one client.
$L$ The L-smooth parameter.
$b_{r}$ The bit error rate in $r$-th communication round.
$\eta$ The learning rate.
$M$ The biggest absolute weight value.
$Q$ The LDPC maximum decoding iteration.
(a) Wireless FL system. Multiple mobile clients are coordinated by one remote server through the noisy wireless channels. The communicated parameters are model updates in the uplink and model weights in the downlink.
(b) System operations. Server: aggregation, quantization, LDPC encoding, client selection and broadcasting. Mobile devices: receiving model weights, LDPC decoding,local training and uplink transmission.
Figure 1: Wireless FL in digital communication, system outlook and underlying operations.
II-A Wireless FL
Fig. 1 depicts a wireless FL system, where model training of different clients is coordinated by the server. The FL problem can be formulated as follows [26, 27, 28]:
$\min_{\mathbf{w}\in{\mathbb{R}^{d}}}f(\mathbf{w})\triangleq\sum_{k=1}^{n}p_{k}f_{k}(\mathbf{w}),$ (1)
where $\mathbf{w}$ denotes the model parameters, $n$ represents the number of clients, $p_{k}$ is the ratio between the amount of training data samples of the $k$-th client and that of all the clients, and $f_{k}(\mathbf{w})$ is the local loss function of the $k$-th client.
To train a shared global model, the server communicates with the mobile clients for multiple rounds through noisy wireless channels. In each communication round, the global model at the server is broadcast to a set of selected clients. Each of the selected clients then performs local training based on the received global model and transmits the model updates to the server for aggregation. Specifically, in the $r$-th communication round, the global model is quantized by a stochastic quantizer, encoded by a channel code, and broadcast to clients. The model parameters received by the $k$-th client is given as follows:
$\mathbf{w}^{k}_{r,0}=\mathbf{w}_{r}^{\prime},$ (2)
where $\mathbf{w}_{r}^{\prime}$ denotes the noisy version of $\mathbf{w}_{r}$. The impact of BER can be quantified by considering different error bit positions in the model parameters and the detailed analysis is given in Appendix B. A selected client updates the received global model by conducting $E$ steps of local mini-batch Stochastic Gradient Descent (SGD) with a learning rate $\eta$ as follows:
$\mathbf{w}^{k}_{r,e+1}=\mathbf{w}^{k}_{r,e}-\eta{\nabla}f_{k}(\mathbf{w}^{k}_{r,e},\xi),e=0,\cdots,E-1,k\in\mathcal{S}_{r},$ (3)
where ${\nabla}f_{k}(\mathbf{w}^{k}_{r,e},\xi)$ denotes the gradient computed from the local mini-batch datasets $\xi$ in the $k$-th client, and $\mathcal{S}_{r}$ denotes the set of selected clients in the $r$-th communication round.
By completing $E$ steps of local training, the local model update of the $k$-th client can be expressed as follows:
${\Delta\mathbf{w}^{k}_{r}=\mathbf{w}^{k}_{r,E}-\mathbf{w}^{k}_{r,0}},$ (4)
which is quantized, encoded, and uploaded to the server. For simplicity, we assume an orthogonal uplink channel is available for each selected client. After receiving these noisy local updates, denoted as $(\Delta\mathbf{w}^{k}_{r})^{\prime}$, the server performs aggregation to update the global model as follows:
$\mathbf{w}_{r+1}=\mathbf{w}_{r}+\frac{1}{|\mathcal{S}_{r}|}\sum_{k\in{\mathcal{S}_{r}}}p_{k}(\Delta\mathbf{w}^{k}_{r})^{\prime}.$ (5)
In this work, we focus on the downlink communication rather than the uplink for two reasons. First, our objective is to design channel decoding schemes to save the energy consumption of mobile devices. For the uplink, the channel decoding happens on the server side which falls out of the scope of this study. Second, the server has a continuous power supply and is able to conduct a sufficient number of decoding iterations to achieve satisfactory BER performance.
II-B Energy Consumption Evaluation
There are two main energy-consuming operations for clients participating in wireless FL, namely model training and model communication. For the latter, it is believed that the energy is mostly consumed by the data transmission. In this section, we evaluate the energy consumption of different components, taking the training of AlexNet for image classification tasks as an example [29]. The analysis is based on actual measurements obtained from fabricated chips published in recent years [30, 31, 32]. For fair comparisons, all data are collected for chips fabricated in the 65 nm complementary metal-oxide semiconductor (CMOS) process.
In[30], a fabricated AI hardware accelerator was reported to achieve the energy efficiency of 13.7 mJ/epoch for 8-bit quantized AlexNet. With five epochs of local training, the processing energy consumption in each communication round is 68.5 mJ. In [31], a fifth-generation (5G) New Radio wireless transmitter and receiver were reported to achieve the energy efficiencies of 47.9 pJ/b and 33.3 pJ/b for transferring uncoded data, respectively, under a communication distance of around 200 meters. Therefore, the energy required to transmit and receive the 8-bit AlexNet with 60 million parameters in a single communication round is about 39 mJ. Assuming a code rate of 1/2, the communication energy consumption for coded transmission of the AlexNet model parameters in each round is 78 mJ. Besides, the LDPC decoder in [32] achieves the energy efficiency of 20.1 pJ/b/iteration, which was measured with 10 decoding iterations on average. Accordingly, the energy consumption of the LDPC decoder for decoding the AlexNet model parameters is 96.5 mJ. The detailed comparison is presented in Table II, which indicates that the energy consumption of LDPC decoding accounts for a large portion of the overall energy consumption in wireless FL.
TABLE II: Energy Consumption of Main Components at a Mobile Unit in Wireless FL for Training an AlexNet
Component Energy efficiency Energy consumption
AI Chip 13.7 mJ/epoch[30] 68.5 mJ
RF Transceiver 81.2 pJ/b[31] 78 mJ
LDPC Decoder 20.1 pJ/b/iter[32] 96.5 mJ
II-C LDPC Decoding
Figure 2: Typical LDPC decoder architecture: the variable nodes and check nodes are connected, and the messages are transmitted between them iteratively.
In this work, we use LDPC codes as an example to illustrate how adaptive channel decoding can help reduce the energy consumption of wireless FL. To that end, we first introduce a typical LDPC decoder architecture in Fig. 2, which contains the variable nodes representing the encoded bits to be corrected and check nodes representing the parity check equation. The check nodes exchange information with variable nodes and perform computations iteratively until achieving the maximum decoding iteration [33]. Such iterative processing leads to high energy consumption of LDPC decoding. To save decoding energy, early termination techniques are usually adopted. Specifically, if the parity-check matrix $\mathbf{H}$ and the codeword $\mathbf{x}$ satisfy $\mathbf{Hx}^{\rm{T}}=0$, the decoding process will stop even if the maximum decoding iteration has not been reached[34]. The early termination scheme can save decoding iterations while preserving BER performance. The energy consumption of LDPC decoding depends on the target BER. As a result, to minimize the energy consumption, it is essential to determine the maximum tolerable BER by FL systems, considering its unique statistical learning nature.
III Energy-Efficient LDPC Decoding Design
In this section, we propose an energy-efficient LDPC decoding scheme to minimize the decoding energy consumption while guaranteeing the same convergence rate for wireless FL as the case with error-free communication. For that purpose, we explore the relationship between BER and convergence rate, and that between the maximum number of decoding iterations and BER. By connecting these two relationships, we demonstrate how to regulate the LDPC decoding scheme.
III-A Effect of BER on the Convergence Rate of Wireless FL
With error-free communication, the convergence rate of FL with non-convex loss functions is proved as $\mathcal{O}(1/\sqrt{T})$ [35], where $T$ is the total number of training epochs. In this subsection, we investigate how BER affects the convergence rate of FL and determine the maximum admissible BER for maintaining the convergence rate of $\mathcal{O}(1/\sqrt{T})$.
Unlike FL with analog communication[36, 37, 38, 39, 40, 41, 42], quantifying the impact of BER on FL with digital communication is not straightforward. In particular, we need to consider different cases of bit errors and analyze the impact of different bit errors, e.g., most significant bits and least significant bits.
To conduct the theoretical analysis, we take the following assumptions that are widely adopted in the literature.
Assumption 1.
($L$-Smoothness [12, 35]) The local loss function $f_{k}\left(\cdot\right)$ at the $k$-th client is $L$-smooth, i.e., $\forall\mathbf{w},\hat{\mathbf{w}}$, we have $||\nabla f_{k}(\mathbf{w})-\nabla f_{k}(\hat{\mathbf{w}})||\leq L||\mathbf{w}-\hat{\mathbf{w}}||$.
Assumption 2.
(Unbiased and Bounded-variance Local Gradient Estimator [43]) Let $\xi_{r}^{k}$ be the batch of randomly sampled data at the $k$-th client in the $r$-th communication round. The gradient estimator $\nabla f_{k}(\mathbf{w};\xi_{r}^{k})$ is unbiased, i.e., $\mathbb{E}[\nabla f_{k}(\mathbf{w};\xi_{r}^{k})]=\nabla f_{k}(\mathbf{w})$, and has a bounded variance, i.e., $\mathbb{E}[||\nabla f_{k}(\mathbf{w};\xi_{r}^{k})-\nabla f_{k}(\mathbf{w})||^{2}]\leq\sigma_{L}^{2}$.
Assumption 3.
(Bounded Local and Global Gradient Dissimilarity [43]) The global dissimilarity between the local and global gradient is bounded, i.e., $||\nabla f_{k}(\mathbf{w})-\nabla f(\mathbf{w})||^{2}\leq\sigma_{G}^{2}$.
Assumption 4.
(Unbiased Quantization [12, 28, 35]) The random quantizer Q($\cdot$) is unbiased, i.e., $\mathbb{E}[Q(\mathbf{w})|\mathbf{w}]=\mathbf{w}$.
With Assumptions 1-4, the convergence bound of wireless FL with a given BER is given by the following theorem.
Theorem 1.
When the learning rate is set as $\eta=\frac{1}{L\sqrt{RE}}=\frac{1}{L\sqrt{T}}$, the average expected squared norm of global stochastic gradient is bounded as follows:
$\begin{split}\frac{1}{T}\sum_{r=1}^{R}\sum_{e=1}^{E}&\mathbb{E}\|\nabla f(\bar{\mathbf{w}}_{r,e})\|^{2}\\ \leq&\frac{C_{0}}{\sqrt{T}}+\frac{C_{1}\sum_{r=1}^{R}b_{r}}{\sqrt{T}}+\frac{C_{2}}{\sqrt{T}}+\frac{C_{3}}{\sqrt{T}},\end{split}$ (6)
where $\bar{\mathbf{w}}_{r,e}\triangleq\frac{1}{n}\sum_{k=1}^{n}\mathbf{w}_{r,e}^{k}$, $C_{0}\triangleq 2L(f(\mathbf{w}_{0})-f^{*})$, $C_{1}\triangleq\frac{8L^{2}dM^{2}}{3}$, $C_{2}=\sigma_{L}^{2}(\frac{1}{n}+\frac{4(n-s)}{s(n-1)})$, $C_{3}\triangleq\frac{\sigma_{G}^{2}}{L}$, $f^{*}$ is the minimum global training loss, $b_{r}$ is the BER in the $r$-th communication round, and $\mathbf{w}_{0}$ is the initial global model.
Proof.
The key idea is to first quantify the BER’s impact on the model error and then introduce the model error to the convergence analysis of vanilla learning process. The detailed proof is available in the Appendix. ∎
Remark 1.
It is worthy noting that the first term on the right-hand side (RHS) of (6) captures the difference between the initial and minimum global training loss. The second term indicates how bit errors in each communication round affect learning convergence. The third term comes from the random mini-batch sampling and partial client participation, and the fourth term shows the impact of imbalanced data distribution.
According to Theorem 1, in order to guarantee the convergence rate of $\mathcal{O}(1/\sqrt{T})$, it is crucial to control the BER in each communication round. Although the explicit relation between $b_{r}$ and $r$ is difficult to obtain, we know from the above discussion that, $b_{r}$ should be a decreasing function with $r$ as otherwise the second term on the RHS of (6) cannot be diminshed with $T$ increases. For ease of implementation, we propose a simple BER regulation with $b_{r}\propto\frac{1}{r^{2}}$, which is shown in the following corollary as a sufficient condition for a practical wireless FL system to achieve the same convergence rate as that with error-free communication.
Corollary 1.
If the BER in each communication round is regulated as follows:
$b_{r}\propto\frac{1}{r^{2}},r=1,2,\cdots,R,$ (7)
the RHS of (6) diminishes to zero at the rate of $\mathcal{O}(1/\sqrt{T})$, i.e., wireless FL converges to a local optimal solution of the global training loss funcation at a rate of $\mathcal{O}(1/\sqrt{T})$.
Proof.
To prove the RHS of (6) diminishes to zero at the rate of $\mathcal{O}(1/\sqrt{T})$, it is sufficient to verify
$\lim_{T\rightarrow\infty}\frac{\sum_{r=1}^{R}b_{r}}{\sqrt{T}}=0$ (8)
when $b_{r}\propto 1/r^{2}\triangleq\delta/r^{2}$. Since
$\begin{split}0&\leq\frac{\sum_{r=1}^{R}b_{r}}{\sqrt{T}}=\delta\frac{\sum_{r=1}^{R}1/r^{2}}{\sqrt{T}}\\ &=\frac{\delta}{\sqrt{T}}\left(1+\sum_{r=2}^{R}\frac{1}{r^{2}}\right)\\ &<\frac{\delta}{\sqrt{T}}\left(1+\sum_{r=2}^{R}\frac{1}{r(r-1)}\right)=\frac{\delta}{\sqrt{T}}\left(2-\frac{1}{R}\right)\\ &=\frac{\delta}{\sqrt{T}}\left(2-\frac{E}{T}\right),\end{split}$ (9)
we conclude (8) by letting $T\rightarrow\infty$ for both sides of (9) and appying the Sandwich Theorem. ∎
Corollary 1 indicates that gradually decreasing the BER at a rate of $\mathcal{O}(1/r^{2})$ can guarantee the convergence rate of wireless FL at a rate of $\mathcal{O}(1/\sqrt{T})$. The proposed BER scaling rule achieves an improving communication quality with communication rounds. This is consistent with the solution developed in [40], which improves the communication quality by increasing the signal-to-noise ratio (SNR) with communication round.
III-B BER vs. Number of Iterations
The analysis in Section III-A indicates that the convergence rate of wireless FL with error-free communication can be achieved if the BER is properly controlled. In this subsection, we analyze the relationship between BER and the maximum number of LDPC decoding iterations, based on which, we develop an adaptive decoding scheme to maintain the same convergence rate.
1 Input: The initial global model $\mathbf{w}_{0}$
2 Output: The converged global model $\mathbf{w}_{R}$
3 for  Communication round $r=1,2,...,R$ do
4       for Client index $k$ in $\{1,2,...,K\}$ do
5             $b_{r}\propto(1/r^{2})$ with $b_{1},b_{R}$
6             $Q=f^{-1}(b_{r})$
7             for $decoding.iter\leq Q$ do
8                  one iteration of LDPC decoding
9                   if Early termination is satisfied then
10                        break
11                   end if
12                  $decoding.iter+=1$
13             end for
14            output: $\mathbf{w}_{r}^{\prime}$
15             for local training step $e=1,...,E$ do
16                  $\mathbf{w}^{k}_{r,0}=\mathbf{w}_{r}^{\prime}$
17                   $\mathbf{w}^{k}_{r,e+1}=\mathbf{w}^{k}_{r,e}-\eta_{r}\nabla f(\mathbf{w}^{k}_{r,e})$
18             end for
19            $\Delta\mathbf{w}_{r}^{k}=\mathbf{w}^{k}_{r,E}-\mathbf{w}_{r,0}^{k}$
20       end for
21      $\mathbf{w}_{r+1}=\mathbf{w}_{r}+\sum_{k=1}^{K}p_{k}\Delta\mathbf{w}_{r}^{k}$
22 end for
Algorithm 1 Regulating LDPC decoder in FL
It has been shown in [44] that for message-passing-based channel decoders to achieve a fraction of 1-$\epsilon$ of the channel capacity over a binary erasure channel, the relationship between the decoding complexity (denoted as $\chi_{D}$) and BER (denoted as $b$) can be modeled as follows:
$\chi_{D}=\mathcal{O}\left(\log\frac{1}{b}\right)+\mathcal{O}\left(\frac{1}{\epsilon}\log\frac{1}{\epsilon}\right).$ (10)
According to [45], the relationship between the decoding complexity and the maximum number of decoding iterations ($Q$) can be modeled as follows:
$\chi_{D}=Q\cdot\chi_{0},$ (11)
where $\chi_{0}$ denotes the decoding complexity per iteration. By combining (10) and (11), it leads to the following relationship between the maximum number of decoding iterations and BER:
$Q=\frac{1}{\chi_{0}}\mathcal{O}\left(\log\frac{1}{b}\right)+\frac{1}{\chi_{0}}\mathcal{O}\left(\frac{1}{\epsilon}\log\frac{1}{\epsilon}\right).$ (12)
In the next subsection, we develop an adaptive scheme to control the maximum number of decoding iterations for clients in a wireless FL system based on Corollary 1 and (12).
III-C Decoding Iteration Control
In this subsection, we show how to adaptively control the maximum number of LDPC decoding iterations over communication rounds to achieve the desired BER.
In principle, by combining (7) and (12), we obtain the following relationship between the maximum number of LDPC decoding iterations and communication rounds:
$Q_{\color[rgb]{0,0,0}r}=\frac{1}{\chi_{0}}\mathcal{O}\left(2\log\ r\right)+\frac{1}{\chi_{0}}\mathcal{O}\left(\frac{1}{\epsilon}\log\frac{1}{\epsilon}\right),$ (13)
which provides a guideline to select the maximum number of LDPC decoding iterations in the $r$-th communication round. Specifically, $Q_{r}$ can be set as $\alpha\log r+\beta$, where $\alpha$ and $\beta$ are constants. However, (13) cannot be directly utilized because $\alpha$ and $\beta$ are code-specific and depend on the receive SNR as well as channel decoding algorithms. Thus, we propose a practical approach by a lookup table, which captures the relationship between BER and the maximum number of decoding iterations at different levels of receive SNR. Entries in the lookup table can be obtained through empirical measurements specific to the LDPC code used in the wireless FL system. Fig. 3 shows the sample simulation results for a rate 1/2 LDPC code with a block length of 1008 bits and binary phase shift keying (BPSK) modulation over an additional white Gaussian noise (AWGN) channel with $\text{SNR}=1.5$ and 2.5 dB. Data in this figure can be stored in a three-dimensional lookup table [46]. From the lookup table, we can get the corresponding maximum number of decoding iteration for a target BER for clients with different receive SNR.
Figure 3: The relation between BER and LDPC maximum decoding iterations under different SNRs.
The wireless FL procedures with adaptive LDPC decoding are summarized in Algorithm 1. The algorithm starts with an initial target BER, denoted as $b_{0}$, which is set to a relatively high value, i.e., $10^{-1}$ in our experiments. The target BER at the last communication round, denoted as $b_{R}$, is set as a small value, i.e., $10^{-4}$ in our experiments. The concrete expression of $b_{r}$ in our experiment is $b_{r}=\frac{0.0999R^{2}}{(R^{2}-1)r^{2}}+\frac{10^{-4}R^{2}-0.1}{R^{2}-1}$. For each communication round, the BER is scaled according to (7), and the corresponding maximum number of LDPC decoding iterations in that round is obtained from the lookup table.
Remark 2.
Note that the number of entries in the lookup table depends on the possible discrete BER and SNR values. Since this number is not large, the memory overhead for storing the lookup table is expected to be small. Additionally, reading data from the lookup table takes constant time and does not introduce additional LDPC decoding latency.
IV Experimental Results
In this section, we validate the effectiveness of the proposed adaptive LDPC decoding algorithm via numerical experiments. First, we introduce the experimental settings. Then, the performance of the proposed algorithm, including test accuracy and energy consumption, is compared with a fixed-iteration LDPC decoding baseline.
Figure 4: Fixed decoding schemes: On the mobile side, setting the LDPC maximum decoding iterations to 6, 12, 18, 24, and 52 can achieve a BER of $10^{-1}$, $10^{-2}$, $10^{-3}$, $10^{-4}$, and $10^{-5}$, respectively.
IV-A Experimental Settings
(a) Experiment of LeNet-300-100 on MNIST.
(b) Experiment of vanilla-CNN on Fashion-MNIST.
(c) Experiment of CNN on CIFAR-10.
(d) Experiment of vanilla-CNN on non-IID Fashion-MNIST.
Figure 5: Experiments on different datasets showing how fixed BER affects FL accuracy.
(a) Consumed decoding iteration for each frame.
(b) The distributed for different consumed iterations.
Figure 6: Example of vanilla-CNN on Fashion-MNIST. The 8-bit quantized vanilla CNN model goes through AWGN channel with $E_{b}/N_{0}$ of 2.5dB, and is decoded by LDPC of block-length of 1008 bit.
System Setups: We consider an FL system with 10 clients. Clients are assumed active in every training round and they communicate with the server using rate 1/2 min-sum LDPC code with a code length of 1008[47]. Besides, binary phase shift keying (BPSK) is adopted as the modulation scheme. In each training round, the number of local SGD steps is set as 5 with the learning rate and batch size as 0.01 and 64, respectively.
Datasets: We conduct experiments on three dataset, including MNIST[48], Fashion-MNIST[49], and CIFAR-10[50]. The training datasets were distributed among mobile devices in an independent and identically distributed (IID) manner. We further consider one non-IID experiment on Fashion-MNIST where each client was assigned only two categories of training data, similar to the setup in [27]. The test datasets were utilized for validation on the server side.
DNN Models: For the MNIST dataset, the LeNet-300-100 model[51] is employed, which is a fully connected neural network consisting of two hidden layers and one softmax layer, with approximately 0.27 million parameters. For the Fashion-MNIST dataset, a vanilla CNN [3] with two 5×5 convolution layers, one fully connected layer, and one softmax layer is employed, comprising around 0.58 million parameters. For the CIFAR-10 dataset, a 7-layer CNN model with four 3×3 convolution layers, two fully connected layers, and one softmax layer is employed, which contains approximately 9.07 million parameters.
IV-B Learning Performance of Conventional Schemes
(a) Max.iter.
(b) Learning curves.
(c) Average iterations.
(d) BER.
(e) Final learning accuracy.
(f) Total decoding energy.
Figure 7: Experiment for LeNet-300-100 on MNIST. Compared with the conventional fixed schemes, our method that adaptively increases the maximum decoding iteration to 24 can achieve the same final accuracy of 98%, and reduce the decoding energy by 6.7mJ.
(a) Max.iter.
(b) Learning curves.
(c) Average iterations.
(d) BER.
(e) Final learning accuracy.
(f) Total decoding energy.
Figure 8: Experiment for vanilla CNN on Fashion-MNIST. Compared with the conventional fixed schemes, our method that adaptively increases the maximum decoding iteration to 24 can achieve a similar final accuracy of 91%, and reduce the decoding energy by 14.1mJ.
In this subsection, the experimental results for conventional schemes that use fixed decoding iterations are presented. As shown in Fig. 4, the fixed decoding iterations lead to fixed BERs in the whole training process.
Next, the impact of fixed BER on FL performance was investigated. Fig. 5 summarizes the test accuracy with different fixed number of maximum decoding iterations. It can be observed that for all four FL tasks, when BER is higher than $10^{-3}$, the learning performance was significantly impacted. Lowering the BER to $10^{-4}$ leads to improved accuracy and convergence towards perfect communication levels. However, further decreasing the BER to $10^{-5}$ will not yield any significant improvement in learning performance.
IV-C Learning Performance of the Proposed Approach
In this subsection, the proposed BER scaling algorithm is compared to the fixed BER scheme. The proposed approach uses a scaling BER of $\mathcal{O}(1/r^{2})$ by controlling the maximum decoding iteration in each communication round based on a lookup table that summarizes the relationship between the BER and decoding iterations.
(a) Max.iter.
(b) Learning curves.
(c) Average iterations.
(d) BER.
(e) Final learning accuracy.
(f) Total decoding energy.
Figure 9: Experiment for 7-layer CNN on CIFAR-10. Compared with the conventional fixed schemes, our method that adaptively increases the maximum decoding iteration to 24 can achieve a similar final accuracy of 77.6%, and reduce the decoding energy by 190mJ.
(a) Max.iter.
(b) Learning curves.
(c) Average iterations.
(d) BER.
(e) Final learning accuracy.
(f) Total decoding energy.
Figure 10: Experiment for vanilla CNN on non-IID Fashion-MNIST. Compared with the conventional fixed schemes, our method that adaptively increases the maximum decoding iteration to 24 can achieve a similar final accuracy of 54.1%, and reduce the decoding energy by 13.9mJ.
Specifically, the learning process is assigned with an initial BER value of $10^{-1}$ and a target BER value of $10^{-4}$. Based on the BER scaling rule of $\mathcal{O}(1/r^{2})$, the desired BER at every communication round can be calculated. Then based on the lookup table, we can control the LDPC maximum decoding iteration for every communication round. In the experiments, the initial and final maximum decoding iterations are set to 6 and 24, respectively. The learning performance and energy consumption of the proposed adaptive scheme are compared to those of fixed schemes with fixed BER of $10^{-4}$. The corresponding fixed maximum decoding iterations is set to 24. To mitigate energy consumption, the LDPC decoder employs an early termination scheme, where the decoding process stops when the bitstream in a frame is deemed correct. Thus, only frames that cannot be correctly decoded require the maximum decoding iteration. Fig. 6 shows an example for the early termination scheme by using the experiment of vanilla-CNN on Fashion-MNIST, illustrating the actual decoding iterations for each bitstream. It can observed that only a few bitstreams use up the maximum decoding iterations which is set to 52. The average decoding iterations is the average number of real consumed decoding iterations. Given the decoding energy is proportional to the decoding iterations, the decoding energy consumption can be reflected by the average number of decoding iterations used for the FL task.
Fig. 7 summarises the experimental results for LeNet-300-100 on MNIST. Fig. 7(a) shows how the maximum number of LDPC decoding iterations varies with the FL communication rounds for the proposed approach compared with other fixed schemes. The traditional schemes adopt a fixed number of maximum decoding iterations, resulting in a fixed BER, while the proposed scheme adopts a roughly linearly increasing number of maximum decoding iterations to achieve the designed BER. Fig. 7(b) shows the learning accuracy curves for different schemes. Setting the fixed maximum decoding iteration to 6, which corresponds to a BER of $10^{-1}$, severely affects the training accuracy. Similarly, when the maximum decoding iteration is set to 12, resulting in a BER of about $10^{-2}$, the training still fails to converge. With a maximum decoding iteration of 18, corresponding to a BER of about $10^{-3}$, the training can converge, but with a noticeable loss in accuracy. The training converges well to high learning accuracy when the maximum decoding iteration is increased to 24, resulting in a BER of $10^{-4}$. Further increasing the maximum decoding iteration to 52 does not lead to further accuracy improvement.
The proposed scheme, which approximately increases the maximum decoding iteration from 6 to 24 in a linear manner, enables the training accuracy to improve rapidly and converge to the same level as the fixed scheme with 24/52 iterations. Fig. 7(c) illustrates how the average number of decoding iterations varies during training. The proposed scheme can reduce the average number of iterations and thus the overall energy consumption. Fig. 7(d) depicts the BER of the proposed scheme, which scales as $\mathcal{O}(1/r^{2})$ from $10^{-1}$ to $10^{-4}$. Fig. 7(e) shows the final accuracy of different schemes. It can be observed that the proposed approach achieves the same accuracy as the fixed schemes that always maintain a very low BER of $10^{-4}$ and $10^{-5}$ throughout the training process. Fig. 7(f) illustrates the decoding energy consumption on the mobile side for the different schemes. It is obtained by multiplying the total number of LDPC decoding iterations with the LDPC decoding energy efficiency reported in [32]. Compared to the fixed schemes using a maximum decoding iteration of 24, the proposed scaling scheme can reduce the overall decoding energy by 19.4%.
Figs. 8, 9, and 10 present the experimental results for Fashion-MNIST, CIFAR-10, and non-IID Fashion-MNIST, respectively. The results are similar to those obtained by the MNIST experiment and show that the adaptive scheme can achieve similar or even slightly better accuracy compared to the fixed schemes. The observation of slightly better learning accuracy was also reported in [52, 53, 54], which suggests that adding noise to the model could enhance its generalization ability, and annealed noise could help achieve better learning accuracy. Overall, the proposed scheme reduces the LDPC decoding energy for Fashion-MNIST, CIFAR-10, and non-IID Fashion-MNIST by 18.3%, 20.1%, and 18.0%, respectively. Table III provides a summary of the experimental data.
V Conclusion
In this paper, we considered the energy consumption issue in federated learning (FL) over digital communication networks. The study recognized that channel decoding contributes to a large portion of the energy consumption for mobile FL, which has not been previously highlighted. It was shown that, due to the robustness of FL to communication errors, properly control the BER over different communication rounds can achieve the same convergence rate as error-free FL. Based on this analysis, we proposed an energy-efficient LDPC decoding iteration control algorithm. Experimental results demonstrated that the proposed method can reduce approximately 20% the total energy consumed by LDPC decoding, while maintaining the same or even a slightly higher learning accuracy compared to the scheme without decoding iteration control.
TABLE III: Experimental Results of Different Schemes. Our scheme can achieve similar accuracy while saving around 20% of the decoding energy compared to the best-fixed scheme.
Final learning accuracy
Fixed-18 Fixed-24 Fixed-52 Ours
MNIST 96.74% 97.96% 98.06% 98.02%
Fashion-MNIST 86.18% 90.63% 91.24% 91.16%
CIFAR-10 76.72% 77.38% 77.57% 77.68%
non-IID Fashion 51.3% 54.1% 54.3% 54.2%
Total consumed decoding energy (mJ)
Fixed-18 Fixed-24 Fixed-52 Ours
MNIST 33.87 34.51 34.68 27.81
Fashion-MNIST 74.23 76.10 76.29 62.17
CIFAR-10 925.10 948.44 951.36 757.80
non-IID Fashion 74.22 76.11 76.27 62.38
Appendix
V-A Proof of Theorem 1
Proof Outline: To prove Theorem 1, we first derive four lemmas, namely Lemmas 1$\sim$4, to establish the relationship between the convergence rate and model error. Then Lemma 5 builds the connection between model error and BER. By combining these relations, Theorem 1 is proved, which presents the relationship between convergence rate and BER.
For each communication round $r=1,...,R$ and local iteration $e=1,...,E$, we have
$\mathbf{w}_{r+1}=\mathbf{w}_{r}+\frac{1}{s}\sum_{k\in{S_{r}}}(\mathbf{w}_{r,E}^{k}-(Q(\mathbf{w}_{r}))^{\prime}),$
$\hat{\mathbf{w}}_{r+1}=\mathbf{w}_{r}+\frac{1}{n}\sum_{k\in{[n]}}(\mathbf{w}_{r,E}^{k}-(Q(\mathbf{w}_{r}))^{\prime}),$
$\bar{\mathbf{w}}_{r,e}=\frac{1}{n}\sum_{k\in{[n]}}\mathbf{w}_{r,e}^{k},$ (14)
where $\mathbf{w}_{r+1}$ represents the updated global model based on $s$ selected clients, $\hat{\mathbf{w}}_{r+1}$ denotes the updated global model based on all $n$ clients, and $\bar{\mathbf{w}}_{r,e}$ denotes the average model of $n$ local models. $(Q(\mathbf{w}_{r}))^{\prime}$ represents that the downlink model $\mathbf{w}_{r}$ will be impacted by the quantization operation $Q(\cdot)$ and channel noise $(\cdot)^{\prime}$ before being received by the clients. To simplify the representation, we will use $Q^{\prime}(\cdot)$ to replace $(Q(\cdot))^{\prime}$ in subsequent derivations.
Lemma 1. Assuming that Assumptions 1 and 4 hold, for any communication round $r=1,...,R$, we have
$\mathbb{E}f(\mathbf{w}_{r+1})$ $\leq\mathbb{E}f(\bar{\mathbf{w}}_{r,E})+\frac{L}{2}\mathbb{E}\|\hat{\mathbf{w}}_{r+1}-\bar{\mathbf{w}}_{r,E}\|^{2}$
$+\frac{L}{2}\mathbb{E}\|\hat{\mathbf{w}}_{r+1}-\mathbf{w}_{r+1}\|^{2}.$ (15)
This lemma provides a bound on the training loss of the updated global model. This bound can be derived from the $L$-smooth property of the loss function on the server side. The proof is given in Appendix-B.
In the next three lemmas, we derive the upper bounds for the three terms on the RHS of (15).
Lemma 2. Assuming that Assumptions 1 - 3 hold and considering the sequence of updates with step size $\eta$, the following inequality holds for the case of noisy downlink transmission:
$\mathbb{E}f(\bar{\mathbf{w}}_{r,E})\leq\frac{L}{2}\mathbb{E}\|Q^{\prime}(\mathbf{w}_{r})-\mathbf{w}_{r}\|^{2}-\frac{\eta}{2}\sum_{e=1}^{E}\mathbb{E}\|\nabla f(\bar{\mathbf{w}}_{r,e})\|^{2}$
$+\mathbb{E}f(\mathbf{w}_{r})-\frac{\eta}{2n}(1-L\eta)\sum_{e=1}^{E}\sum_{k\in[n]}\mathbb{E}\|\nabla f(\mathbf{w}_{r,e}^{k})\|^{2}$
$+\frac{LE\eta^{2}\sigma_{L}^{2}}{2n}+\frac{E\eta^{2}\sigma_{G}^{2}}{2}.$ (16)
The proof is given in Appendix-B. This lemma captures the process of noisy broadcasting and local updates. As shown in the RHS of Eq.(16), the first two terms arise from the noisy broadcasting, while the subsequent terms come from the local training process.
Lemma 3. For $\hat{\mathbf{w}}_{r+1}$ and $\bar{\mathbf{w}}_{r,E}$ defined in (14), we have
$\mathbb{E}\|\hat{\mathbf{w}}_{r+1}-\bar{\mathbf{w}}_{r,E}\|^{2}=\mathbb{E}\|\mathbf{w}_{r}-Q^{\prime}(\mathbf{w}_{r})\|^{2}.$ (17)
This lemma indicates that the expectation of the global model error on the server side is the same as the downlink model error on client side. The proof can be found in Appendix-B.
Lemma 4. For $\mathbf{w}_{r+1}$ and $\hat{\mathbf{w}}_{r+1}$ defined in (14), we have
$\mathbb{E}\|\mathbf{w}_{r+1}-$ $\hat{\mathbf{w}}_{r+1}\|^{2}\leq\frac{4}{s(n-1)}(1-\frac{s}{n})\cdot$
$\left[\eta^{2}\sigma_{L}^{2}En+\eta^{2}E\sum_{e=1}^{E}\sum_{k\in[n]}\|\nabla f(\mathbf{w}_{r,e}^{k})\|^{2}\right].$ (18)
This lemma captures the effect of partial participation of clients. The proof is provided in Appendix-B.
By substituting Lemmas 2-4 into Lemma 1, we derive the following recursive inequality for the updated model at the server:
$\mathbb{E}f(\mathbf{w}_{r+1})\leq L\mathbb{E}\|\mathbf{w}_{r}-Q^{\prime}(\mathbf{w}_{r})\|^{2}-\frac{\eta}{2}\sum_{e=1}^{E}\mathbb{E}\|\nabla f(\bar{\mathbf{w}}_{r,e})\|^{2}$
$-\frac{\eta}{2n}\left\{1-L\eta-\frac{4LE(n-s)}{s(n-1)}\right\}\cdot\sum_{e=1}^{E}\sum_{k\in[n]}\mathbb{E}\|\nabla f(\mathbf{w}_{r,e}^{k})\|^{2}$
$+\mathbb{E}f(\mathbf{w}_{r})+\frac{\eta E\sigma_{G}^{2}}{2}+L\eta^{2}\sigma_{L}^{2}E(\frac{1}{2n}+\frac{2(n-s)}{s(n-1)}).$ (19)
When $\eta$ satisfies the condition
$1-L\eta-\frac{4LE(n-s)}{s(n-1)}\geq 0,$ (20)
we have
$\mathbb{E}f(\mathbf{w}_{r+1})\leq L\mathbb{E}\|\mathbf{w}_{r}-Q^{\prime}(\mathbf{w}_{r})\|^{2}-\frac{\eta}{2}\sum_{e=1}^{E}\mathbb{E}\|\nabla f(\bar{\mathbf{w}}_{r,e})\|^{2}$
$+\mathbb{E}f(\mathbf{w}_{r})+\frac{E\eta^{2}\sigma_{G}^{2}}{2}+L\eta^{2}\sigma_{L}^{2}E(\frac{1}{2n}+\frac{2(n-s)}{s(n-1)}).$ (21)
Summing (21) over $r=1,...,R$ and rearranging the terms, we obtain the following expression:
$\frac{\eta}{2}\sum_{r=1}^{R}\sum_{e=1}^{E}\mathbb{E}\|\nabla f(\bar{\mathbf{w}}_{r,e})\|^{2}\leq L\sum_{r=1}^{R}\mathbb{E}\|\mathbf{w}_{r}-Q^{\prime}(\mathbf{w}_{r})\|^{2}$
$+f(\mathbf{w}_{0})-f^{*}+\frac{RE\eta^{2}\sigma_{G}^{2}}{2}+RL\eta^{2}\sigma_{L}^{2}E(\frac{1}{2n}+\frac{2(n-s)}{s(n-1)}).$ (22)
Multiplying $\frac{2}{RE\eta}$ on both sides of (22), we get
$\frac{1}{RE}\sum_{r=1}^{R}\sum_{e=1}^{E}\mathbb{E}\|\nabla f(\bar{\mathbf{w}}_{r,e})\|^{2}\leq\frac{2L}{RE\eta}\sum_{r=1}^{R}\mathbb{E}\|\mathbf{w}_{r}-Q^{\prime}(\mathbf{w}_{r})\|^{2}$
$+\frac{2(f(\mathbf{w}_{0})-f^{*})}{RE\eta}+\eta\sigma_{G}^{2}+L\eta\sigma_{L}^{2}(\frac{1}{n}+\frac{4(n-s)}{s(n-1)}).$ (23)
Since a high quantization bit-length is assumed to be used in this work with negligible quantization error, we have
$\mathbb{E}\|\mathbf{w}_{r}-Q^{\prime}(\mathbf{w}_{r})\|^{2}$ $=\mathbb{E}\|Q^{\prime}(\mathbf{w}_{r})-Q(\mathbf{w}_{r})+Q(\mathbf{w}_{r})-\mathbf{w}_{r}\|^{2}$
$=\mathbb{E}\|(\mathbf{w}_{r})^{\prime}-\mathbf{w}_{r}\|^{2}.$ (24)
Combining (23) and (24), we obtain
$\frac{1}{RE}\sum_{r=1}^{R}\sum_{e=1}^{E}\mathbb{E}\|\nabla f(\bar{\mathbf{w}}_{r,e})\|^{2}\leq\frac{2L}{RE\eta}\sum_{r=1}^{R}\mathbb{E}\|(\mathbf{w}_{r})^{\prime}-\mathbf{w}_{r}\|^{2}$
$+\frac{2(f(\mathbf{w}_{0})-f^{*})}{RE\eta}+\eta\sigma_{G}^{2}+L\eta\sigma_{L}^{2}(\frac{1}{n}+\frac{4(n-s)}{s(n-1)}).$ (25)
Lemma 5. For the perfect global model $\mathbf{w}$ and the received noisy model $(\mathbf{w})^{\prime}$, the model error can be bounded by
$\mathbb{E}||(\mathbf{w})^{\prime}-\mathbf{w}||^{2}\leq\frac{4d(4^{N}-1)}{3(2^{N}-1)^{2}}\cdot b\cdot M^{2},$ (26)
where $M$ denotes the maximum absolute value among all weight elements, i.e. $,M=max(|w_{1}|,|w_{2}|,$$,|w_{d}|)$. Lemma 5 gives the relationship between model variance and BER. The proof can be found in Appendix-B.
By substituting Eq.(26) into Eq.(25), we establish the relationship between the FL convergence rate and BER:
$\frac{1}{RE}\sum_{r=1}^{R}\sum_{e=1}^{E}\mathbb{E}\|\nabla f(\bar{\mathbf{w}}_{r,e})\|^{2}\leq\frac{8LdM^{2}}{3RE\eta}\sum_{r=1}^{R}\frac{b_{r}(4^{N}-1)}{(2^{N}-1)^{2}}$
$+\frac{2(f(\mathbf{w}_{0})-f^{*})}{RE\eta}+\eta\sigma_{G}^{2}+L\eta\sigma_{L}^{2}(\frac{1}{n}+\frac{4(n-s)}{s(n-1)}).$ (27)
When the quantization bit-length is high, we have $\frac{4^{N}-1}{(2^{N}-1)^{2}}\approx 1$, and Eq.(27) can be simplified as follows:
$\frac{1}{RE}\sum_{r=1}^{R}\sum_{e=1}^{E}\mathbb{E}\|\nabla f(\bar{\mathbf{w}}_{r,e})\|^{2}\leq\frac{2(f(\mathbf{w}_{0})-f^{*})}{RE\eta}$
$+\frac{8LdM^{2}}{3RE\eta}\sum_{r=1}^{R}b_{r}+L\eta\sigma_{L}^{2}(\frac{1}{n}+\frac{4(n-s)}{s(n-1)})+\eta\sigma_{G}^{2},$ (28)
which completes the proof of Theorem 1.
V-B Proof of Lemmas
1) Proof of Lemma 1
Quantization is unbiased according to Assumption 3, i.e., $\mathbb{E}_{Q}{(\mathbf{w})}=\mathbf{w}$. It is proved in Eq.(44) that the impact of the BER is also unbiased. So we have $\mathbb{E}_{Q^{\prime}}{Q^{\prime}(\mathbf{w})}=\mathbf{w}$. Based on the above unbiased properties and the notations defined in Eq.(14), we have
$\mathbb{E}_{Q^{\prime}}\hat{\mathbf{w}}_{r+1}$ $=\mathbf{w}_{r}+\frac{1}{n}\sum_{k\in[n]}(\mathbf{w}_{r,E}^{k}-\mathbb{E}_{Q^{\prime}}Q^{\prime}(\mathbf{w}_{r}))$
$=\mathbf{w}_{r}+\frac{1}{n}\sum_{k\in[n]}(\mathbf{w}_{r,E}^{k}-\mathbf{w}_{r})$
$=\frac{1}{n}\sum_{k\in[n]}\mathbf{w}_{r,E}^{k}$
$=\bar{\mathbf{w}}_{r,E}.$ (29)
We can also get
$\mathbb{E}_{S_{r}}\mathbf{w}_{r+1}$ $=\mathbf{w}_{r}+\frac{1}{n}\sum_{k\in[n]}(\mathbf{w}_{r,E}^{k}-Q^{\prime}(\mathbf{w}_{r}))=\hat{\mathbf{w}}_{r+1},$ (30)
whose proof is similar to Section 7.1 in [35].
Since the loss function $f$ is $L$-smooth, for any variables $\mathbf{x,y}$, we have
$f(\mathbf{x})\leq f(\mathbf{y})+\left\langle\nabla f(\mathbf{y}),\mathbf{x}-\mathbf{y}\right\rangle+\frac{L}{2}||\mathbf{x}-\mathbf{y}||^{2}.$ (31)
Taking $\mathbf{x}=\hat{\mathbf{w}}_{r+1}$ and $\mathbf{y}=\mathbf{\bar{w}}_{r,E}$ in (31), we can get
$f(\hat{\mathbf{w}}_{r+1})\leq$ $f(\mathbf{\bar{w}}_{r,E})+\left\langle\nabla f(\mathbf{\bar{w}}_{r,E}),\hat{\mathbf{w}}_{r+1}-\mathbf{\bar{w}}_{r,E}\right\rangle$
$+\frac{L}{2}||\hat{\mathbf{w}}_{r+1}-\mathbf{\bar{w}}_{r,E}||^{2}.$ (32)
We take the expectation on both sides of (32) and, since $\mathbb{E}_{Q^{\prime}}\hat{\mathbf{w}}_{r+1}=\bar{\mathbf{w}}_{r,E}$ (see (29)), we have
$\mathbb{E}f(\hat{\mathbf{w}}_{r+1})\leq\mathbb{E}f(\mathbf{\bar{w}}_{r,E})+\frac{L}{2}\mathbb{E}||\hat{\mathbf{w}}_{r+1}-\mathbf{\bar{w}}_{r,E}||^{2}.$ (33)
Taking $\mathbf{x}=\mathbf{w}_{r+1}$ and $\mathbf{y}=\hat{\mathbf{w}}_{r+1}$ in (31), and following the same steps as in (32) and (33), we have
$\mathbb{E}f(\mathbf{w}_{r+1})\leq\mathbb{E}f(\hat{\mathbf{w}}_{r+1})+\frac{L}{2}\mathbb{E}||\hat{\mathbf{w}}_{r+1}-\mathbf{w}_{r+1}||^{2}.$ (34)
By combining (33) and (34), we can get Lemma 1.
2) Proof of Lemma 2
Based on the proof in Section 8.3 of [35], and considering the non-iid case as in Assumption 3, if the downlink communication is perfect, we can obtain the upper bound of FL convergence as
$\mathbb{E}$ $f(\bar{\mathbf{w}}_{r,E})\leq\mathbb{E}f(\mathbf{w}_{r})-\frac{\eta}{2}\sum_{e=1}^{E}\mathbb{E}\|\nabla f(\bar{\mathbf{w}}_{r,e})\|^{2}$
$-\eta(\frac{1}{2n}-\frac{L\eta}{2n}-\frac{L^{2}\eta^{2}E(E-1)}{n})\sum_{e=1}^{E}\sum_{k\in[n]}\mathbb{E}\|\nabla f(\mathbf{w}_{r,e}^{k})\|^{2}$
$+\frac{LE\eta^{2}\sigma_{L}^{2}}{2n}+\frac{L^{2}\eta^{3}\sigma_{L}^{2}(n+1)E(E-1)}{2n}.$ (35)
For noisy downlink communication, the received model will have quantization and BER noise. As a result, (35) can be rewritten as
$\mathbb{E}$ $f(\bar{\mathbf{w}}_{r,E})\leq\mathbb{E}f(Q^{\prime}(\mathbf{w}_{r}))-\frac{\eta}{2}\sum_{e=1}^{E}\mathbb{E}\|\nabla f(\bar{\mathbf{w}}_{r,e})\|^{2}$
$-\eta(\frac{1}{2n}-\frac{L\eta}{2n}-\frac{L^{2}\eta^{2}E(E-1)}{n})\sum_{e=1}^{E}\sum_{k\in[n]}\mathbb{E}\|\nabla f(\mathbf{w}_{r,e}^{k})\|^{2}$
$+\frac{LE\eta^{2}\sigma_{L}^{2}}{2n}+\frac{E\eta^{2}\sigma_{G}^{2}}{2}.$ (36)
Since $f$ is $L$-smooth, according to (31), we have
$f(Q^{\prime}(\mathbf{w}_{r}))\leq$ $f(\mathbf{w}_{r})+\left\langle\nabla f(\mathbf{w}_{r}),Q^{\prime}(\mathbf{w}_{r})-\mathbf{w}_{r}\right\rangle$
$+\frac{L}{2}||Q^{\prime}(\mathbf{w}_{r})-\mathbf{w}_{r}||^{2}.$ (37)
Since the quantization and BER are both unbiased, i.e., $\mathbb{E}_{Q^{\prime}}{Q^{\prime}(\mathbf{w})}=\mathbb{E}_{Q}{Q(\mathbf{w})}=\mathbf{w}$, we take the expectation on both sides of (37) to obtain
$\mathbb{E}f(Q^{\prime}(\mathbf{w}_{r}))\leq\mathbb{E}f(\mathbf{w}_{r})+\frac{L}{2}\mathbb{E}||Q^{\prime}(\mathbf{w}_{r})-\mathbf{w}_{r}||^{2}.$ (38)
Combining (36) and (38), we prove Lemma 2.
3) Proof of Lemma 3
According to the notations defined in Eq.(14), we can get the expanded representation of $\hat{\mathbf{w}}_{r+1}$ and $\bar{\mathbf{w}}_{r,E}$ and derive
$\mathbb{E}\|\hat{\mathbf{w}}_{r+1}-\bar{\mathbf{w}}_{r,E}\|^{2}$
$=\mathbb{E}\|\mathbf{w}_{r}+\frac{1}{n}\sum_{k\in[n]}(\mathbf{w}_{r,E}^{k}-Q^{\prime}(\mathbf{w}_{r}))-\frac{1}{n}\sum_{k\in[n]}\mathbf{w}_{r,E}^{k}\|^{2}$
$=\mathbb{E}\|\mathbf{w}_{r}+\frac{1}{n}\sum_{k\in[n]}\mathbf{w}_{r,E}^{k}-\frac{1}{n}\sum_{k\in[n]}Q^{\prime}(\mathbf{w}_{r})-\frac{1}{n}\sum_{k\in[n]}\mathbf{w}_{r,E}^{k}\|^{2}$
$=\mathbb{E}\|\mathbf{w}_{r}-Q^{\prime}(\mathbf{w}_{r})\|^{2},$ (39)
which concludes the proof of Lemma 3.
4) Proof of Lemma 4
Similar to Section 8.5 in [35] and considering the noise in the downlink global model, we have
$\mathbb{E}\|\mathbf{w}_{r+1}-\hat{\mathbf{w}}_{r+1}\|^{2}$ $\leq\frac{4}{s(n-1)}(1-\frac{s}{n})\cdot$
$\sum_{k\in[n]}\mathbb{E}\|\mathbf{w}^{k}_{r,E}-Q^{\prime}(\mathbf{w}_{r})\|^{2}.$ (40)
Based on $E$ steps of local training with learning rate $\eta$, we have
$\mathbb{E}\|\mathbf{w}^{k}_{r,E}-Q^{\prime}(\mathbf{w}_{r})\|^{2}\leq\eta^{2}E(\sigma_{L}^{2}+\sum_{e=1}^{E}\|\nabla f(\mathbf{w}_{r,e}^{k})\|^{2}).$ (41)
Combining (40) and (41), we can conclude Lemma 4.
5) Proof of Lemma 5
Due to communication errors, the received DNN model for local training will contain random errors and differ from the global model, as shown in Fig. 11. In the following, we analyze the effect of the BER on the model error.
Figure 11: The difference between received noisy model and broadcasted perfect global model.
(a) BER=1E-2
(b) BER=1E-3
(c) BER=1E-4
Figure 12: Validation of the model error formula.
For a quantization scheme using $N$ bits to represent the weight $w$, the probability of having $x$ bit errors is given by
$P(X=x)=C_{N}^{x}b^{x}(1-b)^{N-x}.$ (42)
Since $b$ is much smaller than 1, the following relations can be established:
$P(X=0)\gg P(X=1)\gg P(X=2)\gg...\gg P(X=8).$ (43)
Therefore, the focus is on the case with at most one error bit.
The bit error can occur at various positions, ranging from the Least Significant Bit (LSB) to the Most Significant Bit (MSB), leading to different levels of error. For instance, for unsigned quantization, if the erroneous bit appears in the MSB and flips from 0 to 1, the error would be $+2^{N-1}\Delta$, where $\Delta=\frac{max(\mathbf{w})-min(\mathbf{w})}{2^{N}-1}=\frac{range(\mathbf{w})}{2^{N}-1}$, denoting the bin value for a given quantization. Alternatively, if the MSB flips from 1 to 0, the quantified error will be $-2^{N-1}\Delta$. The results for the case that the error occurs at the LSB will be $+2^{0}\Delta$ and $-2^{0}\Delta$. Additionally, it is assumed that the numbers of 0 and 1 in the bitstreams are similar. Based on this analysis, we can determine all the cases of bit flipping, and calculate the resulting values of the noisy weight with the corresponding probabilities, as shown in Table IV.
TABLE IV: Probability mass function of $(w)^{\prime}$.
$(w)^{\prime}_{i}$ $w+2^{0}\Delta$ $w-2^{0}\Delta$ $w+2^{1}\Delta$ $w-2^{1}\Delta$
$P_{i}$ b/2 b/2 b/2 b/2
$(w)^{\prime}_{i}$ $w+2^{N-1}\Delta$ $w-2^{N-1}\Delta$ $w$
[…]
Based on Table IV, the mean and variance of the received noisy weight $(w)^{\prime}$ can be calculated as follows:
$\mathbb{E}((w)^{\prime})=\sum_{i}((w)^{\prime}_{i}\cdot P_{i})=w,$ (44)
$\mathbb{E}||(w)^{\prime}-w||^{2}$ $=\sum_{i}[((w)^{\prime}_{i}-w)^{2}\cdot P_{i}]$
$=\frac{4^{N}-1}{3(2^{N}-1)^{2}}\cdot b\cdot range(\mathbf{w})^{2}.$ (45)
The above equations indicate that, for a weight parameter in the DNN model, the mean value of the noisy parameter is equal to the broadcasted perfect parameter. Furthermore, the variance between the two parameters is a function of the number of quantization bits, the BER level, and the range of the model parameters. For a $d-$dimensional model $\mathbf{w}$, i.e., $\mathbf{w}=[w_{1},w_{2},...,w_{d}]$, the following equations can be derived:
$\mathbb{E}||(\mathbf{w})^{\prime}-\mathbf{w}||^{2}$ $=\sum_{j=1}^{d}\mathbb{E}||(w_{j})^{\prime}-w_{j}||^{2}$
$=\frac{(4^{N}-1)d}{3(2^{N}-1)^{2}}\cdot b\cdot range(\mathbf{w})^{2}.$ (46)
To verify the theoretical results of Eq.(46), simulations were conducted using an 8-bit vanilla-CNN on the Fashion-MNIST dataset. The experimental settings are described in Section IV-A. Three different BER conditions were simulated, and the model errors obtained from the simulation are plotted in Fig. 12. The expectation on the left-hand side (LHS) of Eq.(46) corresponds to different cases of errors. The model errors predicted using the proposed formula on the RHS of Eq.(46) are also plotted in the same figure. It can be seen that, for different BER conditions and different communication rounds, the results obtained from the theoretical formula closely match those obtained from simulation, confirming the accuracy of the derived Eq.(46).
If we denote $M$ as the maximum absolute value of all weight elements, i.e. $,M=max(|w_{1}|,|w_{2}|,$$,|w_{d}|)$, we have $range(\mathbf{w})\leq 2M$. The upper bound of Eq.(46) can be obtained as follows:
$\mathbb{E}||(\mathbf{w})^{\prime}-\mathbf{w}||^{2}\leq\frac{4d(4^{N}-1)}{3(2^{N}-1)^{2}}\cdot b\cdot M^{2},$ (47)
which concludes the proof of Lemma 5.

### v2

@@
 Abstract
-One of the most critical challenges for deploying distributed learning, such as federated learning (FL), in wireless networks is the limited battery capacity of mobile devices.
-While it is a common belief that the major energy consumption of mobile devices comes from the uplink data transmission, this paper presents a novel finding, namely the channel decoding operation also contributes significantly to the overall energy consumption of mobile devices in FL.
+One of the most critical challenges for deploying distributed learning solutions, such as federated learning (FL), in wireless networks is the limited battery capacity of mobile clients.
+While it is a common belief that the major energy consumption of mobile clients comes from the uplink data transmission, this paper presents a novel finding, namely the channel decoding operation also contributes significantly to the overall energy consumption of mobile clients in FL.
 Motivated by this new observation, we propose an energy-efficient adaptive channel decoding scheme that leverages the intrinsic robustness of FL to model errors.
-In particular, the robustness is exploited to reduce the energy consumption of channel decoders at mobile devices by adaptively adjusting the number of decoding iterations.
-We theoretically prove that FL with communication errors can converge at the same rate as error-free communication as long as the bit error rate (BER) is properly constrained.
-An adaptive channel decoding scheme is then proposed to improve the energy efficiency of FL systems.
+In particular, the robustness is exploited to reduce the energy consumption of channel decoders at mobile clients by adaptively adjusting the number of decoding iterations.
+We theoretically prove that wireless FL with communication errors can converge at the same rate as the case with error-free communication as long as the bit error rate (BER) is properly constrained.
+An adaptive channel decoding scheme is then proposed to improve the energy efficiency of wireless FL systems.
 Experimental results demonstrate that the proposed method maintains the same learning accuracy while reducing the channel decoding energy consumption by 20% when compared to existing approaches.
@@
 As a result, uploading the distributed data, which may contain privacy-sensitive information, for centralized learning becomes prohibited.
-To address this issue, federated learning (FL) was proposed [3] to enable the training of deep learning models without sharing the data of users (e.g., IoT devices).
+To address this issue, federated learning (FL) was proposed [3] to enable the distributed training of deep learning models without sharing the data of users (e.g., IoT devices).
 However, to fully unlock the potential of FL, we need to tackle several critical challenges, such as data and device heterogeneity [4, 5, 6], as well as the communication bottleneck[7].
-Another issue for wireless FL is the energy consumption of mobile devices due to heavy computation and communication workloads for training and transferring the deep neural network (DNN) parameters[8].
-Many innovative approaches have been proposed to reduce the energy consumption of mobile devices in FL from different perspectives.
-For instance, reinforcement learning was utilized in [9] to identify a near-optimal clustering of participating devices to reduce the overall energy consumption of mobile devices.
-In [10], the trade-off between learning accuracy and users’ total energy consumption was optimized through computational and communication resource allocation.
-Besides system-level resource management, some works focused on the power consumption of specific hardware components such as transmitters, receivers, and artificial intelligence (AI) chips.
-Techniques such as model sparsification [11] and quantization [12, 13, 14, 15, 16, 17] have been proposed to minimize the communication overhead, thus reducing the energy consumption of the transmitters.
-The authors of [18] and [19] minimized the transmission energy while maintaining the learning performance.
-From computation perspective, model sparsification and quantization techniques are also beneficial for minimizing the processing energy consumption [20, 21, 22].
-Additionally, some works, such as [10], [23], and [24], considered the joint optimization of the uplink transmission and local training energy consumption.
-However, very few works have paid attention to the energy consumption of channel decoder.
-As will be shown in Section II-B, channel decoders consume a substantial portion of energy because of the iterative process to decode massive DNN parameters.
-To address the research gap, in this work, we first quantitatively analyze the energy consumption of different components in mobile clients and show that the channel decoder contributes significantly to the overall energy consumption.
-Building on this observation, an energy-efficient channel decoding strategy is designed for wireless FL systems, by leveraging the fact that FL is robust to model errors due to its statistical learning nature[25].
-Furthermore, we conduct a theoretical analysis on the relation between the bit error rate (BER) and the convergence rate of FL, and determine the maximum tolerable BER that can maintain the same convergence rate as the case with error-free communication.
-Finally, we propose a practical approach to adjust the decoding iterations in each communication round to achieve the maximum tolerable BER, while reducing the energy consumption.
+Another issue of wireless FL is the energy consumption of mobile clients due to heavy computation and communication workloads for training and transferring the deep neural network (DNN) parameters[8].
+Many innovative approaches have been proposed to reduce the energy consumption of mobile clients in FL from different perspectives.
+For instance, reinforcement learning was utilized in [9] to identify near-optimal client selection and resource allocation to reduce both the convergence time and energy consumption of mobile clients.
+In [10], the trade-off between learning accuracy and total energy consumption of users was optimized through computational and communication resource allocation.
+In [11], the transmission and computation energy was minimized by bandwidth allocation and power control.
+Besides system-level resource management, techniques such as model sparsification [12] and quantization [13, 14, 15, 16, 17, 18] have been proposed to minimize the communication overhead, thus reducing the energy consumption of the transmitters.
+Authors of [19] minimized the transmission energy consumption of mobile clients while maintaining the learning performance by optimizing the transmit power, user selection and uplink resource allocation.
+Authors of [20] adopted quantization and periodic aggregation to reduce the transmission volume and energy, which also provided a theoretical framework about the impact of quantization and periodic aggregation on the convergence rate of FL.
+Model sparsification and quantization techniques are also beneficial for minimizing the processing energy consumption [21, 8].
+Authors of [21] proposed to introduce sparsity at training time to reduce the on-device training energy.
+Authors of [8] reduced the local computation energy by quantization on both DNN weights and activations.
+However, most works on energy-efficient FL focused on the energy consumption of mobile transmitters while few of them paid attention to the energy consumption of channel decoders in mobile receivers.
+As will be shown in Section II-B, channel decoders consume a substantial proportion of energy because of the iterative processes to decode massive DNN parameters in each communication round of wireless FL.
+To address the research gap, in this work, we first quantitatively analyze the energy consumption of different components in a mobile unit and show that the channel decoder contributes significantly to the overall energy consumption.
+Building on this observation, an energy-efficient adaptive channel decoding scheme is proposed for wireless FL systems, by leveraging the fact that FL is robust to model errors due to its statistical learning nature[22].
+Furthermore, by extending the convergence analysis in[20] to the scenario of FL with downlink communication erros, we investigate the relationship between the bit error rate (BER) and the convergence rate of wireless FL, and determine the maximum tolerable BER that can maintain the same convergence rate as the case with error-free communication.
+Finally, we propose a practical approach to adjust the number of channel decoding iterations in each communication round to achieve the maximum tolerable BER, while reducing the energy consumption of mobile clients in wireless FL.
 I-B Contributions
-This paper investigates the impact of channel decoding on the energy consumption of wireless FL. The major contributions are summarized as follows:
-  • We theoretically analyze the relation between BER and the convergence rate of FL. The maximum tolerable BER to achieve the same convergence rate as error-free FL is determined.
-  • Based on the analysis, an adaptive low-density parity checking (LDPC) decoding mechanism is proposed, which controls the maximum number of decoding iterations to reduce the energy consumption of wireless FL while satisfying the BER requirement for a target convergence rate.
-  • Experimental results show that the proposed adaptive scheme effectively reduces the energy consumption by 20% when compared with the traditional scheme, while guaranteeing the same learning performance.
+This paper investigates the impact of channel decoding on the energy efficiency of wireless FL. Our major contributions are summarized as follows:
+  • We theoretically analyze the relationship between BER and the convergence rate of wireless FL.
+Specially, an upper bound of the average expected gradient norm of the global training loss with respect to BERs over communication rounds is derived by extending the analysis in [20], based on which, the maximum tolerable BER to achieve the same convergence rate as error-free FL is determined.
+  • Based on the theoretical analysis, an adaptive low-density parity checking (LDPC) decoding mechanism is proposed, which controls the maximum number of decoding iterations to satisfy the BER requirement in every communication round and maintain the target convergence rate as error-free wireless FL, while reducing the energy consumption of channel decoders in wireless FL.
+  • Experimental results on both IID and non-IID datasets show that the proposed adaptive scheme effectively reduces the energy consumption of channel decoders by 20% when compared with the traditional schemes which set a fixed maximum number of channel decoding iterations, while guaranteeing the same learning performance.
 I-C Organization
 The remainder of this paper is organized as follows.
-Section II introduces the system model and presents a comprehensive energy consumption analysis for mobile devices.
-Section III develops an energy-efficient LDPC decoding strategy for FL systems by exploiting the relationship between the convergence rate of FL and the achieved BER of LDPC decoding.
+Section II introduces the system model and presents a comprehensive energy consumption analysis for mobile units.
+Section III develops an energy-efficient adaptive LDPC decoding scheme for wireless FL systems by exploiting the relationship between the convergence rate and the achieved BER of LDPC decoding.
 Experimental results are presented in Section IV.
-Finally, concluding remarks are drawn in Section V.
+Finally, conclusions are drawn in Section V.
 II System Model and Preliminaries
 In this section, we first introduce the wireless FL system.
-An overview regarding the energy consumption of different components at a mobile unit is then presented, which highlights the significant contribution of channel decoding to the total energy consumption of wireless FL.
-For ease of reference, we list the key notations of the system model in Table I.
+An overview regarding the energy consumption of different components at a mobile unit (i.e., client) is then presented, which highlights the significant contribution of channel decoding to the total energy consumption of wireless FL.
+For ease of reference, we list the key notations in Table I.
 TABLE I: Key Notations
 Notation Definition
-$\mathbf{w}$ DNN model.
-$\Delta\mathbf{w}$ The local model update.
-$f$ The loss function of training.
-$f^{*}$ The minimum value of the training loss.
-${\nabla}f$ The gradient.
-$r$ The index of the communication round.
-$R$ The total communication rounds.
-$e$ The index of the local training step.
-$E$ The local training steps.
-$\sigma_{L}^{2}$ The variance of local gradients estimator.
-$\sigma_{G}^{2}$ The dissimilarity between local and global gradients.
-$k$ The index of the mobile device.
-$n$ The number of total mobile devices.
-$s$ The number of selected clients in every round
-$S_{r}$ The set of selected clients in the $r$-th round.
-$d$ The dimensions of the DNN model.
-$p_{k}$ The data size ratio of the $k$-th client.
-$T$ The total training iterations of one client.
-$L$ The L-smooth parameter.
-$b_{r}$ The bit error rate in $r$-th communication round.
-$\eta$ The learning rate.
-$M$ The biggest absolute weight value.
-$Q$ The LDPC maximum decoding iteration.
-(a) Wireless FL system.
-Multiple mobile clients are coordinated by one remote server through the noisy wireless channels.
-The communicated parameters are model updates in the uplink and model weights in the downlink.
-(b) System operations.
-Server: aggregation, quantization, LDPC encoding, client selection and broadcasting.
-Mobile devices: receiving model weights, LDPC decoding,local training and uplink transmission.
-Figure 1: Wireless FL in digital communication, system outlook and underlying operations.
+$\mathbf{w}$ DNN model
+$\Delta\mathbf{w}_{r}^{k}$ Local model update of client $k$ at the $r$-th communication round
+$f$ The loss function of training
+$f^{*}$ Minimum value of the global training loss
+$R$ The total communication rounds
+$E$ Number of local training steps in each communication round
+$\sigma_{L}^{2}$ Variance of local gradient estimator
+$\sigma_{G}^{2}$ Dissimilarity between local and global gradient
+$n$ Number of total clients
+$N$ Each DNN parameter is with $N$ bit representation
+$d$ Dimension of the DNN model
+$T$ Total training iterations, i.e., $T=RE$.
+$L$ The $L$-smooth parameter
+$b_{r}$ Bit error rate in the $r$-th communication round
+$\eta$ Learning rate
+$Q$ Maximum number of LDPC decoding iterations
+$M_{r}$ Range of the global model parameters in the $r$-th communication round
+$\hat{M}$ Maximum range of global model parameters for all communication rounds
+(a) Wireless FL system: Multiple mobile clients are coordinated by one remote server through noisy wireless channels, where model updates and model weights are communicated in the uplink and downlink, respectively.
+(b) System operations: The server performs model aggregation, digitalization, LDPC encoding, and model broadcasting.
+Mobile clients perform LDPC decoding, local training, and uploading the model updates.
+Figure 1: Wireless FL in digital communication network and its operations.
 II-A Wireless FL
-Fig. 1 depicts a wireless FL system, where model training of different clients is coordinated by the server. The FL problem can be formulated as follows [26, 27, 28]:
-$\min_{\mathbf{w}\in{\mathbb{R}^{d}}}f(\mathbf{w})\triangleq\sum_{k=1}^{n}p_{k}f_{k}(\mathbf{w}),$ (1)
-where $\mathbf{w}$ denotes the model parameters, $n$ represents the number of clients, $p_{k}$ is the ratio between the amount of training data samples of the $k$-th client and that of all the clients, and $f_{k}(\mathbf{w})$ is the local loss function of the $k$-th client.
-To train a shared global model, the server communicates with the mobile clients for multiple rounds through noisy wireless channels.
-In each communication round, the global model at the server is broadcast to a set of selected clients.
-Each of the selected clients then performs local training based on the received global model and transmits the model updates to the server for aggregation.
-Specifically, in the $r$-th communication round, the global model is quantized by a stochastic quantizer, encoded by a channel code, and broadcast to clients.
-The model parameters received by the $k$-th client is given as follows:
-$\mathbf{w}^{k}_{r,0}=\mathbf{w}_{r}^{\prime},$ (2)
-where $\mathbf{w}_{r}^{\prime}$ denotes the noisy version of $\mathbf{w}_{r}$.
-The impact of BER can be quantified by considering different error bit positions in the model parameters and the detailed analysis is given in Appendix B.
-A selected client updates the received global model by conducting $E$ steps of local mini-batch Stochastic Gradient Descent (SGD) with a learning rate $\eta$ as follows:
-$\mathbf{w}^{k}_{r,e+1}=\mathbf{w}^{k}_{r,e}-\eta{\nabla}f_{k}(\mathbf{w}^{k}_{r,e},\xi),e=0,\cdots,E-1,k\in\mathcal{S}_{r},$ (3)
-where ${\nabla}f_{k}(\mathbf{w}^{k}_{r,e},\xi)$ denotes the gradient computed from the local mini-batch datasets $\xi$ in the $k$-th client, and $\mathcal{S}_{r}$ denotes the set of selected clients in the $r$-th communication round.
+Fig. 1(a) depicts a wireless FL system, where model training of different clients is coordinated by a remote server. The FL problem can be formulated as follows [23, 24, 25]:
+$\min_{\mathbf{w}\in{\mathbb{R}^{d}}}f(\mathbf{w})\triangleq\frac{1}{n}\sum_{k=1}^{n}f_{k}(\mathbf{w}),$ (1)
+where $\mathbf{w}\in{\mathbb{R}^{d}}$ denotes the DNN model with $d$ parameters, $n$ represents the number of mobile clients, $f_{k}\left(\cdot\right)$ is the local loss function of the $k$-th client, and $f\left(\cdot\right)$ is the global loss function.
+Figure 2: Digitalization scheme for model parameters. This part gives an example, say [min, max] is digitalized by 3-bits and shows how each value inside is mapped to ”000”, ”001”, ”…”, ”111”.
+To train a shared global model, the server communicates with the clients for multiple rounds through noisy wireless channels.
+As shown in Fig.
+1(b), in each communication round, the global model at the server is broadcast to the clients.
+Each client then performs local training based on the received global model and transmits the model updates after local training to the server for aggregation.
+Specifically, in the $r$-th communication round, the global model is digitalized by $N$ bits, encoded by a channel code, and broadcast to clients.
+The model parameter digitalization scheme is illustrated in Fig.
+2, where the range of [min, max] is divided into $12^{N}-$ 1 intervals and the width of each interval is $\Delta=\frac{\text{max-min}}{2^{N}-1}$.
+In this paper, we assume $N$ is large enough so the continuous and digital representations of model parameters have negligible differences.
+The model parameters received by the $k$-th client in the $r$-th communication round is given as follows:
+$\mathbf{w}^{k}_{r,0}=\mathbf{\tilde{w}}_{r},\ k=1,\cdots,n,\ r=0,\cdots,R-1,$ (2)
+where $\mathbf{\tilde{w}}_{r}$ denotes the distorted version of $\mathbf{w}_{r}$ due to channel decoding errors.
+The impact of channel decoding errors can be quantified by considering the different error bit positions in the $N$-bit representations of model parameters, which will be analyzed in Section III-A.
+Each client updates the received global model by conducting $E$ steps of local mini-batch Stochastic Gradient Descent (SGD) with a learning rate $\eta$ as follows:
+$\begin{split}\mathbf{w}^{k}_{r,e+1}&=\mathbf{w}^{k}_{r,e}-\eta{\nabla}f_{k}(\mathbf{w}^{k}_{r,e};\xi_{r,e}^{k}),\\ &e=0,\cdots,E-1,\ k=1,\cdots,n,\end{split}$ (3)
+where ${\nabla}f_{k}(\mathbf{w}^{k}_{r,e};\xi_{r,e}^{k})$ denotes the gradient computed from a mini-batch datasets $\xi_{r,e}^{k}$ at the $k$-th client.
 By completing $E$ steps of local training, the local model update of the $k$-th client can be expressed as follows:
-${\Delta\mathbf{w}^{k}_{r}=\mathbf{w}^{k}_{r,E}-\mathbf{w}^{k}_{r,0}},$ (4)
-which is quantized, encoded, and uploaded to the server.
-For simplicity, we assume an orthogonal uplink channel is available for each selected client.
-After receiving these noisy local updates, denoted as $(\Delta\mathbf{w}^{k}_{r})^{\prime}$, the server performs aggregation to update the global model as follows:
-$\mathbf{w}_{r+1}=\mathbf{w}_{r}+\frac{1}{|\mathcal{S}_{r}|}\sum_{k\in{\mathcal{S}_{r}}}p_{k}(\Delta\mathbf{w}^{k}_{r})^{\prime}.$ (5)
-In this work, we focus on the downlink communication rather than the uplink for two reasons.
-First, our objective is to design channel decoding schemes to save the energy consumption of mobile devices.
-For the uplink, the channel decoding happens on the server side which falls out of the scope of this study.
-Second, the server has a continuous power supply and is able to conduct a sufficient number of decoding iterations to achieve satisfactory BER performance.
+$\begin{split}\Delta\mathbf{w}^{k}_{r}&=\mathbf{w}^{k}_{r,E}-\mathbf{w}^{k}_{r,0},\\ &r=0,\cdots,R-1,\ k=1,\cdots,n,\end{split}$ (4)
+which is also digitalized, encoded, and uploaded to the server.
+For simplicity, we assume an orthogonal uplink channel is available for each client.
+After receiving these local updates, denoted as $\Delta\mathbf{\tilde{w}}^{k}_{r}$, the server performs model aggregation to update the global model as follows:
+$\mathbf{w}_{r+1}=\mathbf{w}_{r}+\frac{1}{n}\sum_{k=1}^{n}\Delta\mathbf{\tilde{w}}^{k}_{r},\ r=0,\cdots,R-1.$ (5)
+In this work, we focus on the downlink communication rather than the uplink since our objective is to design novel channel decoding schemes to save the energy consumption of mobile clients.
+For the uplink, the channel decoding happens at the server, where a powerful channel decoder as well as a sufficient number of decoding iterations can always be applied to achieve satisfactory BER performance.
+As such, we consider $\Delta\mathbf{\tilde{w}}_{r}^{k}=\Delta\mathbf{w}_{r}^{k}$ in the following.
 II-B Energy Consumption Evaluation
@@
 Accordingly, the energy consumption of the LDPC decoder for decoding the AlexNet model parameters is 96.5 mJ.
-The detailed comparison is presented in Table II, which indicates that the energy consumption of LDPC decoding accounts for a large portion of the overall energy consumption in wireless FL.
+The detailed comparison is presented in Table II, which indicates that the energy consumption of LDPC decoding accounts for a large proportion of the overall energy consumption at clients in wireless FL.
 TABLE II: Energy Consumption of Main Components at a Mobile Unit in Wireless FL for Training an AlexNet
@@
 II-C LDPC Decoding
-Figure 2: Typical LDPC decoder architecture: the variable nodes and check nodes are connected, and the messages are transmitted between them iteratively.
 In this work, we use LDPC codes as an example to illustrate how adaptive channel decoding can help reduce the energy consumption of wireless FL.
-To that end, we first introduce a typical LDPC decoder architecture in Fig.
-2, which contains the variable nodes representing the encoded bits to be corrected and check nodes representing the parity check equation.
-The check nodes exchange information with variable nodes and perform computations iteratively until achieving the maximum decoding iteration [33].
-Such iterative processing leads to high energy consumption of LDPC decoding.
-To save decoding energy, early termination techniques are usually adopted.
-Specifically, if the parity-check matrix $\mathbf{H}$ and the codeword $\mathbf{x}$ satisfy $\mathbf{Hx}^{\rm{T}}=0$, the decoding process will stop even if the maximum decoding iteration has not been reached[34].
-The early termination scheme can save decoding iterations while preserving BER performance.
-The energy consumption of LDPC decoding depends on the target BER.
+In LDPC, the iterative decoding process continues until $Q$ decoding iterations, denoted as the maximum number of decoding iterations, are completed [30].
+Such iterative LDPC decoding processing leads to high energy consumption.
+To save the decoding energy, early termination techniques are usually adopted.
+Specifically, if the decoded codewords can fully pass the parity check, the decoding process will stop even if not all the $Q$ decoding iterations has been completed[31].
+The early termination scheme can save decoding iterations without compromising the BER performance.
+The energy consumption of LDPC decoding depends on the target BER and thus the pre-defined maximum number of decoding iterations.
 As a result, to minimize the energy consumption, it is essential to determine the maximum tolerable BER by FL systems, considering its unique statistical learning nature.
@@
 III-A Effect of BER on the Convergence Rate of Wireless FL
-With error-free communication, the convergence rate of FL with non-convex loss functions is proved as $\mathcal{O}(1/\sqrt{T})$ [35], where $T$ is the total number of training epochs.
+With error-free communication, the convergence rate of FL with non-convex loss functions is proved as $\mathcal{O}(1/\sqrt{T})$ [20], where $T$ is the total number of SGD iterations.
 In this subsection, we investigate how BER affects the convergence rate of FL and determine the maximum admissible BER for maintaining the convergence rate of $\mathcal{O}(1/\sqrt{T})$.
@@
 Assumption 1.
-($L$-Smoothness [12, 35]) The local loss function $f_{k}\left(\cdot\right)$ at the $k$-th client is $L$-smooth, i.e., $\forall\mathbf{w},\hat{\mathbf{w}}$, we have $||\nabla f_{k}(\mathbf{w})-\nabla f_{k}(\hat{\mathbf{w}})||\leq L||\mathbf{w}-\hat{\mathbf{w}}||$.
+($L$-Smoothness [13, 20]) The loss function $f\left(\cdot\right)$ is $L$-smooth, i.e., $\forall\mathbf{w},\mathbf{v}$, we have $||\nabla f(\mathbf{w})-\nabla f(\mathbf{v})||\leq L||\mathbf{w}-\mathbf{v}||$.
 Assumption 2.
-(Unbiased and Bounded-variance Local Gradient Estimator [43]) Let $\xi_{r}^{k}$ be the batch of randomly sampled data at the $k$-th client in the $r$-th communication round.
-The gradient estimator $\nabla f_{k}(\mathbf{w};\xi_{r}^{k})$ is unbiased, i.e., $\mathbb{E}[\nabla f_{k}(\mathbf{w};\xi_{r}^{k})]=\nabla f_{k}(\mathbf{w})$, and has a bounded variance, i.e., $\mathbb{E}[||\nabla f_{k}(\mathbf{w};\xi_{r}^{k})-\nabla f_{k}(\mathbf{w})||^{2}]\leq\sigma_{L}^{2}$.
+(Unbiased and Variance-bounded Local Gradient Estimator [39]) Let $\xi_{k}$ denote the batch of randomly sampled data at client $k$.
+The gradient estimator $\nabla f_{k}(\mathbf{w};\xi_{k})$ is unbiased, i.e., $\mathbb{E}_{\xi_{k}}[\nabla f_{k}(\mathbf{w};\xi_{k})]=\nabla f_{k}(\mathbf{w})$, and has bounded variance, i.e., $\mathbb{E}_{\xi_{k}}[||\nabla f_{k}(\mathbf{w};\xi_{k})-\nabla f_{k}(\mathbf{w})||^{2}]\leq\sigma_{L}^{2}$.
 Assumption 3.
-(Bounded Local and Global Gradient Dissimilarity [43]) The global dissimilarity between the local and global gradient is bounded, i.e., $||\nabla f_{k}(\mathbf{w})-\nabla f(\mathbf{w})||^{2}\leq\sigma_{G}^{2}$.
+(Bounded Local and Global Gradient Dissimilarity [39]) The dissimilarity between the local and global gradient is bounded, i.e., $||\nabla f_{k}(\mathbf{w})-\nabla f(\mathbf{w})||^{2}\leq\sigma_{G}^{2}$.
 Assumption 4.
-(Unbiased Quantization [12, 28, 35]) The random quantizer Q($\cdot$) is unbiased, i.e., $\mathbb{E}[Q(\mathbf{w})|\mathbf{w}]=\mathbf{w}$.
-With Assumptions 1-4, the convergence bound of wireless FL with a given BER is given by the following theorem.
+(Independent Bit Errors [40]) Each bit in the bitstream of the decoded model parameters encounters error independently.
+With the $N$-bit representation, each model parameter has $12^{N}-$ 1 possible distorted patterns.
+However, our empirical results in Fig.
+4 show that it is sufficient to focus on cases with at most one bit error of each model parameter.
+In particular, Fig.
+4 compares the actual model error with the analytical model error expression in (7) assuming at most 1-bit error, which match closely with various BER.
+Therefore, we impose the following assumption to simplify our analysis.
+Assumption 5.
+(At Most 1-bit Error of each Model Parameter) There is at most 1-bit error in the $N$-bit representation of each model parameter.
+Based on Assumptions 4 and 5, we are able to establish the relationship between the decoded model at a client and the broadcasted global model in the following lemma.
+Lemma 1.
+Denote $\tilde{\mathbf{w}}$ as the decoded model of $\mathbf{w}$. Based on Assumptions 4 and 5, the bias of each dimension in $\mathbf{w}$ is given as
+$\tilde{w}_{d}=w_{d}+\delta\left(w_{d}\right).$ (6)
+The mean square error caused by bit errors is given as
+$\mathbb{E}[||\mathbf{\tilde{w}}-\mathbf{w}||^{2}|\mathbf{w}]=\frac{d\left(4^{N}-1\right)}{3\left(2^{N}-1\right)^{2}}\cdot b(1-b)^{N-1}M^{2},$ (7)
+where $b$ is the BER.
+Proof.
+Please refer to Appendix V-A. ∎
+Figure 3: The difference between received noisy model and broadcasted perfect global model.
+(a) BER=$10^{-2}$
+(b) BER=$10^{-3}$
+(c) BER=$10^{-4}$
+Figure 4: Validation of the model error formula by the experiment of vanilla-CNN on Fashion-MNIST dataset.
+The experimental setting is in Section IV-A.
+In every communication round, we compute the actual model error between the noisy DNN model and perfect model, and compare with the value from the analytical function in (7).
+It is shown that the value from analytical function is very close to the actual value, for different BER levels and model range in different communication rounds.
+After understanding the effect of BER, we analyze the convergence property of wireless FL with communication errors.
+Lemma 2.
+With Assumptions 1-5 and the learning rate $\eta$ satisfies
+$1-\frac{n+L}{n}L\eta-2L^{2}\eta^{2}E(E-1)\geq 0,$ (8)
+the average expected gradient norm of $\{\nabla f(\mathbf{\bar{w}}_{r,e})\}$ is bounded as follows:
+$\begin{split}&\frac{1}{RE}\sum_{r=0}^{R-1}\sum_{e=0}^{E-1}\mathbb{E}\|\nabla f(\mathbf{\bar{w}}_{r,e})\|^{2}\leq\frac{2(\mathbb{E}f(\mathbf{w}_{0})-f^{*})}{\eta RE}\\ &+\frac{L(L+2)d\hat{M}^{2}(4^{N}-1)}{3(2^{N}-1)^{2}\eta RE}\sum_{r=0}^{R-1}b_{r}(1-b_{r})^{N-1}\\ &+\frac{d\hat{M}^{2}}{\eta RE}\sum_{r=0}^{R-1}(b_{r}(1-b_{r})^{N-1})^{2}\\ &+\frac{L(L+1)\eta(\sigma_{L}^{2}+\sigma_{G}^{2})}{n}+\frac{L^{2}\eta^{2}\sigma_{L}^{2}(n+1)(E-1)}{n},\end{split}$ (9)
+where $f^{*}$ is the minimum global training loss, $\hat{M}$ is the maximum range of DNN weights over all communication rounds, $b_{r}$ is the BER in the $r$-th communication round, and $\mathbf{w}_{0}$ is the initial global model.
+Proof.
+Please refer to Appendix V-B. ∎
+Remark 1.
+It is worthy noting that the first term on the RHS of (9) captures the distance between the initial and minimum global training loss.
+The second and third terms indicates how bit errors in each communication round affect learning convergence.
+The last two terms stem from non-IID data among clients and the random mini-batch sampling.
+We next simplify (9) by applying $\eta=\frac{1}{L\sqrt{RE}}=\frac{1}{L\sqrt{T}}$, which satisfies (8) with a sufficient large value of $R$.
 Theorem 1.
-When the learning rate is set as $\eta=\frac{1}{L\sqrt{RE}}=\frac{1}{L\sqrt{T}}$, the average expected squared norm of global stochastic gradient is bounded as follows:
-$\begin{split}\frac{1}{T}\sum_{r=1}^{R}\sum_{e=1}^{E}&\mathbb{E}\|\nabla f(\bar{\mathbf{w}}_{r,e})\|^{2}\\ \leq&\frac{C_{0}}{\sqrt{T}}+\frac{C_{1}\sum_{r=1}^{R}b_{r}}{\sqrt{T}}+\frac{C_{2}}{\sqrt{T}}+\frac{C_{3}}{\sqrt{T}},\end{split}$ (6)
-where $\bar{\mathbf{w}}_{r,e}\triangleq\frac{1}{n}\sum_{k=1}^{n}\mathbf{w}_{r,e}^{k}$, $C_{0}\triangleq 2L(f(\mathbf{w}_{0})-f^{*})$, $C_{1}\triangleq\frac{8L^{2}dM^{2}}{3}$, $C_{2}=\sigma_{L}^{2}(\frac{1}{n}+\frac{4(n-s)}{s(n-1)})$, $C_{3}\triangleq\frac{\sigma_{G}^{2}}{L}$, $f^{*}$ is the minimum global training loss, $b_{r}$ is the BER in the $r$-th communication round, and $\mathbf{w}_{0}$ is the initial global model.
+With Assumptions 1-5 and a sufficient large value of $R$ such that $\eta=\frac{1}{L\sqrt{RE}}$ satisfies (8), (9) can be simplified as follows:
+$\begin{split}\frac{1}{T}&\sum_{r=0}^{R-1}\sum_{e=0}^{E-1}\mathbb{E}\|\nabla f(\mathbf{\bar{w}}_{r,e})\|^{2}\\ \leq&\frac{C_{0}}{\sqrt{T}}+\frac{C_{1}\sum_{r=0}^{R-1}b_{r}(1-b_{r})^{N-1}}{\sqrt{T}}\\ &+\frac{C_{2}\sum_{r=0}^{R-1}(b_{r}(1-b_{r})^{N-1})^{2}}{\sqrt{T}}+\frac{C_{3}}{\sqrt{T}}+\frac{C_{4}}{T},\end{split}$ (10)
+where $C_{0}\triangleq 2L(\mathbb{E}f(\mathbf{w}_{0})-f^{*})$, $C_{1}\triangleq\frac{L^{2}(L+2)dM^{2}(4^{N}-1)}{3(2^{N}-1)^{2}}$, $C_{2}\triangleq LdM^{2}$, $C_{3}\triangleq\frac{(L+1)(\sigma_{L}^{2}+\sigma_{G}^{2})}{n}$, $C_{4}\triangleq\frac{L^{2}\sigma_{L}^{2}(n+1)(E-1)}{n}$.
+According to Theorem 1, in order to guarantee the convergence rate of $\mathcal{O}(1/\sqrt{T})$, it is crucial to control the BER in each communication round such that the RHS of (10) reduces to zero at a rate of $\mathcal{O}(1/\sqrt{T})$.
+Since the explicit relationship between $b_{r}$ and $r$ is difficult to obtain, we propose a simple BER regulation method in Corollary 1.
+As such, wireless FL systems can achieve the same convergence rate as those with error-free communication.
+Corollary 1.
+If the BER in each communication round is regulated as
+$b_{r}\propto\frac{1}{(r+1)^{2}},r=0,\cdots,R-1,$ (11)
+the RHS of (10) diminishes to zero at a rate of $\mathcal{O}(1/\sqrt{T})$, i.e., wireless FL converges to a local optimal solution of the global training loss funcation at a rate of $\mathcal{O}(1/\sqrt{T})$.
 Proof.
-The key idea is to first quantify the BER’s impact on the model error and then introduce the model error to the convergence analysis of vanilla learning process.
-The detailed proof is available in the Appendix.
-∎
-Remark 1.
-It is worthy noting that the first term on the right-hand side (RHS) of (6) captures the difference between the initial and minimum global training loss.
-The second term indicates how bit errors in each communication round affect learning convergence.
-The third term comes from the random mini-batch sampling and partial client participation, and the fourth term shows the impact of imbalanced data distribution.
-According to Theorem 1, in order to guarantee the convergence rate of $\mathcal{O}(1/\sqrt{T})$, it is crucial to control the BER in each communication round.
-Although the explicit relation between $b_{r}$ and $r$ is difficult to obtain, we know from the above discussion that, $b_{r}$ should be a decreasing function with $r$ as otherwise the second term on the RHS of (6) cannot be diminshed with $T$ increases.
-For ease of implementation, we propose a simple BER regulation with $b_{r}\propto\frac{1}{r^{2}}$, which is shown in the following corollary as a sufficient condition for a practical wireless FL system to achieve the same convergence rate as that with error-free communication.
-Corollary 1.
-If the BER in each communication round is regulated as follows:
-$b_{r}\propto\frac{1}{r^{2}},r=1,2,\cdots,R,$ (7)
-the RHS of (6) diminishes to zero at the rate of $\mathcal{O}(1/\sqrt{T})$, i.e., wireless FL converges to a local optimal solution of the global training loss funcation at a rate of $\mathcal{O}(1/\sqrt{T})$.
-Proof.
-To prove the RHS of (6) diminishes to zero at the rate of $\mathcal{O}(1/\sqrt{T})$, it is sufficient to verify
-$\lim_{T\rightarrow\infty}\frac{\sum_{r=1}^{R}b_{r}}{\sqrt{T}}=0$ (8)
-when $b_{r}\propto 1/r^{2}\triangleq\delta/r^{2}$. Since
-$\begin{split}0&\leq\frac{\sum_{r=1}^{R}b_{r}}{\sqrt{T}}=\delta\frac{\sum_{r=1}^{R}1/r^{2}}{\sqrt{T}}\\ &=\frac{\delta}{\sqrt{T}}\left(1+\sum_{r=2}^{R}\frac{1}{r^{2}}\right)\\ &<\frac{\delta}{\sqrt{T}}\left(1+\sum_{r=2}^{R}\frac{1}{r(r-1)}\right)=\frac{\delta}{\sqrt{T}}\left(2-\frac{1}{R}\right)\\ &=\frac{\delta}{\sqrt{T}}\left(2-\frac{E}{T}\right),\end{split}$ (9)
-we conclude (8) by letting $T\rightarrow\infty$ for both sides of (9) and appying the Sandwich Theorem. ∎
-Corollary 1 indicates that gradually decreasing the BER at a rate of $\mathcal{O}(1/r^{2})$ can guarantee the convergence rate of wireless FL at a rate of $\mathcal{O}(1/\sqrt{T})$.
-The proposed BER scaling rule achieves an improving communication quality with communication rounds.
+When $b_{r}\propto\frac{1}{(r+1)^{2}}\triangleq\frac{\vartheta}{(r+1)^{2}}$, to prove the RHS of (10) diminishes to zero at a rate of $\mathcal{O}(1/\sqrt{T})$, it is sufficient to verify
+$\lim_{T\rightarrow\infty}\frac{\sum_{r=0}^{R-1}b_{r}(1-b_{r})^{N-1}}{\sqrt{T}}=0$ (12)
+due to the fact that $0\leq\left(b_{r}(1-b_{r})^{N-1}\right)^{2}\leq b_{r}(1-b_{r})^{N-1}\leq 1$ for $0\leq b_{r}\leq 1$. Specifically, since
+$\begin{split}0&\leq\frac{\sum_{r=0}^{R-1}b_{r}(1-b_{r})^{N-1}}{\sqrt{T}}\leq\frac{\sum_{r=0}^{R-1}b_{r}}{\sqrt{T}}=\vartheta\frac{\sum_{r=0}^{R-1}1/(r+1)^{2}}{\sqrt{T}}\\ &=\frac{\vartheta}{\sqrt{T}}\left(1+\sum_{r=1}^{R-1}\frac{1}{(r+1)^{2}}\right)\\ &<\frac{\vartheta}{\sqrt{T}}\left(1+\sum_{r=1}^{R-1}\frac{1}{r(r+1)}\right)=\frac{\vartheta}{\sqrt{T}}\left(2-\frac{1}{R}\right)\\ &=\frac{\vartheta}{\sqrt{T}}\left(2-\frac{E}{T}\right),\end{split}$ (13)
+we conclude the proof by letting $T\rightarrow\infty$ for both sides of (13) and applying the Sandwich Theorem. ∎
+Corollary 1 indicates that gradually decreasing the BER at a rate of $\mathcal{O}(1/(r+1)^{2})$ can guarantee wireless FL converges at a rate of $\mathcal{O}(1/\sqrt{T})$.
+The proposed BER regulation rule achieves an improving communication quality with communication rounds.
 This is consistent with the solution developed in [36], which improves the communication quality by increasing the signal-to-noise ratio (SNR) with communication round.
@@
 2 Output: The converged global model $\mathbf{w}_{R}$
-3 for  Communication round $r=1,2,...,R$ do
-4       for Client index $k$ in $\{1,2,...,K\}$ do
-5             $b_{r}\propto(1/r^{2})$ with $b_{1},b_{R}$
+3 for  Communication round $r=0,\cdots,R-1$ do
+4       for Client index $k=1,\cdots,n$ do
+5             $b_{r}\propto(1/(r+1)^{2})$ with $b_{0},b_{R-1}$
 6             $Q=f^{-1}(b_{r})$
@@
 13             end for
-14            output: $\mathbf{w}_{r}^{\prime}$
-15             for local training step $e=1,...,E$ do
-16                  $\mathbf{w}^{k}_{r,0}=\mathbf{w}_{r}^{\prime}$
-17                   $\mathbf{w}^{k}_{r,e+1}=\mathbf{w}^{k}_{r,e}-\eta_{r}\nabla f(\mathbf{w}^{k}_{r,e})$
+14            output: $\mathbf{\tilde{w}}_{r}$
+15             for local training step $e=0,...,E-1$ do
+16                  $\mathbf{w}^{k}_{r,0}=\mathbf{\tilde{w}}_{r}$
+17                   $\mathbf{w}^{k}_{r,e+1}=\mathbf{w}^{k}_{r,e}-\eta\nabla f_{k}(\mathbf{w}^{k}_{r,e};\xi^{k}_{r,e})$
 18             end for
@@
 20       end for
-21      $\mathbf{w}_{r+1}=\mathbf{w}_{r}+\sum_{k=1}^{K}p_{k}\Delta\mathbf{w}_{r}^{k}$
+21      $\mathbf{w}_{r+1}=\mathbf{w}_{r}+\frac{1}{n}\sum^{n}_{k=1}\Delta\mathbf{w}_{r}^{k}$
 22 end for
@@
 In principle, by combining (11) and (16), we obtain the following relationship between the maximum number of LDPC decoding iterations and communication rounds:
-$Q_{\color[rgb]{0,0,0}r}=\frac{1}{\chi_{0}}\mathcal{O}\left(2\log\ r\right)+\frac{1}{\chi_{0}}\mathcal{O}\left(\frac{1}{\epsilon}\log\frac{1}{\epsilon}\right),$ (13)
+$Q_{r}=\frac{1}{\chi_{0}}\mathcal{O}\left(2\log\ r\right)+\frac{1}{\chi_{0}}\mathcal{O}\left(\frac{1}{\epsilon}\log\frac{1}{\epsilon}\right),$ (17)
 which provides a guideline to select the maximum number of LDPC decoding iterations in the $r$-th communication round.
@@
 However, (17) cannot be directly utilized because $\alpha$ and $\beta$ are code-specific and depend on the receive SNR as well as channel decoding algorithms.
-Thus, we propose a practical approach by a lookup table, which captures the relationship between BER and the maximum number of decoding iterations at different levels of receive SNR.
-Entries in the lookup table can be obtained through empirical measurements specific to the LDPC code used in the wireless FL system.
+Thus, we propose a practical approach by the mapping relationship which captures the relationship between BER and the maximum number of decoding iterations at different levels of receive SNR.
+Entries in the mapping relationship can be obtained through empirical measurements specific to the LDPC code used in the wireless FL system.
 Fig.
-3 shows the sample simulation results for a rate 1/2 LDPC code with a block length of 1008 bits and binary phase shift keying (BPSK) modulation over an additional white Gaussian noise (AWGN) channel with $\text{SNR}=1.5$ and 2.5 dB.
-Data in this figure can be stored in a three-dimensional lookup table [46].
-From the lookup table, we can get the corresponding maximum number of decoding iteration for a target BER for clients with different receive SNR.
+5 shows the exampled simulation results for a rate 1/2 LDPC code with a block length of 1008 bits and binary phase shift keying (BPSK) modulation over an additional white Gaussian noise (AWGN) channel with $\text{SNR}=1.5$ and 2.5 dB.
+Data in this figure can be stored in memory [43].
+From the mapping relationship, we can get the corresponding maximum number of decoding iteration for a target BER for clients with different receive SNR.
 Figure 5: The relation between BER and LDPC maximum decoding iterations under different SNRs.
@@
 The algorithm starts with an initial target BER, denoted as $b_{0}$, which is set to a relatively high value, i.e., $10^{-1}$ in our experiments.
-The target BER at the last communication round, denoted as $b_{R}$, is set as a small value, i.e., $10^{-4}$ in our experiments.
-The concrete expression of $b_{r}$ in our experiment is $b_{r}=\frac{0.0999R^{2}}{(R^{2}-1)r^{2}}+\frac{10^{-4}R^{2}-0.1}{R^{2}-1}$.
-For each communication round, the BER is scaled according to (7), and the corresponding maximum number of LDPC decoding iterations in that round is obtained from the lookup table.
+The target BER at the last communication round, denoted as $b_{R-1}$, is set as a small value, i.e., $10^{-4}$ in our experiments.
+The concrete expression of $b_{r}$ in our experiment is $b_{r}=\frac{0.0999R^{2}}{(R^{2}-1)(r+1)^{2}}+\frac{10^{-4}R^{2}-0.1}{R^{2}-1}$.
+For each communication round, the BER is scaled according to (11), and the corresponding maximum number of LDPC decoding iterations in that round is obtained from the mapping relationship.
 Remark 2.
-Note that the number of entries in the lookup table depends on the possible discrete BER and SNR values.
-Since this number is not large, the memory overhead for storing the lookup table is expected to be small.
-Additionally, reading data from the lookup table takes constant time and does not introduce additional LDPC decoding latency.
+Note that the number of entries in the mapping relationship depends on the possible discrete BER and SNR values.
+Since this number is not large, the memory overhead for storing the mapping relationship is expected to be small.
+Additionally, reading data from the memory takes constant time and does not introduce additional LDPC decoding latency.
 IV Experimental Results
@@
 First, we introduce the experimental settings.
-Then, the performance of the proposed algorithm, including test accuracy and energy consumption, is compared with a fixed-iteration LDPC decoding baseline.
-Figure 4: Fixed decoding schemes: On the mobile side, setting the LDPC maximum decoding iterations to 6, 12, 18, 24, and 52 can achieve a BER of $10^{-1}$, $10^{-2}$, $10^{-3}$, $10^{-4}$, and $10^{-5}$, respectively.
+Then, the performance of the proposed algorithm, including test accuracy and energy consumption, is compared with a baseline with a fixed number of LDPC decoding iterations.
+Figure 6: Fixed decoding schemes: On the client side, setting the LDPC maximum decoding iterations to 6, 12, 18, 24, and 52 can achieve a BER of around $10^{-1}$, $10^{-2}$, $10^{-3}$, $10^{-4}$, and $10^{-5}$, respectively.
 IV-A Experimental Settings
@@
 System Setups: We consider an FL system with 10 clients.
-Clients are assumed active in every training round and they communicate with the server using rate 1/2 min-sum LDPC code with a code length of 1008[47].
+All clients are assumed to be selected in every communication round and they communicate with the server using rate 1/2 min-sum LDPC code with a code length of 1008 bits[44].
 Besides, binary phase shift keying (BPSK) is adopted as the modulation scheme.
-In each training round, the number of local SGD steps is set as 5 with the learning rate and batch size as 0.01 and 64, respectively.
+In each communication round, the number of local SGD steps is set as 5 with the learning rate and batch size as 0.01 and 64, respectively.
+The receive SNR of all the wireless channels is assumed to be 2.5 dB.
 Datasets: We conduct experiments on three dataset, including MNIST[45], Fashion-MNIST[46], and CIFAR-10[47].
-The training datasets were distributed among mobile devices in an independent and identically distributed (IID) manner.
-We further consider one non-IID experiment on Fashion-MNIST where each client was assigned only two categories of training data, similar to the setup in [27].
-The test datasets were utilized for validation on the server side.
+The training datasets are distributed among clients in an IID manner.
+We further consider one experiment where the Fashion-MNIST data are non-IID among clients.
+In particular, each client was assigned with two categories of training data, similar to the setup in [24].
+The test datasets are used for validation at the server.
 DNN Models: For the MNIST dataset, the LeNet-300-100 model[48] is employed, which is a fully connected neural network consisting of two hidden layers and one softmax layer, with approximately 0.27 million parameters.
@@
 (f) Total decoding energy.
-Figure 7: Experiment for LeNet-300-100 on MNIST.
+Figure 9: Experiment for LeNet-300-100 on MNIST dataset.
 Compared with the conventional fixed schemes, our method that adaptively increases the maximum decoding iteration to 24 can achieve the same final accuracy of 98%, and reduce the decoding energy by 6.7mJ.
@@
 (f) Total decoding energy.
-Figure 8: Experiment for vanilla CNN on Fashion-MNIST.
+Figure 10: Experiment for vanilla CNN on IID Fashion-MNIST dataset.
 Compared with the conventional fixed schemes, our method that adaptively increases the maximum decoding iteration to 24 can achieve a similar final accuracy of 91%, and reduce the decoding energy by 14.1mJ.
-In this subsection, the experimental results for conventional schemes that use fixed decoding iterations are presented.
-As shown in Fig.
-4, the fixed decoding iterations lead to fixed BERs in the whole training process.
-Next, the impact of fixed BER on FL performance was investigated.
+In this subsection, experimental results for the baseline scheme are presented. As shown in Fig. 6, it leads to approximately constant BERs over the training process.
+Next, we investigate the impact of a constant BER on the performance of wireless FL.
 Fig.
-5 summarizes the test accuracy with different fixed number of maximum decoding iterations.
-It can be observed that for all four FL tasks, when BER is higher than $10^{-3}$, the learning performance was significantly impacted.
-Lowering the BER to $10^{-4}$ leads to improved accuracy and convergence towards perfect communication levels.
-However, further decreasing the BER to $10^{-5}$ will not yield any significant improvement in learning performance.
-IV-C Learning Performance of the Proposed Approach
-In this subsection, the proposed BER scaling algorithm is compared to the fixed BER scheme.
-The proposed approach uses a scaling BER of $\mathcal{O}(1/r^{2})$ by controlling the maximum decoding iteration in each communication round based on a lookup table that summarizes the relationship between the BER and decoding iterations.
+7 summarizes the test accuracy with different maximum numbers of decoding iterations.
+It can be observed that for all the four FL tasks, when the BER is higher than $10^{-3}$, the learning performance was significantly impaired.
+Lowering the BER to $10^{-4}$ leads to improved accuracy and faster convergence towards wireless FL with error-free communication.
+However, further decreasing the BER to $10^{-5}$ does not yield any significant improvement in test accuracy.
+IV-C Learning Performance of the Proposed Scheme
 (a) Max.iter.
@@
 (f) Total decoding energy.
-Figure 9: Experiment for 7-layer CNN on CIFAR-10.
+Figure 11: Experiment for 7-layer CNN on CIFAR-10 dataset.
 Compared with the conventional fixed schemes, our method that adaptively increases the maximum decoding iteration to 24 can achieve a similar final accuracy of 77.6%, and reduce the decoding energy by 190mJ.
@@
 (f) Total decoding energy.
-Figure 10: Experiment for vanilla CNN on non-IID Fashion-MNIST.
+Figure 12: Experiment for vanilla CNN on non-IID Fashion-MNIST dataset.
 Compared with the conventional fixed schemes, our method that adaptively increases the maximum decoding iteration to 24 can achieve a similar final accuracy of 54.1%, and reduce the decoding energy by 13.9mJ.
-Specifically, the learning process is assigned with an initial BER value of $10^{-1}$ and a target BER value of $10^{-4}$.
-Based on the BER scaling rule of $\mathcal{O}(1/r^{2})$, the desired BER at every communication round can be calculated.
-Then based on the lookup table, we can control the LDPC maximum decoding iteration for every communication round.
-In the experiments, the initial and final maximum decoding iterations are set to 6 and 24, respectively.
-The learning performance and energy consumption of the proposed adaptive scheme are compared to those of fixed schemes with fixed BER of $10^{-4}$.
-The corresponding fixed maximum decoding iterations is set to 24.
-To mitigate energy consumption, the LDPC decoder employs an early termination scheme, where the decoding process stops when the bitstream in a frame is deemed correct.
-Thus, only frames that cannot be correctly decoded require the maximum decoding iteration.
+In this subsection, the proposed adaptive LDPC decoding scheme is compared to the baseline scheme.
+Specifically, the learning process of the proposed scheme is assigned with an initial BER of $10^{-1}$ and a final BER value of $10^{-4}$.
+The desired BER at every communication round is calculated via (11) and the maximum number of LDPC decoding iterations is determined according to the mapping with BER as stated in Section III.
+In the experiments, the initial and final maximum numbers of LDPC decoding iterations are thus set to 6 and 24, respectively.
+For the baseline scheme, a maximum of 24 LDPC decoding iterations are used in each communication round, which corresponds to a BER of $10^{-4}$.
 Fig.
-6 shows an example for the early termination scheme by using the experiment of vanilla-CNN on Fashion-MNIST, illustrating the actual decoding iterations for each bitstream.
-It can observed that only a few bitstreams use up the maximum decoding iterations which is set to 52.
-The average decoding iterations is the average number of real consumed decoding iterations.
-Given the decoding energy is proportional to the decoding iterations, the decoding energy consumption can be reflected by the average number of decoding iterations used for the FL task.
+8(a) shows an example on the Fashion-MNIST (IID) datasets to illustrate the actual number of decoding iterations for each frame of data bits.
+It can observed that only a few frames exhaust all the decoding iterations which is set to a large value of 52.
+Note that the decoding energy consumption is proportional to the actual number of decoding iterations used for the FL tasks.
+From the distribution on the actual number of decoding iterations shown in Fig.
+8(b), it is observed that 18 decoding iterations are sufficient for most cases, implying the significant reduction of energy consumption.
 Fig.
-7 summarises the experimental results for LeNet-300-100 on MNIST.
+9 summarizes the experimental results for LeNet-300-100 on the MNIST (IID) dataset.
+Specifically, Fig.
+9(a) shows how the maximum number of LDPC decoding iterations varies with communication rounds in the proposed scheme.
+While the baseline scheme with a fixed number of maximum decoding iterations results in a constant BER, the proposed scheme achieves the desired BER by adopting a linearly increasing maximum number of decoding iterations.
 Fig.
-7(a) shows how the maximum number of LDPC decoding iterations varies with the FL communication rounds for the proposed approach compared with other fixed schemes.
-The traditional schemes adopt a fixed number of maximum decoding iterations, resulting in a fixed BER, while the proposed scheme adopts a roughly linearly increasing number of maximum decoding iterations to achieve the designed BER.
+9(b) shows the test accuracy achieved by different schemes.
+It is observed that setting the maximum number of decoding iterations to 6 (which corresponds to a BER of $\sim 10^{-1}$) severely affects the learning accuracy.
+Similarly, when the maximum number of decoding iterations is set to 12 (which corresponds to a BER of $\sim 10^{-2}$), the training process still fails to converge.
+By setting the maximum number of decoding iterations as 18 (which corresponds to a BER of $\sim 10^{-3}$), the training process converges, but with a noticeable loss in test accuracy compared to the case when the maximum number of decoding iterations is 24 (which corresponds to a BER of $\sim 10^{-4}$).
+However, further increasing the maximum number of decoding iterations to 52 does not lead to much accuracy improvement.
+The proposed scheme, which linearly increases the maximum number of decoding iterations from 6 to 24, achieves the same test accuracy as those with fixed 24 or 52 decoding iterations.
 Fig.
-7(b) shows the learning accuracy curves for different schemes.
-Setting the fixed maximum decoding iteration to 6, which corresponds to a BER of $10^{-1}$, severely affects the training accuracy.
-Similarly, when the maximum decoding iteration is set to 12, resulting in a BER of about $10^{-2}$, the training still fails to converge.
-With a maximum decoding iteration of 18, corresponding to a BER of about $10^{-3}$, the training can converge, but with a noticeable loss in accuracy.
-The training converges well to high learning accuracy when the maximum decoding iteration is increased to 24, resulting in a BER of $10^{-4}$.
-Further increasing the maximum decoding iteration to 52 does not lead to further accuracy improvement.
-The proposed scheme, which approximately increases the maximum decoding iteration from 6 to 24 in a linear manner, enables the training accuracy to improve rapidly and converge to the same level as the fixed scheme with 24/52 iterations.
+9(c) illustrates how the average number of decoding iterations over all DNN parameters varies during the training process.
+The proposed scheme can reduce the average number of decoding iterations and thus the overall energy consumption.
 Fig.
-7(c) illustrates how the average number of decoding iterations varies during training.
-The proposed scheme can reduce the average number of iterations and thus the overall energy consumption.
+9(d) depicts the BER of the proposed scheme over communication rounds, which scales as $\mathcal{O}(1/(r+1)^{2})$ from $10^{-1}$ to $10^{-4}$.
 Fig.
-7(d) depicts the BER of the proposed scheme, which scales as $\mathcal{O}(1/r^{2})$ from $10^{-1}$ to $10^{-4}$.
+9(e) shows the final test accuracy of different schemes at the 50-th communication round.
+It can be observed that the proposed scheme achieves the same accuracy as the baseline schemes that always maintain a very low BER of $10^{-4}$ and $10^{-5}$ throughout the training process.
 Fig.
-7(e) shows the final accuracy of different schemes.
-It can be observed that the proposed approach achieves the same accuracy as the fixed schemes that always maintain a very low BER of $10^{-4}$ and $10^{-5}$ throughout the training process.
-Fig.
-7(f) illustrates the decoding energy consumption on the mobile side for the different schemes.
+9(f) illustrates the decoding energy consumption on clients under different schemes.
 It is obtained by multiplying the total number of LDPC decoding iterations with the LDPC decoding energy efficiency reported in [29].
-Compared to the fixed schemes using a maximum decoding iteration of 24, the proposed scaling scheme can reduce the overall decoding energy by 19.4%.
-Figs.
-8, 9, and 10 present the experimental results for Fashion-MNIST, CIFAR-10, and non-IID Fashion-MNIST, respectively.
-The results are similar to those obtained by the MNIST experiment and show that the adaptive scheme can achieve similar or even slightly better accuracy compared to the fixed schemes.
-The observation of slightly better learning accuracy was also reported in [52, 53, 54], which suggests that adding noise to the model could enhance its generalization ability, and annealed noise could help achieve better learning accuracy.
-Overall, the proposed scheme reduces the LDPC decoding energy for Fashion-MNIST, CIFAR-10, and non-IID Fashion-MNIST by 18.3%, 20.1%, and 18.0%, respectively.
-Table III provides a summary of the experimental data.
-V Conclusion
-In this paper, we considered the energy consumption issue in federated learning (FL) over digital communication networks.
-The study recognized that channel decoding contributes to a large portion of the energy consumption for mobile FL, which has not been previously highlighted.
-It was shown that, due to the robustness of FL to communication errors, properly control the BER over different communication rounds can achieve the same convergence rate as error-free FL.
-Based on this analysis, we proposed an energy-efficient LDPC decoding iteration control algorithm.
-Experimental results demonstrated that the proposed method can reduce approximately 20% the total energy consumed by LDPC decoding, while maintaining the same or even a slightly higher learning accuracy compared to the scheme without decoding iteration control.
+Compared to the baseline scheme with a maximum of 24 decoding iterations, the proposed scheme reduces the overall decoding energy consumption by 19.4%.
 TABLE III: Experimental Results of Different Schemes. Our scheme can achieve similar accuracy while saving around 20% of the decoding energy compared to the best-fixed scheme.
@@
 non-IID Fashion 74.22 76.11 76.27 62.38
+Figs.
+10, 11, and 12 present the experimental results for Fashion-MNIST (IID), CIFAR-10 (IID), and Fashion-MNIST (non-IID) datasets, respectively.
+The results are similar to those on the MNIST (IID) dataset in Fig.
+9, showing that adapting the maximum number of LDPC decoding iterations can achieve similar or even slightly better accuracy compared to the baseline scheme.
+This observation was also reported in [49, 50, 51], which suggests that adding annealed noise to the DNN model could enhance its generalization ability.
+Moreover, the proposed scheme reduces the LDPC decoding energy consumption on the Fashion-MNIST (IID), CIFAR-10 (IID), and Fashion-MNIST (non-IID) datasets by 18.3%, 20.1%, and 18.0%, respectively.
+Table III provides a summary of the experimental data.
+V Conclusion
+In this paper, we considered the energy consumption issue in wireless federated learning (FL) over digital communication networks.
+Our study recognized that channel decoding contributes to a large portion of the energy consumption at mobile clients in wireless FL systems.
+To achieve energy-efficient wireless FL, we first show that by properly controlling the bit error rate (BER) over different communication rounds, wireless FL can achieve the same convergence rate as FL with error-free communication.
+Based on this analysis, we proposed an energy-efficient LDPC decoding scheme to regulate the maximum number of decoding iterations over communication rounds.
+Experimental results demonstrated that the proposed method can reduce approximately 20% of the total energy consumed by LDPC decoding, while maintaining the same or even a slightly higher learning accuracy compared to the scheme without adaptive decoding iteration control.
 Appendix
-V-A Proof of Theorem 1
-Proof Outline: To prove Theorem 1, we first derive four lemmas, namely Lemmas 1$\sim$4, to establish the relationship between the convergence rate and model error.
-Then Lemma 5 builds the connection between model error and BER.
-By combining these relations, Theorem 1 is proved, which presents the relationship between convergence rate and BER.
-For each communication round $r=1,...,R$ and local iteration $e=1,...,E$, we have
-$\mathbf{w}_{r+1}=\mathbf{w}_{r}+\frac{1}{s}\sum_{k\in{S_{r}}}(\mathbf{w}_{r,E}^{k}-(Q(\mathbf{w}_{r}))^{\prime}),$
-$\hat{\mathbf{w}}_{r+1}=\mathbf{w}_{r}+\frac{1}{n}\sum_{k\in{[n]}}(\mathbf{w}_{r,E}^{k}-(Q(\mathbf{w}_{r}))^{\prime}),$
-$\bar{\mathbf{w}}_{r,e}=\frac{1}{n}\sum_{k\in{[n]}}\mathbf{w}_{r,e}^{k},$ (14)
-where $\mathbf{w}_{r+1}$ represents the updated global model based on $s$ selected clients, $\hat{\mathbf{w}}_{r+1}$ denotes the updated global model based on all $n$ clients, and $\bar{\mathbf{w}}_{r,e}$ denotes the average model of $n$ local models.
-$(Q(\mathbf{w}_{r}))^{\prime}$ represents that the downlink model $\mathbf{w}_{r}$ will be impacted by the quantization operation $Q(\cdot)$ and channel noise $(\cdot)^{\prime}$ before being received by the clients.
-To simplify the representation, we will use $Q^{\prime}(\cdot)$ to replace $(Q(\cdot))^{\prime}$ in subsequent derivations.
-Lemma 1. Assuming that Assumptions 1 and 4 hold, for any communication round $r=1,...,R$, we have
-$\mathbb{E}f(\mathbf{w}_{r+1})$ $\leq\mathbb{E}f(\bar{\mathbf{w}}_{r,E})+\frac{L}{2}\mathbb{E}\|\hat{\mathbf{w}}_{r+1}-\bar{\mathbf{w}}_{r,E}\|^{2}$
-$+\frac{L}{2}\mathbb{E}\|\hat{\mathbf{w}}_{r+1}-\mathbf{w}_{r+1}\|^{2}.$ (15)
-This lemma provides a bound on the training loss of the updated global model.
-This bound can be derived from the $L$-smooth property of the loss function on the server side.
-The proof is given in Appendix-B.
-In the next three lemmas, we derive the upper bounds for the three terms on the RHS of (15).
-Lemma 2. Assuming that Assumptions 1 - 3 hold and considering the sequence of updates with step size $\eta$, the following inequality holds for the case of noisy downlink transmission:
-$\mathbb{E}f(\bar{\mathbf{w}}_{r,E})\leq\frac{L}{2}\mathbb{E}\|Q^{\prime}(\mathbf{w}_{r})-\mathbf{w}_{r}\|^{2}-\frac{\eta}{2}\sum_{e=1}^{E}\mathbb{E}\|\nabla f(\bar{\mathbf{w}}_{r,e})\|^{2}$
-$+\mathbb{E}f(\mathbf{w}_{r})-\frac{\eta}{2n}(1-L\eta)\sum_{e=1}^{E}\sum_{k\in[n]}\mathbb{E}\|\nabla f(\mathbf{w}_{r,e}^{k})\|^{2}$
-$+\frac{LE\eta^{2}\sigma_{L}^{2}}{2n}+\frac{E\eta^{2}\sigma_{G}^{2}}{2}.$ (16)
-The proof is given in Appendix-B.
-This lemma captures the process of noisy broadcasting and local updates.
-As shown in the RHS of Eq.(16), the first two terms arise from the noisy broadcasting, while the subsequent terms come from the local training process.
-Lemma 3. For $\hat{\mathbf{w}}_{r+1}$ and $\bar{\mathbf{w}}_{r,E}$ defined in (14), we have
-$\mathbb{E}\|\hat{\mathbf{w}}_{r+1}-\bar{\mathbf{w}}_{r,E}\|^{2}=\mathbb{E}\|\mathbf{w}_{r}-Q^{\prime}(\mathbf{w}_{r})\|^{2}.$ (17)
-This lemma indicates that the expectation of the global model error on the server side is the same as the downlink model error on client side. The proof can be found in Appendix-B.
-Lemma 4. For $\mathbf{w}_{r+1}$ and $\hat{\mathbf{w}}_{r+1}$ defined in (14), we have
-$\mathbb{E}\|\mathbf{w}_{r+1}-$ $\hat{\mathbf{w}}_{r+1}\|^{2}\leq\frac{4}{s(n-1)}(1-\frac{s}{n})\cdot$
-$\left[\eta^{2}\sigma_{L}^{2}En+\eta^{2}E\sum_{e=1}^{E}\sum_{k\in[n]}\|\nabla f(\mathbf{w}_{r,e}^{k})\|^{2}\right].$ (18)
-This lemma captures the effect of partial participation of clients. The proof is provided in Appendix-B.
-By substituting Lemmas 2-4 into Lemma 1, we derive the following recursive inequality for the updated model at the server:
-$\mathbb{E}f(\mathbf{w}_{r+1})\leq L\mathbb{E}\|\mathbf{w}_{r}-Q^{\prime}(\mathbf{w}_{r})\|^{2}-\frac{\eta}{2}\sum_{e=1}^{E}\mathbb{E}\|\nabla f(\bar{\mathbf{w}}_{r,e})\|^{2}$
-$-\frac{\eta}{2n}\left\{1-L\eta-\frac{4LE(n-s)}{s(n-1)}\right\}\cdot\sum_{e=1}^{E}\sum_{k\in[n]}\mathbb{E}\|\nabla f(\mathbf{w}_{r,e}^{k})\|^{2}$
-$+\mathbb{E}f(\mathbf{w}_{r})+\frac{\eta E\sigma_{G}^{2}}{2}+L\eta^{2}\sigma_{L}^{2}E(\frac{1}{2n}+\frac{2(n-s)}{s(n-1)}).$ (19)
-When $\eta$ satisfies the condition
-$1-L\eta-\frac{4LE(n-s)}{s(n-1)}\geq 0,$ (20)
-we have
-$\mathbb{E}f(\mathbf{w}_{r+1})\leq L\mathbb{E}\|\mathbf{w}_{r}-Q^{\prime}(\mathbf{w}_{r})\|^{2}-\frac{\eta}{2}\sum_{e=1}^{E}\mathbb{E}\|\nabla f(\bar{\mathbf{w}}_{r,e})\|^{2}$
-$+\mathbb{E}f(\mathbf{w}_{r})+\frac{E\eta^{2}\sigma_{G}^{2}}{2}+L\eta^{2}\sigma_{L}^{2}E(\frac{1}{2n}+\frac{2(n-s)}{s(n-1)}).$ (21)
-Summing (21) over $r=1,...,R$ and rearranging the terms, we obtain the following expression:
-$\frac{\eta}{2}\sum_{r=1}^{R}\sum_{e=1}^{E}\mathbb{E}\|\nabla f(\bar{\mathbf{w}}_{r,e})\|^{2}\leq L\sum_{r=1}^{R}\mathbb{E}\|\mathbf{w}_{r}-Q^{\prime}(\mathbf{w}_{r})\|^{2}$
-$+f(\mathbf{w}_{0})-f^{*}+\frac{RE\eta^{2}\sigma_{G}^{2}}{2}+RL\eta^{2}\sigma_{L}^{2}E(\frac{1}{2n}+\frac{2(n-s)}{s(n-1)}).$ (22)
-Multiplying $\frac{2}{RE\eta}$ on both sides of (22), we get
-$\frac{1}{RE}\sum_{r=1}^{R}\sum_{e=1}^{E}\mathbb{E}\|\nabla f(\bar{\mathbf{w}}_{r,e})\|^{2}\leq\frac{2L}{RE\eta}\sum_{r=1}^{R}\mathbb{E}\|\mathbf{w}_{r}-Q^{\prime}(\mathbf{w}_{r})\|^{2}$
-$+\frac{2(f(\mathbf{w}_{0})-f^{*})}{RE\eta}+\eta\sigma_{G}^{2}+L\eta\sigma_{L}^{2}(\frac{1}{n}+\frac{4(n-s)}{s(n-1)}).$ (23)
-Since a high quantization bit-length is assumed to be used in this work with negligible quantization error, we have
-$\mathbb{E}\|\mathbf{w}_{r}-Q^{\prime}(\mathbf{w}_{r})\|^{2}$ $=\mathbb{E}\|Q^{\prime}(\mathbf{w}_{r})-Q(\mathbf{w}_{r})+Q(\mathbf{w}_{r})-\mathbf{w}_{r}\|^{2}$
-$=\mathbb{E}\|(\mathbf{w}_{r})^{\prime}-\mathbf{w}_{r}\|^{2}.$ (24)
-Combining (23) and (24), we obtain
-$\frac{1}{RE}\sum_{r=1}^{R}\sum_{e=1}^{E}\mathbb{E}\|\nabla f(\bar{\mathbf{w}}_{r,e})\|^{2}\leq\frac{2L}{RE\eta}\sum_{r=1}^{R}\mathbb{E}\|(\mathbf{w}_{r})^{\prime}-\mathbf{w}_{r}\|^{2}$
-$+\frac{2(f(\mathbf{w}_{0})-f^{*})}{RE\eta}+\eta\sigma_{G}^{2}+L\eta\sigma_{L}^{2}(\frac{1}{n}+\frac{4(n-s)}{s(n-1)}).$ (25)
-Lemma 5. For the perfect global model $\mathbf{w}$ and the received noisy model $(\mathbf{w})^{\prime}$, the model error can be bounded by
-$\mathbb{E}||(\mathbf{w})^{\prime}-\mathbf{w}||^{2}\leq\frac{4d(4^{N}-1)}{3(2^{N}-1)^{2}}\cdot b\cdot M^{2},$ (26)
-where $M$ denotes the maximum absolute value among all weight elements, i.e.
-$,M=max(|w_{1}|,|w_{2}|,$$,|w_{d}|)$.
-Lemma 5 gives the relationship between model variance and BER.
-The proof can be found in Appendix-B.
-By substituting Eq.(26) into Eq.(25), we establish the relationship between the FL convergence rate and BER:
-$\frac{1}{RE}\sum_{r=1}^{R}\sum_{e=1}^{E}\mathbb{E}\|\nabla f(\bar{\mathbf{w}}_{r,e})\|^{2}\leq\frac{8LdM^{2}}{3RE\eta}\sum_{r=1}^{R}\frac{b_{r}(4^{N}-1)}{(2^{N}-1)^{2}}$
-$+\frac{2(f(\mathbf{w}_{0})-f^{*})}{RE\eta}+\eta\sigma_{G}^{2}+L\eta\sigma_{L}^{2}(\frac{1}{n}+\frac{4(n-s)}{s(n-1)}).$ (27)
-When the quantization bit-length is high, we have $\frac{4^{N}-1}{(2^{N}-1)^{2}}\approx 1$, and Eq.(27) can be simplified as follows:
-$\frac{1}{RE}\sum_{r=1}^{R}\sum_{e=1}^{E}\mathbb{E}\|\nabla f(\bar{\mathbf{w}}_{r,e})\|^{2}\leq\frac{2(f(\mathbf{w}_{0})-f^{*})}{RE\eta}$
-$+\frac{8LdM^{2}}{3RE\eta}\sum_{r=1}^{R}b_{r}+L\eta\sigma_{L}^{2}(\frac{1}{n}+\frac{4(n-s)}{s(n-1)})+\eta\sigma_{G}^{2},$ (28)
-which completes the proof of Theorem 1.
-V-B Proof of Lemmas
-1) Proof of Lemma 1
-Quantization is unbiased according to Assumption 3, i.e., $\mathbb{E}_{Q}{(\mathbf{w})}=\mathbf{w}$.
-It is proved in Eq.(44) that the impact of the BER is also unbiased.
-So we have $\mathbb{E}_{Q^{\prime}}{Q^{\prime}(\mathbf{w})}=\mathbf{w}$.
-Based on the above unbiased properties and the notations defined in Eq.(14), we have
-$\mathbb{E}_{Q^{\prime}}\hat{\mathbf{w}}_{r+1}$ $=\mathbf{w}_{r}+\frac{1}{n}\sum_{k\in[n]}(\mathbf{w}_{r,E}^{k}-\mathbb{E}_{Q^{\prime}}Q^{\prime}(\mathbf{w}_{r}))$
-$=\mathbf{w}_{r}+\frac{1}{n}\sum_{k\in[n]}(\mathbf{w}_{r,E}^{k}-\mathbf{w}_{r})$
-$=\frac{1}{n}\sum_{k\in[n]}\mathbf{w}_{r,E}^{k}$
-$=\bar{\mathbf{w}}_{r,E}.$ (29)
-We can also get
-$\mathbb{E}_{S_{r}}\mathbf{w}_{r+1}$ $=\mathbf{w}_{r}+\frac{1}{n}\sum_{k\in[n]}(\mathbf{w}_{r,E}^{k}-Q^{\prime}(\mathbf{w}_{r}))=\hat{\mathbf{w}}_{r+1},$ (30)
-whose proof is similar to Section 7.1 in [35].
-Since the loss function $f$ is $L$-smooth, for any variables $\mathbf{x,y}$, we have
-$f(\mathbf{x})\leq f(\mathbf{y})+\left\langle\nabla f(\mathbf{y}),\mathbf{x}-\mathbf{y}\right\rangle+\frac{L}{2}||\mathbf{x}-\mathbf{y}||^{2}.$ (31)
-Taking $\mathbf{x}=\hat{\mathbf{w}}_{r+1}$ and $\mathbf{y}=\mathbf{\bar{w}}_{r,E}$ in (31), we can get
-$f(\hat{\mathbf{w}}_{r+1})\leq$ $f(\mathbf{\bar{w}}_{r,E})+\left\langle\nabla f(\mathbf{\bar{w}}_{r,E}),\hat{\mathbf{w}}_{r+1}-\mathbf{\bar{w}}_{r,E}\right\rangle$
-$+\frac{L}{2}||\hat{\mathbf{w}}_{r+1}-\mathbf{\bar{w}}_{r,E}||^{2}.$ (32)
-We take the expectation on both sides of (32) and, since $\mathbb{E}_{Q^{\prime}}\hat{\mathbf{w}}_{r+1}=\bar{\mathbf{w}}_{r,E}$ (see (29)), we have
-$\mathbb{E}f(\hat{\mathbf{w}}_{r+1})\leq\mathbb{E}f(\mathbf{\bar{w}}_{r,E})+\frac{L}{2}\mathbb{E}||\hat{\mathbf{w}}_{r+1}-\mathbf{\bar{w}}_{r,E}||^{2}.$ (33)
-Taking $\mathbf{x}=\mathbf{w}_{r+1}$ and $\mathbf{y}=\hat{\mathbf{w}}_{r+1}$ in (31), and following the same steps as in (32) and (33), we have
-$\mathbb{E}f(\mathbf{w}_{r+1})\leq\mathbb{E}f(\hat{\mathbf{w}}_{r+1})+\frac{L}{2}\mathbb{E}||\hat{\mathbf{w}}_{r+1}-\mathbf{w}_{r+1}||^{2}.$ (34)
-By combining (33) and (34), we can get Lemma 1.
-2) Proof of Lemma 2
-Based on the proof in Section 8.3 of [35], and considering the non-iid case as in Assumption 3, if the downlink communication is perfect, we can obtain the upper bound of FL convergence as
-$\mathbb{E}$ $f(\bar{\mathbf{w}}_{r,E})\leq\mathbb{E}f(\mathbf{w}_{r})-\frac{\eta}{2}\sum_{e=1}^{E}\mathbb{E}\|\nabla f(\bar{\mathbf{w}}_{r,e})\|^{2}$
-$-\eta(\frac{1}{2n}-\frac{L\eta}{2n}-\frac{L^{2}\eta^{2}E(E-1)}{n})\sum_{e=1}^{E}\sum_{k\in[n]}\mathbb{E}\|\nabla f(\mathbf{w}_{r,e}^{k})\|^{2}$
-$+\frac{LE\eta^{2}\sigma_{L}^{2}}{2n}+\frac{L^{2}\eta^{3}\sigma_{L}^{2}(n+1)E(E-1)}{2n}.$ (35)
-For noisy downlink communication, the received model will have quantization and BER noise. As a result, (35) can be rewritten as
-$\mathbb{E}$ $f(\bar{\mathbf{w}}_{r,E})\leq\mathbb{E}f(Q^{\prime}(\mathbf{w}_{r}))-\frac{\eta}{2}\sum_{e=1}^{E}\mathbb{E}\|\nabla f(\bar{\mathbf{w}}_{r,e})\|^{2}$
-$-\eta(\frac{1}{2n}-\frac{L\eta}{2n}-\frac{L^{2}\eta^{2}E(E-1)}{n})\sum_{e=1}^{E}\sum_{k\in[n]}\mathbb{E}\|\nabla f(\mathbf{w}_{r,e}^{k})\|^{2}$
-$+\frac{LE\eta^{2}\sigma_{L}^{2}}{2n}+\frac{E\eta^{2}\sigma_{G}^{2}}{2}.$ (36)
-Since $f$ is $L$-smooth, according to (31), we have
-$f(Q^{\prime}(\mathbf{w}_{r}))\leq$ $f(\mathbf{w}_{r})+\left\langle\nabla f(\mathbf{w}_{r}),Q^{\prime}(\mathbf{w}_{r})-\mathbf{w}_{r}\right\rangle$
-$+\frac{L}{2}||Q^{\prime}(\mathbf{w}_{r})-\mathbf{w}_{r}||^{2}.$ (37)
-Since the quantization and BER are both unbiased, i.e., $\mathbb{E}_{Q^{\prime}}{Q^{\prime}(\mathbf{w})}=\mathbb{E}_{Q}{Q(\mathbf{w})}=\mathbf{w}$, we take the expectation on both sides of (37) to obtain
-$\mathbb{E}f(Q^{\prime}(\mathbf{w}_{r}))\leq\mathbb{E}f(\mathbf{w}_{r})+\frac{L}{2}\mathbb{E}||Q^{\prime}(\mathbf{w}_{r})-\mathbf{w}_{r}||^{2}.$ (38)
-Combining (36) and (38), we prove Lemma 2.
-3) Proof of Lemma 3
-According to the notations defined in Eq.(14), we can get the expanded representation of $\hat{\mathbf{w}}_{r+1}$ and $\bar{\mathbf{w}}_{r,E}$ and derive
-$\mathbb{E}\|\hat{\mathbf{w}}_{r+1}-\bar{\mathbf{w}}_{r,E}\|^{2}$
-$=\mathbb{E}\|\mathbf{w}_{r}+\frac{1}{n}\sum_{k\in[n]}(\mathbf{w}_{r,E}^{k}-Q^{\prime}(\mathbf{w}_{r}))-\frac{1}{n}\sum_{k\in[n]}\mathbf{w}_{r,E}^{k}\|^{2}$
-$=\mathbb{E}\|\mathbf{w}_{r}+\frac{1}{n}\sum_{k\in[n]}\mathbf{w}_{r,E}^{k}-\frac{1}{n}\sum_{k\in[n]}Q^{\prime}(\mathbf{w}_{r})-\frac{1}{n}\sum_{k\in[n]}\mathbf{w}_{r,E}^{k}\|^{2}$
-$=\mathbb{E}\|\mathbf{w}_{r}-Q^{\prime}(\mathbf{w}_{r})\|^{2},$ (39)
-which concludes the proof of Lemma 3.
-4) Proof of Lemma 4
-Similar to Section 8.5 in [35] and considering the noise in the downlink global model, we have
-$\mathbb{E}\|\mathbf{w}_{r+1}-\hat{\mathbf{w}}_{r+1}\|^{2}$ $\leq\frac{4}{s(n-1)}(1-\frac{s}{n})\cdot$
-$\sum_{k\in[n]}\mathbb{E}\|\mathbf{w}^{k}_{r,E}-Q^{\prime}(\mathbf{w}_{r})\|^{2}.$ (40)
-Based on $E$ steps of local training with learning rate $\eta$, we have
-$\mathbb{E}\|\mathbf{w}^{k}_{r,E}-Q^{\prime}(\mathbf{w}_{r})\|^{2}\leq\eta^{2}E(\sigma_{L}^{2}+\sum_{e=1}^{E}\|\nabla f(\mathbf{w}_{r,e}^{k})\|^{2}).$ (41)
-Combining (40) and (41), we can conclude Lemma 4.
-5) Proof of Lemma 5
-Due to communication errors, the received DNN model for local training will contain random errors and differ from the global model, as shown in Fig.
-11.
-In the following, we analyze the effect of the BER on the model error.
-Figure 11: The difference between received noisy model and broadcasted perfect global model.
-(a) BER=1E-2
-(b) BER=1E-3
-(c) BER=1E-4
-Figure 12: Validation of the model error formula.
-For a quantization scheme using $N$ bits to represent the weight $w$, the probability of having $x$ bit errors is given by
-$P(X=x)=C_{N}^{x}b^{x}(1-b)^{N-x}.$ (42)
-Since $b$ is much smaller than 1, the following relations can be established:
-$P(X=0)\gg P(X=1)\gg P(X=2)\gg...\gg P(X=8).$ (43)
-Therefore, the focus is on the case with at most one error bit.
-The bit error can occur at various positions, ranging from the Least Significant Bit (LSB) to the Most Significant Bit (MSB), leading to different levels of error.
-For instance, for unsigned quantization, if the erroneous bit appears in the MSB and flips from 0 to 1, the error would be $+2^{N-1}\Delta$, where $\Delta=\frac{max(\mathbf{w})-min(\mathbf{w})}{2^{N}-1}=\frac{range(\mathbf{w})}{2^{N}-1}$, denoting the bin value for a given quantization.
-Alternatively, if the MSB flips from 1 to 0, the quantified error will be $-2^{N-1}\Delta$.
-The results for the case that the error occurs at the LSB will be $+2^{0}\Delta$ and $-2^{0}\Delta$.
-Additionally, it is assumed that the numbers of 0 and 1 in the bitstreams are similar.
-Based on this analysis, we can determine all the cases of bit flipping, and calculate the resulting values of the noisy weight with the corresponding probabilities, as shown in Table IV.
-TABLE IV: Probability mass function of $(w)^{\prime}$.
-$(w)^{\prime}_{i}$ $w+2^{0}\Delta$ $w-2^{0}\Delta$ $w+2^{1}\Delta$ $w-2^{1}\Delta$
-$P_{i}$ b/2 b/2 b/2 b/2
-$(w)^{\prime}_{i}$ $w+2^{N-1}\Delta$ $w-2^{N-1}\Delta$ $w$
-[…]
-Based on Table IV, the mean and variance of the received noisy weight $(w)^{\prime}$ can be calculated as follows:
-$\mathbb{E}((w)^{\prime})=\sum_{i}((w)^{\prime}_{i}\cdot P_{i})=w,$ (44)
-$\mathbb{E}||(w)^{\prime}-w||^{2}$ $=\sum_{i}[((w)^{\prime}_{i}-w)^{2}\cdot P_{i}]$
-$=\frac{4^{N}-1}{3(2^{N}-1)^{2}}\cdot b\cdot range(\mathbf{w})^{2}.$ (45)
-The above equations indicate that, for a weight parameter in the DNN model, the mean value of the noisy parameter is equal to the broadcasted perfect parameter.
-Furthermore, the variance between the two parameters is a function of the number of quantization bits, the BER level, and the range of the model parameters.
-For a $d-$dimensional model $\mathbf{w}$, i.e., $\mathbf{w}=[w_{1},w_{2},...,w_{d}]$, the following equations can be derived:
-$\mathbb{E}||(\mathbf{w})^{\prime}-\mathbf{w}||^{2}$ $=\sum_{j=1}^{d}\mathbb{E}||(w_{j})^{\prime}-w_{j}||^{2}$
-$=\frac{(4^{N}-1)d}{3(2^{N}-1)^{2}}\cdot b\cdot range(\mathbf{w})^{2}.$ (46)
-To verify the theoretical results of Eq.(46), simulations were conducted using an 8-bit vanilla-CNN on the Fashion-MNIST dataset.
-The experimental settings are described in Section IV-A.
-Three different BER conditions were simulated, and the model errors obtained from the simulation are plotted in Fig.
-12.
-The expectation on the left-hand side (LHS) of Eq.(46) corresponds to different cases of errors.
-The model errors predicted using the proposed formula on the RHS of Eq.(46) are also plotted in the same figure.
-It can be seen that, for different BER conditions and different communication rounds, the results obtained from the theoretical formula closely match those obtained from simulation, confirming the accuracy of the derived Eq.(46).
-If we denote $M$ as the maximum absolute value of all weight elements, i.e.
-$,M=max(|w_{1}|,|w_{2}|,$$,|w_{d}|)$, we have $range(\mathbf{w})\leq 2M$.
-The upper bound of Eq.(46) can be obtained as follows:
-$\mathbb{E}||(\mathbf{w})^{\prime}-\mathbf{w}||^{2}\leq\frac{4d(4^{N}-1)}{3(2^{N}-1)^{2}}\cdot b\cdot M^{2},$ (47)
-which concludes the proof of Lemma 5.
+V-A Proof of Lemma 1
+We express a decoded model $\tilde{\mathbf{w}}$ in terms of $\mathbf{w}$ as follows:
+$\tilde{\mathbf{w}}=\mathbf{w}+\bm{\delta}\left(\mathbf{w}\right),$ (18)
+where $\bm{\delta}\left(\mathbf{w}\right)$ is a $d$-dimensional function of $\mathbf{w}$ denoting the model distortion caused by bit errors.
+Under Assumption 4, each entry of $\bm{\delta}\left(\mathbf{w}\right)$ depends only on $w_{d}$, i.e., the $d$-th dimension of $\mathbf{w}$.
+Denote $\delta\left(w_{d}\right)$ as the $d$-th dimension of $\bm{\delta}\left(\mathbf{w}\right)$.
+Therefore, the $d$-th dimension of $\tilde{\mathbf{w}}$ is given as
+$\tilde{w}_{d}=w_{d}+\delta\left(w_{d}\right).$ (19)
+Under Assumptions 4 and 5, the 1-bit error may occur at different bits, leading to different levels of model distortion.
+Let $w_{d}^{(i)}\in\{0,1\}$ be the $i$-th digit ($i=0$ and $N-1$ respectively denote the least and most significant bits) in the $N$-bit representation of model parameter $w_{d}$.
+Hence,
+$\mathbb{E}\left[\tilde{w}_{d}|w_{d}\right]=w_{d}+b\left(1-b\right)^{N-1}\Delta\sum_{i=0}^{N-1}\left(1-2w_{d}^{(i)}\right)2^{i},$ (20)
+The expected bias of $w_{d}$ is bounded as follows:
+$\begin{split}|\mathbb{E}\left[\tilde{w}_{d}|w_{d}\right]-w_{d}|&\leq b\left(1-b\right)^{N-1}\left(2^{N}-1\right)\Delta\\ &=b\left(1-b\right)^{N-1}M.\end{split}$ (21)
+The mean square error of $w_{d}$ is given as follows:
+$\begin{split}\mathbb{E}\left[||\tilde{w}_{d}-w_{d}||^{2}|w_{d}\right]&=b\left(1-b\right)^{N-1}\sum_{i=0}^{N-1}\left(\Delta 2^{i}\right)^{2}\\ &=b\left(1-b\right)^{N-1}\frac{4^{N}-1}{3(2^{N}-1)^{2}}M^{2}.\end{split}$ (22)
+Therefore,
+$\begin{split}\mathbb{E}\left[||\mathbf{\tilde{w}}-\mathbf{w}||^{2}|\mathbf{w}\right]&=\sum_{j=1}^{d}\mathbb{E}\left[||\tilde{w}_{j}-w_{j}||^{2}|\mathbf{w}\right]\\ &=\frac{d(4^{N}-1)}{3(2^{N}-1)^{2}}\cdot b\left(1-b\right)^{N}M^{2}.\end{split}$ (23)
+V-B Proof of Lemma 2
+We define
+$\mathbf{w}_{r+1}$ $=\mathbf{w}_{r}+\frac{1}{n}\sum_{k=1}^{n}(\mathbf{w}_{r,E}^{k}-\mathbf{\tilde{w}}_{r}),$ (24)
+$\mathbf{\bar{w}}_{r,e}$ $=\frac{1}{n}\sum_{k=1}^{n}\mathbf{w}_{r,e}^{k},$ (25)
+for each communication round $r=0,...,R-1$ and each local iteration $e=0,\cdots,E-1$, where $\mathbf{w}_{r+1}$ denotes the updated global model and $\mathbf{\bar{w}}_{r,e}$ is the average of client models after the $e$-th local iteration.
+To prove the Lemma 2, we first present Lemmas 26$\sim$5.
+Lemma 3.
+For $\mathbf{\bar{w}}_{r,E}$ and $\mathbf{w}_{r+1}$ defined in (24) and (25), we have
+$\mathbf{w}_{r+1}-\mathbf{\bar{w}}_{r,E}=\mathbf{w}_{r}-\mathbf{\tilde{w}}_{r},\ r=0,\cdots,R-1.$ (26)
+Proof.
+The proof can be obtained by substituting the definition of $\mathbf{\bar{w}}_{r,E}$ to the LHS of (26). ∎
+Lemma 26 translates the difference between $\mathbf{w}_{r+1}$ and $\mathbf{\bar{w}}_{r,E}$ to the model distortion incurred by downlink bit errors.
+Lemma 4.
+With Assumption 1, for $r=0,...,R-1$,
+$\begin{split}\mathbb{E}f(\mathbf{w}_{r+1})&\leq\mathbb{E}f(\mathbf{\bar{w}}_{r,E})+\frac{L}{2}\mathbb{E}||\mathbf{w}_{r+1}-\mathbf{\bar{w}}_{r,E}||^{2}\\ &-\mathbb{E}\left\langle\nabla f(\mathbf{\bar{w}}_{r,E}),\bm{\delta}(\mathbf{w}_{r})\right\rangle.\end{split}$ (27)
+Proof.
+The proof can be obtained using the $L$-smoothness of $f(\cdot)$, which is omitted for brevity. ∎
+Lemma 27 provides an upper bound on the expected value of the global loss function at $\mathbf{w}_{r+1}$.
+Lemma 5.
+With Assumptions 1 - 3, for $r=0,\cdots,R-1$,
+$\begin{split}&\mathbb{E}f(\mathbf{\bar{w}}_{r,E})\leq\mathbb{E}f(\mathbf{w}_{r})+\frac{L}{2}\mathbb{E}\|\mathbf{\tilde{w}}_{r}-\mathbf{w}_{r}\|^{2}+\mathbb{E}\left\langle\nabla f(\mathbf{w}_{r}),\bm{\delta}(\mathbf{w}_{r})\right\rangle\\ &-\frac{\eta}{2n}\left[1-L\eta-2L^{2}\eta^{2}E(E-1)\right]\sum_{k=1}^{n}\sum_{e=0}^{E-1}\mathbb{E}\|\nabla f(\mathbf{w}_{r,e}^{k})\|^{2}\\ &-\frac{\eta}{2}\sum_{e=0}^{E-1}\mathbb{E}\|\nabla f(\mathbf{\bar{w}}_{r,e})\|^{2}+\frac{LE\eta^{2}(\sigma_{L}^{2}+\sigma_{G}^{2})}{2n}\\ &+\frac{L^{2}\eta^{3}\sigma_{L}^{2}(n+1)E(E-1)}{2n}.\end{split}$ (28)
+Proof.
+The proof can be obtained following similar lines for Lemma 7 in [20] by further considering the non-IID data distribution among clients (i.e., Assumption 3), which is omitted for brevity. ∎
+Lemma 28 provides an upper bound on the expected value of the global loss function at $\bar{\mathbf{w}}_{r,E}$ in terms of the expected model distortion caused by downlink bit errors, i.e., $\mathbb{E}\|\mathbf{\tilde{w}}_{r}-\mathbf{w}_{r}\|^{2}$.
+By applying Lemmas 26 and 28 to Lemma 27, we obtain the following inequality:
+$\begin{split}&\mathbb{E}f(\mathbf{w}_{r+1})\leq\mathbb{E}f(\mathbf{w}_{r})+L\mathbb{E}\|\mathbf{\tilde{w}}_{r}-\mathbf{w}_{r}\|^{2}\\ &+\mathbb{E}\left\langle\nabla f(\mathbf{w}_{r})-\nabla f(\mathbf{\bar{w}}_{r,E}),\bm{\delta}(\mathbf{w}_{r})\right\rangle-\frac{\eta}{2}\sum_{e=0}^{E-1}\mathbb{E}\|\nabla f(\mathbf{\bar{w}}_{r,e})\|^{2}\\ &-\frac{\eta}{2n}\left[1-L\eta-2L^{2}\eta^{2}E(E-1)\right]\sum_{k=1}^{n}\sum_{e=0}^{E-1}\mathbb{E}\|\nabla f(\mathbf{w}_{r,e}^{k})\|^{2}\\ &+\frac{LE\eta^{2}(\sigma_{L}^{2}+\sigma_{G}^{2})}{2n}+\frac{L^{2}\eta^{3}\sigma_{L}^{2}(n+1)E(E-1)}{2n}.\end{split}$ (29)
+By leveraging $2\left\langle\mathbf{a},\mathbf{b}\right\rangle=\|\mathbf{a}\|^{2}+\|\mathbf{b}\|^{2}-\|\mathbf{a}-\mathbf{b}\|^{2}\leq\|\mathbf{a}\|^{2}+\|\mathbf{b}\|^{2}$ holds for any $\mathbf{a}$, $\mathbf{b}$, an upper bound for the third term on the RHS of (29) is derived as follows:
+$\begin{split}&\mathbb{E}\left\langle\nabla f(\mathbf{w}_{r})-\nabla f(\mathbf{\bar{w}}_{r,E}),\bm{\delta}(\mathbf{w}_{r})\right\rangle\\ &\leq\frac{\mathbb{E}\|\nabla f(\mathbf{w}_{r})-\nabla f(\mathbf{\bar{w}}_{r,E})\|^{2}}{2}+\frac{\mathbb{E}\|\bm{\delta}(\mathbf{w}_{r})\|^{2}}{2}\\ &\overset{(\text{a})}{\leq}\frac{L^{2}}{2}\mathbb{E}\|\mathbf{\bar{w}}_{r,E}-\mathbf{w}_{r}\|^{2}+\frac{\mathbb{E}\|\bm{\delta}(\mathbf{w}_{r})\|^{2}}{2}\\ &=\frac{L^{2}}{2}\mathbb{E}\bigg{\|}\frac{1}{n}\sum_{k=1}^{n}(\mathbf{w}_{r,E}^{k}-\mathbf{w}_{r})\bigg{\|}^{2}+\frac{\mathbb{E}\|\bm{\delta}(\mathbf{w}_{r})\|^{2}}{2}\\ &=\frac{L^{2}}{2}\mathbb{E}\bigg{\|}\frac{1}{n}\sum_{k=1}^{n}(\mathbf{w}_{r,E}^{k}-\tilde{\mathbf{w}}_{r})+\tilde{\mathbf{w}}_{r}-\mathbf{w}_{r}\bigg{\|}^{2}+\frac{\mathbb{E}\|\bm{\delta}(\mathbf{w}_{r})\|^{2}}{2}\\ &\overset{(\text{b})}{=}\frac{L^{2}}{2}\mathbb{E}\bigg{\|}\frac{\eta}{n}\sum_{k=1}^{n}\sum_{e=0}^{E-1}\nabla f_{k}(\mathbf{w}_{e}^{k};\xi_{r,e}^{k})+\tilde{\mathbf{w}}_{r}-\mathbf{w}_{r}\bigg{\|}^{2}\\ &\ \ \ +\frac{\mathbb{E}\|\bm{\delta}(\mathbf{w}_{r})\|^{2}}{2}\end{split}$ (30)
+where (a) is due to the $L$-smoothness of $f(\cdot)$ and (b) is due to the local updating rules. Based on the Assumptions 2 and 3, we have
+$\begin{split}\mathbb{E}\|\nabla&f_{k}(\mathbf{w}_{r,e}^{k};\xi_{r,e}^{k})\|^{2}=\mathbb{E}\|\nabla f_{k}(\mathbf{w}_{r,e}^{k};\xi_{r,e}^{k})\\ &-\nabla f_{k}(\mathbf{w})+\nabla f_{k}(\mathbf{w})-\nabla f(\mathbf{w})+\nabla f(\mathbf{w})\|^{2}\\ &\leq\sigma_{L}^{2}+\sigma_{G}^{2}+\mathbb{E}\|\nabla f(\mathbf{w})\|^{2}.\end{split}$ (31)
+By combing (31) and (30), the bound in (30) can be further updated as
+$\begin{split}&\mathbb{E}\left\langle\nabla f(\mathbf{w}_{r})-\nabla f(\mathbf{\bar{w}}_{r,E}),\bm{\delta}(\mathbf{w}_{r})\right\rangle\\ &\leq\frac{L^{2}\eta^{2}}{2n^{2}}\sum_{k=1}^{n}\sum_{e=0}^{E-1}\mathbb{E}\|\nabla f_{k}(\mathbf{w}_{r,e}^{k};\xi_{r,e}^{k})\|^{2}+\frac{L^{2}}{2}\mathbb{E}\|\tilde{\mathbf{w}}_{r}-\mathbf{w}_{r}\|^{2}\\ &\ \ \ +\frac{\mathbb{E}\|\bm{\delta}(\mathbf{w}_{r})\|^{2}}{2}\\ &\overset{(\text{b})}{\leq}\frac{L^{2}\eta^{2}}{2n^{2}}\sum_{k=1}^{n}\sum_{e=0}^{E-1}\mathbb{E}\|\nabla f(\mathbf{w}_{r,e}^{k})\|^{2}+\frac{L^{2}\eta^{2}E(\sigma_{G}^{2}+\sigma_{L}^{2})}{2n}\\ &\ \ \ +\frac{L^{2}}{2}\mathbb{E}\|\tilde{\mathbf{w}}_{r}-\mathbf{w}_{r}\|^{2}+\frac{\mathbb{E}\|\bm{\delta}(\mathbf{w}_{r})\|^{2}}{2},\end{split}$ (32)
+Substituting (32) in (29), we have
+$\begin{split}&\mathbb{E}f(\mathbf{w}_{r+1})\leq\mathbb{E}f(\mathbf{w}_{r})+\frac{L(L+2)}{2}\mathbb{E}\|\mathbf{w}_{r}-\mathbf{\tilde{w}}_{r}\|^{2}+\frac{\mathbb{E}\|\bm{\delta}(\mathbf{w}_{r})\|^{2}}{2}\\ &-\frac{\eta}{2n}\left[1-\frac{n+L}{n}L\eta-2L^{2}\eta^{2}E(E-1)\right]\\ &\cdot\sum_{k=1}^{n}\sum_{e=0}^{E-1}\mathbb{E}\|\nabla f(\mathbf{w}_{r,e}^{k})\|^{2}-\frac{\eta}{2}\sum_{e=0}^{E-1}\mathbb{E}\|\nabla f(\mathbf{\bar{w}}_{r,e})\|^{2}\\ &+\frac{L(L+1)E\eta^{2}(\sigma_{L}^{2}+\sigma_{G}^{2})}{2n}+\frac{L^{2}\eta^{3}\sigma_{L}^{2}(n+1)E(E-1)}{2n}.\end{split}$ (33)
+For any $\eta$ that satisfies (8), we have
+$\begin{split}&\mathbb{E}f(\mathbf{w}_{r+1})\leq\mathbb{E}f(\mathbf{w}_{r})+\frac{L(L+2)}{2}\mathbb{E}\|\mathbf{w}_{r}-\mathbf{\tilde{w}}_{r}\|^{2}\\ &+\frac{\mathbb{E}\|\bm{\delta}(\mathbf{w}_{r})\|^{2}}{2}-\frac{\eta}{2}\sum_{e=0}^{E-1}\mathbb{E}\|\nabla f(\mathbf{\bar{w}}_{r,e})\|^{2}\\ &+\frac{L(L+1)E\eta^{2}(\sigma_{L}^{2}+\sigma_{G}^{2})}{2n}+\frac{L^{2}\eta^{3}\sigma_{L}^{2}(n+1)E(E-1)}{2n}.\end{split}$ (34)
+Summing (34) over $r=0,\cdots,R-1$, rearranging the terms, and multiplying $\frac{2}{\eta RE}$ on both sides, we obtain the following expression:
+$\begin{split}&\frac{1}{RE}\sum_{r=0}^{R-1}\sum_{e=0}^{E-1}\mathbb{E}\|\nabla f(\mathbf{\bar{w}}_{r,e})\|^{2}\leq\frac{2(\mathbb{E}f(\mathbf{w}_{0})-\mathbb{E}f(\mathbf{w}_{R}))}{\eta RE}\\ &+\frac{L(L+2)}{\eta RE}\sum_{r=0}^{R-1}\mathbb{E}\|\mathbf{\tilde{w}}_{r}-\mathbf{w}_{r}\|^{2}+\frac{1}{\eta RE}\sum_{r=0}^{R-1}\mathbb{E}\|\bm{\delta}(\mathbf{w}_{r})\|^{2}\\ &+\frac{L(L+1)\eta(\sigma_{L}^{2}+\sigma_{G}^{2})}{n}+\frac{L^{2}\eta^{2}\sigma_{L}^{2}(n+1)(E-1)}{n}.\end{split}$ (35)
+Based on (23) and (21) respectively, we have
+$\begin{split}\mathbb{E}\left[||\mathbf{\tilde{w}}_{r}-\mathbf{w}_{r}||^{2}\right]&=\mathbb{E}\left[\mathbb{E}\left[||\mathbf{\tilde{w}}_{r}-\mathbf{w}_{r}||^{2}|\mathbf{w}_{r}\right]\right]\\ &=\frac{d(4^{N}-1)}{3(2^{N}-1)^{2}}\cdot b_{r}\left(1-b_{r}\right)^{N-1}\cdot M_{r}^{2},\end{split}$ (36)
+$\begin{split}\mathbb{E}\|\bm{\delta}(\mathbf{w}_{r})\|^{2}&=\sum_{j=1}^{d}\mathbb{E}\left[|\delta(w_{r,d})|^{2}|{w}_{r,d}\right]\\ &\leq\sum_{j=1}^{d}\mathbb{E}\left[(b_{r}(1-b_{r})^{N-1}M)^{2}|{w}_{r,d}\right]\\ &\leq d(b_{r}(1-b_{r})^{N-1})^{2}M_{r}^{2}.\end{split}$ (37)
+By substituting (36) and (37) into (35), and denote $f^{*}$ as the minimum global training loss, we have
+$\frac{1}{RE}\sum_{r=0}^{R-1}\sum_{e=0}^{E-1}\mathbb{E}\|\nabla f(\mathbf{\bar{w}}_{r,e})\|^{2}\leq\frac{2(\mathbb{E}f(\mathbf{w}_{0})-f^{*})}{\eta RE}$
+$+\frac{L(L+2)d(4^{N}-1)}{3(2^{N}-1)^{2}\eta RE}\sum_{r=0}^{R-1}b_{r}(1-b_{r})^{N-1}M_{r}^{2}$
+$+\frac{d}{\eta RE}\sum_{r=0}^{R-1}(b_{r}(1-b_{r})^{N-1})^{2}M_{r}^{2}$
+$+\frac{L(L+1)\eta(\sigma_{L}^{2}+\sigma_{G}^{2})}{n}+\frac{L^{2}\eta^{2}\sigma_{L}^{2}(n+1)(E-1)}{n}.$ (38)
+Denoting $\hat{M}=max(M_{0},M_{1},\cdots,M_{R-1})$, we have
+$\frac{1}{RE}\sum_{r=0}^{R-1}\sum_{e=0}^{E-1}\mathbb{E}\|\nabla f(\mathbf{\bar{w}}_{r,e})\|^{2}\leq\frac{2(\mathbb{E}f(\mathbf{w}_{0})-f^{*})}{\eta RE}$
+$+\frac{L(L+2)d\hat{M}^{2}(4^{N}-1)}{3(2^{N}-1)^{2}\eta RE}\sum_{r=0}^{R-1}b_{r}(1-b_{r})^{N-1}$
+$+\frac{d\hat{M}^{2}}{\eta RE}\sum_{r=0}^{R-1}(b_{r}(1-b_{r})^{N-1})^{2}$
+$+\frac{L(L+1)\eta(\sigma_{L}^{2}+\sigma_{G}^{2})}{n}+\frac{L^{2}\eta^{2}\sigma_{L}^{2}(n+1)(E-1)}{n}.$ (39)

### v3

@@
 One of the most critical challenges for deploying distributed learning solutions, such as federated learning (FL), in wireless networks is the limited battery capacity of mobile clients.
-While it is a common belief that the major energy consumption of mobile clients comes from the uplink data transmission, this paper presents a novel finding, namely the channel decoding operation also contributes significantly to the overall energy consumption of mobile clients in FL.
+While it is a common belief that the major energy consumption of mobile clients comes from the uplink data transmission, this paper presents a novel finding, namely channel decoding also contributes significantly to the overall energy consumption of mobile clients in FL.
 Motivated by this new observation, we propose an energy-efficient adaptive channel decoding scheme that leverages the intrinsic robustness of FL to model errors.
 In particular, the robustness is exploited to reduce the energy consumption of channel decoders at mobile clients by adaptively adjusting the number of decoding iterations.
-We theoretically prove that wireless FL with communication errors can converge at the same rate as the case with error-free communication as long as the bit error rate (BER) is properly constrained.
+We theoretically prove that wireless FL with communication errors can converge at the same rate as the case with error-free communication provided the bit error rate (BER) is properly constrained.
 An adaptive channel decoding scheme is then proposed to improve the energy efficiency of wireless FL systems.
-Experimental results demonstrate that the proposed method maintains the same learning accuracy while reducing the channel decoding energy consumption by 20% when compared to existing approaches.
+Experimental results demonstrate that the proposed method maintains the same learning accuracy while reducing the channel decoding energy consumption by $\sim$ 20% when compared to an existing approach.
 Index Terms:
@@
 I-A Motivation
-Deep learning has achieved remarkable success in numerous realms, such as computer vision and natural language processing [1].
+Deep learning has achieved remarkable successes in numerous realms, such as computer vision and natural language processing [2].
 However, the conventional centralized learning paradigm encounters the data isolation issue due to the rising concern about privacy [3].
 As a result, uploading the distributed data, which may contain privacy-sensitive information, for centralized learning becomes prohibited.
-To address this issue, federated learning (FL) was proposed [3] to enable the distributed training of deep learning models without sharing the data of users (e.g., IoT devices).
-However, to fully unlock the potential of FL, we need to tackle several critical challenges, such as data and device heterogeneity [4, 5, 6], as well as the communication bottleneck[7].
-Another issue of wireless FL is the energy consumption of mobile clients due to heavy computation and communication workloads for training and transferring the deep neural network (DNN) parameters[8].
-Many innovative approaches have been proposed to reduce the energy consumption of mobile clients in FL from different perspectives.
-For instance, reinforcement learning was utilized in [9] to identify near-optimal client selection and resource allocation to reduce both the convergence time and energy consumption of mobile clients.
-In [10], the trade-off between learning accuracy and total energy consumption of users was optimized through computational and communication resource allocation.
-In [11], the transmission and computation energy was minimized by bandwidth allocation and power control.
-Besides system-level resource management, techniques such as model sparsification [12] and quantization [13, 14, 15, 16, 17, 18] have been proposed to minimize the communication overhead, thus reducing the energy consumption of the transmitters.
-Authors of [19] minimized the transmission energy consumption of mobile clients while maintaining the learning performance by optimizing the transmit power, user selection and uplink resource allocation.
-Authors of [20] adopted quantization and periodic aggregation to reduce the transmission volume and energy, which also provided a theoretical framework about the impact of quantization and periodic aggregation on the convergence rate of FL.
-Model sparsification and quantization techniques are also beneficial for minimizing the processing energy consumption [21, 8].
-Authors of [21] proposed to introduce sparsity at training time to reduce the on-device training energy.
-Authors of [8] reduced the local computation energy by quantization on both DNN weights and activations.
-However, most works on energy-efficient FL focused on the energy consumption of mobile transmitters while few of them paid attention to the energy consumption of channel decoders in mobile receivers.
-As will be shown in Section II-B, channel decoders consume a substantial proportion of energy because of the iterative processes to decode massive DNN parameters in each communication round of wireless FL.
+To address this issue, federated learning (FL) was proposed [4] to enable the distributed training without sharing the data of users (e.g., IoT devices).
+However, to fully unlock the potential of FL, we need to tackle several critical challenges, such as data and device heterogeneity [5, 6], as well as the communication bottleneck[7].
+Another issue of wireless FL is the energy consumption of mobile clients due to heavy computation and communication workloads for training deep neural networks (DNNs) and transferring model parameters[8].
+Many innovative approaches have been proposed to reduce the energy consumption of mobile clients in wireless FL from different perspectives.
+For instance, reinforcement learning was utilized in [9] to determine the near-optimal client selection and resource allocation scheme to reduce both the convergence time and energy consumption of mobile clients.
+In [10], the trade-off between learning accuracy and total energy consumption of clients was optimized through computational and communication resource allocation.
+In [11], the transmission and computation energy was minimized under a latency constraint by optimizing the time, bandwidth allocation, power control, computation frequencies, and learning accuracy.
+Also, authors of [12] minimized the transmission energy consumption of mobile clients while maintaining the learning performance by further designing the client selection strategies.
+Besides system-level resource management, techniques such as model sparsification [13] and quantization [14, 15] have been proposed to minimize the communication overhead, thus reducing the energy consumption of clients’ transmitters.
+Both quantization and periodic aggregation were adopted in [16] to reduce the communication overhead and energy consumption, where a theoretical framework about the impact of parameter quantization on the convergence rate was also provided.
+Model sparsification and quantization techniques are also beneficial for minimizing the DNN processing energy consumption at mobile clients of wireless FL [17, 18].
+However, most works on energy-efficient FL focused on the energy consumption of mobile processors and transmitters, while few of them paid attention to the energy consumption of channel decoders in mobile receivers.
+As will be shown in Section II-B, channel decoders may consume a substantial proportion of energy because of the iterative processes to decode massive DNN parameters in each communication round of wireless FL.
 To address the research gap, in this work, we first quantitatively analyze the energy consumption of different components in a mobile unit and show that the channel decoder contributes significantly to the overall energy consumption.
 Building on this observation, an energy-efficient adaptive channel decoding scheme is proposed for wireless FL systems, by leveraging the fact that FL is robust to model errors due to its statistical learning nature[19].
-Furthermore, by extending the convergence analysis in[20] to the scenario of FL with downlink communication erros, we investigate the relationship between the bit error rate (BER) and the convergence rate of wireless FL, and determine the maximum tolerable BER that can maintain the same convergence rate as the case with error-free communication.
-Finally, we propose a practical approach to adjust the number of channel decoding iterations in each communication round to achieve the maximum tolerable BER, while reducing the energy consumption of mobile clients in wireless FL.
+Furthermore, by extending the convergence analysis in[16] to scenarios with downlink communication erros, we formalize the relationship between the bit error rate (BER) and the convergence performance of wireless FL.
+A decreasing rule of BER with respect to communications rounds is determined, which can maintain the same convergence rate as the case with error-free communication.
+Finally, we propose a practical approach to adjust the number of channel decoding iterations in each communication round to achieve the desired BER, while reducing the energy consumption of mobile clients in wireless FL.
 I-B Contributions
 This paper investigates the impact of channel decoding on the energy efficiency of wireless FL. Our major contributions are summarized as follows:
-  • We theoretically analyze the relationship between BER and the convergence rate of wireless FL.
-Specially, an upper bound of the average expected gradient norm of the global training loss with respect to BERs over communication rounds is derived by extending the analysis in [20], based on which, the maximum tolerable BER to achieve the same convergence rate as error-free FL is determined.
-  • Based on the theoretical analysis, an adaptive low-density parity checking (LDPC) decoding mechanism is proposed, which controls the maximum number of decoding iterations to satisfy the BER requirement in every communication round and maintain the target convergence rate as error-free wireless FL, while reducing the energy consumption of channel decoders in wireless FL.
-  • Experimental results on both IID and non-IID datasets show that the proposed adaptive scheme effectively reduces the energy consumption of channel decoders by 20% when compared with the traditional schemes which set a fixed maximum number of channel decoding iterations, while guaranteeing the same learning performance.
+  • We theoretically analyze the relationship between BER and the convergence performance of wireless FL.
+Specially, an upper bound of the average expected gradient norm of the global training loss function with respect to BERs over communication rounds is derived by extending the analysis in [16].
+Based on the analysis, a decreasing rule of BERs is proposed to achieve the same convergence rate as FL with error-free communication.
+  • With the decreasing rule of BERs, an adaptive low-density parity checking (LDPC) decoding mechanism is proposed, which controls the maximum number of decoding iterations to achieve the desired BER requirement in every communication round.
+This provides a practical approach to maintain the target convergence rate, while reducing the channel decoding energy consumption of mobile clients in wireless FL.
+  • Experimental results on both IID and non-IID datasets show that with a given number of communication rounds, the proposed adaptive LDPC decoding scheme effectively reduces the energy consumption of channel decoders by 20% when compared with a baseline scheme adopting a fixed maximum number of channel decoding iterations, while enjoying the same learning performance.
 I-C Organization
 The remainder of this paper is organized as follows.
-Section II introduces the system model and presents a comprehensive energy consumption analysis for mobile units.
+Section II introduces the system model and presents a comprehensive energy consumption analysis for mobile clients.
 Section III develops an energy-efficient adaptive LDPC decoding scheme for wireless FL systems by exploiting the relationship between the convergence rate and the achieved BER of LDPC decoding.
@@
 Finally, conclusions are drawn in Section V.
-II System Model and Preliminaries
+II System Model and Energy Consumption Evaluation
 In this section, we first introduce the wireless FL system.
@@
 Notation Definition
-$\mathbf{w}$ DNN model
+$\mathbf{w}_{r}$ Global model in the $r$-th communication round
 $\Delta\mathbf{w}_{r}^{k}$ Local model update of client $k$ at the $r$-th communication round
-$f$ The loss function of training
-$f^{*}$ Minimum value of the global training loss
-$R$ The total communication rounds
+$f\left(\cdot\right)$ The global loss function
+$f^{*}$ Minimum global training loss
+$R$ Number of communication rounds
 $E$ Number of local training steps in each communication round
-$\sigma_{L}^{2}$ Variance of local gradient estimator
-$\sigma_{G}^{2}$ Dissimilarity between local and global gradient
-$n$ Number of total clients
-$N$ Each DNN parameter is with $N$ bit representation
-$d$ Dimension of the DNN model
-$T$ Total training iterations, i.e., $T=RE$.
-$L$ The $L$-smooth parameter
+$\sigma_{L}^{2}$ Variance bound of local gradient estimator
+$\sigma_{G}^{2}$ Dissimilarity bound between local and global gradients
+$K$ Number of clients
+$N$ Each global model parameter is digitalized as an $N$-bit sequence
+$D$ Dimension of model $\mathbf{w}$
+$T$ Total number of local training steps, $T=RE$
+$L$ $L$-smoothness parameter
 $b_{r}$ Bit error rate in the $r$-th communication round
 $\eta$ Learning rate
-$Q$ Maximum number of LDPC decoding iterations
-$M_{r}$ Range of the global model parameters in the $r$-th communication round
-$\hat{M}$ Maximum range of global model parameters for all communication rounds
-(a) Wireless FL system: Multiple mobile clients are coordinated by one remote server through noisy wireless channels, where model updates and model weights are communicated in the uplink and downlink, respectively.
+$Q_{r}$ Maximum number of LDPC decoding iterations in the $r$-th communication round
+$M$ Bound of $\max(\mathbf{w}_{r})-\min(\mathbf{w}_{r}),r=0,\cdots,R-1$
+(a) Wireless FL system: Multiple mobile clients are coordinated by a remote server through noisy wireless channels, where model updates and model weights are communicated in the uplink and downlink, respectively.
 (b) System operations: The server performs model aggregation, digitalization, LDPC encoding, and model broadcasting.
-Mobile clients perform LDPC decoding, local training, and uploading the model updates.
-Figure 1: Wireless FL in digital communication network and its operations.
+Mobile clients perform LDPC decoding, local training, and upload the model updates.
+Procedures of uploading model updates are omitted.
+Figure 1: Wireless FL with digital communication and its operations.
 II-A Wireless FL
-Fig. 1(a) depicts a wireless FL system, where model training of different clients is coordinated by a remote server. The FL problem can be formulated as follows [23, 24, 25]:
-$\min_{\mathbf{w}\in{\mathbb{R}^{d}}}f(\mathbf{w})\triangleq\frac{1}{n}\sum_{k=1}^{n}f_{k}(\mathbf{w}),$ (1)
-where $\mathbf{w}\in{\mathbb{R}^{d}}$ denotes the DNN model with $d$ parameters, $n$ represents the number of mobile clients, $f_{k}\left(\cdot\right)$ is the local loss function of the $k$-th client, and $f\left(\cdot\right)$ is the global loss function.
-Figure 2: Digitalization scheme for model parameters. This part gives an example, say [min, max] is digitalized by 3-bits and shows how each value inside is mapped to ”000”, ”001”, ”…”, ”111”.
+Fig. 1(a) depicts a wireless FL system, where model training of different clients is coordinated by a remote server. The FL problem can be formulated as follows [4]:
+$\min_{\mathbf{w}\in{\mathbb{R}^{D}}}f(\mathbf{w})\triangleq\frac{1}{K}\sum_{k=1}^{K}f_{k}(\mathbf{w}),$ (1)
+where $\mathbf{w}\in{\mathbb{R}^{D}}$ denotes the DNN model with $D$ parameters, $K$ represents the number of mobile clients, $f_{k}\left(\cdot\right)$ is the local loss function of the $k$-th client, and $f\left(\cdot\right)$ is the global loss function.
+Figure 2: Digitalization for each parameter in $\tilde{\mathbf{w}}_{r}$ (an example with $N=3$), where a continuous model parameter is mapped to the binary sequence of the nearest boundary.
 To train a shared global model, the server communicates with the clients for multiple rounds through noisy wireless channels.
 As shown in Fig.
-1(b), in each communication round, the global model at the server is broadcast to the clients.
+1(b), in each communication round, the global model at the server (denoted as $\mathbf{w}_{r}$ for the $r$-th communication round) is broadcast to the clients.
 Each client then performs local training based on the received global model and transmits the model updates after local training to the server for aggregation.
-Specifically, in the $r$-th communication round, the global model is digitalized by $N$ bits, encoded by a channel code, and broadcast to clients.
-The model parameter digitalization scheme is illustrated in Fig.
-2, where the range of [min, max] is divided into $12^{N}-$ 1 intervals and the width of each interval is $\Delta=\frac{\text{max-min}}{2^{N}-1}$.
-In this paper, we assume $N$ is large enough so the continuous and digital representations of model parameters have negligible differences.
-The model parameters received by the $k$-th client in the $r$-th communication round is given as follows:
-$\mathbf{w}^{k}_{r,0}=\mathbf{\tilde{w}}_{r},\ k=1,\cdots,n,\ r=0,\cdots,R-1,$ (2)
+Specifically, in the $r$-th communication round, the global model is digitalized by an $N$-bit sequence, encoded by a channel code (e.g., the LDPC codes), and broadcast to clients.
+The parameter digitalization scheme is illustrated in Fig.
+2, where the range of [$\min(\mathbf{w}_{r}),\max(\mathbf{w}_{r})$] is evenly divided into $12^{N}-$ 1 intervals and the width of each interval is $\Delta_{r}=\frac{\max(\mathbf{w}_{r})-\min(\mathbf{w}_{r})}{2^{N}-1}$.
+Each of the $2^{N}$ boundaries of these intervals are marked by a $N$-bit binary sequence.
+A parameter of $\mathbf{w}_{r}$ is mapped to the binary sequence of the nearest boundary.
+We assume $N$ is large enough so the continuous and digital representations of model parameters have negligible difference.
+The model parameters received/decoded by the $k$-th client at the $r$-th communication round is given as follows:
+$\mathbf{w}^{k}_{r,0}=\mathbf{\tilde{w}}_{r},\ k=1,\cdots,K,\ r=0,\cdots,R-1,$ (2)
 where $\mathbf{\tilde{w}}_{r}$ denotes the distorted version of $\mathbf{w}_{r}$ due to channel decoding errors.
+To reconstruct $\tilde{\mathbf{w}}_{r}$ from the decoded bits at clients, the server also broadcasts the scalars $\min(\mathbf{w}_{r})$ and $\max(\mathbf{w}_{r})$ through an error-free downlink channel.
 The impact of channel decoding errors can be quantified by considering the different error bit positions in the $N$-bit representations of model parameters, which will be analyzed in Section III-A.
-Each client updates the received global model by conducting $E$ steps of local mini-batch Stochastic Gradient Descent (SGD) with a learning rate $\eta$ as follows:
-$\begin{split}\mathbf{w}^{k}_{r,e+1}&=\mathbf{w}^{k}_{r,e}-\eta{\nabla}f_{k}(\mathbf{w}^{k}_{r,e};\xi_{r,e}^{k}),\\ &e=0,\cdots,E-1,\ k=1,\cdots,n,\end{split}$ (3)
+Each client updates the received global model $\tilde{\mathbf{w}}_{r}$ by conducting $E$ steps of local mini-batch stochastic gradient descent (SGD) with a learning rate $\eta$ as follows:
+$\begin{split}\mathbf{w}^{k}_{r,e+1}=\mathbf{w}^{k}_{r,e}-\eta&{\nabla}f_{k}(\mathbf{w}^{k}_{r,e};\xi_{r,e}^{k}),\\ &e=0,\cdots,E-1,\ k=1,\cdots,K,\end{split}$ (3)
 where ${\nabla}f_{k}(\mathbf{w}^{k}_{r,e};\xi_{r,e}^{k})$ denotes the gradient computed from a mini-batch datasets $\xi_{r,e}^{k}$ at the $k$-th client.
 By completing $E$ steps of local training, the local model update of the $k$-th client can be expressed as follows:
-$\begin{split}\Delta\mathbf{w}^{k}_{r}&=\mathbf{w}^{k}_{r,E}-\mathbf{w}^{k}_{r,0},\\ &r=0,\cdots,R-1,\ k=1,\cdots,n,\end{split}$ (4)
+$\begin{split}\Delta\mathbf{w}^{k}_{r}=\mathbf{w}^{k}_{r,E}-&\mathbf{w}^{k}_{r,0},\\ &r=0,\cdots,R-1,\ k=1,\cdots,K,\end{split}$ (4)
 which is also digitalized, encoded, and uploaded to the server.
@@
 After receiving these local updates, denoted as $\Delta\mathbf{\tilde{w}}^{k}_{r}$, the server performs model aggregation to update the global model as follows:
-$\mathbf{w}_{r+1}=\mathbf{w}_{r}+\frac{1}{n}\sum_{k=1}^{n}\Delta\mathbf{\tilde{w}}^{k}_{r},\ r=0,\cdots,R-1.$ (5)
+$\mathbf{w}_{r+1}=\mathbf{w}_{r}+\frac{1}{K}\sum_{k=1}^{K}\Delta\mathbf{\tilde{w}}^{k}_{r},\ r=0,\cdots,R-1.$ (5)
 In this work, we focus on the downlink communication rather than the uplink since our objective is to design novel channel decoding schemes to save the energy consumption of mobile clients.
 For the uplink, the channel decoding happens at the server, where a powerful channel decoder as well as a sufficient number of decoding iterations can always be applied to achieve satisfactory BER performance.
-As such, we consider $\Delta\mathbf{\tilde{w}}_{r}^{k}=\Delta\mathbf{w}_{r}^{k}$ in the following.
+Hence, we consider $\Delta\mathbf{\tilde{w}}_{r}^{k}=\Delta\mathbf{w}_{r}^{k}$ in the following.
+Define $\bar{\mathbf{w}}_{r,e}\triangleq\frac{1}{K}\sum_{k=1}^{K}\mathbf{w}_{r,e}^{k}$ as the average of client models after $e$ local training steps in the $r$-th communication round.
+After $R$ communication rounds, the wireless FL system obtains $\bar{\mathbf{w}}_{R,E}$ as the learned model.
 II-B Energy Consumption Evaluation
@@
 For the latter, it is believed that the energy is mostly consumed by the data transmission.
-In this section, we evaluate the energy consumption of different components, taking the training of AlexNet for image classification tasks as an example [26].
+In this subsection, we evaluate the energy consumption of different components, taking the training of AlexNet for image classification tasks as an example [20].
 The analysis is based on actual measurements obtained from fabricated chips published in recent years [21, 22, 23].
@@
 With five epochs of local training, the processing energy consumption in each communication round is 68.5 mJ.
-In [28], a fifth-generation (5G) New Radio wireless transmitter and receiver were reported to achieve the energy efficiencies of 47.9 pJ/b and 33.3 pJ/b for transferring uncoded data, respectively, under a communication distance of around 200 meters.
+In [22], a fifth-generation (5G) New Radio wireless transmitter and receiver were reported to achieve the energy efficiencies of 47.9 pJ/bit and 33.3 pJ/bit for transferring uncoded data, respectively, under a communication distance of around 200 meters.
 Therefore, the energy required to transmit and receive the 8-bit AlexNet with 60 million parameters in a single communication round is about 39 mJ.
 Assuming a code rate of 1/2, the communication energy consumption for coded transmission of the AlexNet model parameters in each round is 78 mJ.
-Besides, the LDPC decoder in [29] achieves the energy efficiency of 20.1 pJ/b/iteration, which was measured with 10 decoding iterations on average.
+Besides, the LDPC decoder in [23] achieves the energy efficiency of 20.1 pJ/bit/iteration, which was measured with 10 decoding iterations on average.
 Accordingly, the energy consumption of the LDPC decoder for decoding the AlexNet model parameters is 96.5 mJ.
 The detailed comparison is presented in Table II, which indicates that the energy consumption of LDPC decoding accounts for a large proportion of the overall energy consumption at clients in wireless FL.
-TABLE II: Energy Consumption of Main Components at a Mobile Unit in Wireless FL for Training an AlexNet
+TABLE II: Energy Consumption of Main Components at a Mobile Unit in Wireless FL for Training an 8-bit AlexNet
 Component Energy efficiency Energy consumption
 AI Chip 13.7 mJ/epoch[21] 68.5 mJ
-RF Transceiver 81.2 pJ/b[28] 78 mJ
-LDPC Decoder 20.1 pJ/b/iter[29] 96.5 mJ
+RF Transceiver 81.2 pJ/bit[22] 78 mJ
+LDPC Decoder 20.1 pJ/bit/iter.[23] 96.5 mJ
 II-C LDPC Decoding
-In this work, we use LDPC codes as an example to illustrate how adaptive channel decoding can help reduce the energy consumption of wireless FL.
-In LDPC, the iterative decoding process continues until $Q$ decoding iterations, denoted as the maximum number of decoding iterations, are completed [30].
+In this paper, we take the LDPC codes as an example to illustrate how adaptive channel decoding can help reduce the energy consumption of wireless FL.
+In LDPC decoders, the iterative decoding process continues until $Q$ decoding iterations, which is termed as the maximum number of decoding iterations, are completed [24].
 Such iterative LDPC decoding processing leads to high energy consumption.
 To save the decoding energy, early termination techniques are usually adopted.
-Specifically, if the decoded codewords can fully pass the parity check, the decoding process will stop even if not all the $Q$ decoding iterations has been completed[31].
-The early termination scheme can save decoding iterations without compromising the BER performance.
+Specifically, if the decoded codewords can fully pass the parity check, the decoding process will stop even if not all the $Q$ decoding iterations have been completed[25].
+Therefore, early termination can save some decoding iterations without compromising the BER performance.
 The energy consumption of LDPC decoding depends on the target BER and thus the pre-defined maximum number of decoding iterations.
-As a result, to minimize the energy consumption, it is essential to determine the maximum tolerable BER by FL systems, considering its unique statistical learning nature.
+As a result, to minimize the decoding energy consumption, it is essential to determine suitable BERs over communication rounds that can be tolerated by wireless FL systems, considering its unique statistical learning nature.
 III Energy-Efficient LDPC Decoding Design
@@
 III-A Effect of BER on the Convergence Rate of Wireless FL
-With error-free communication, the convergence rate of FL with non-convex loss functions is proved as $\mathcal{O}(1/\sqrt{T})$ [20], where $T$ is the total number of SGD iterations.
-In this subsection, we investigate how BER affects the convergence rate of FL and determine the maximum admissible BER for maintaining the convergence rate of $\mathcal{O}(1/\sqrt{T})$.
-Unlike FL with analog communication[32, 33, 34, 35, 36, 37, 38], quantifying the impact of BER on FL with digital communication is not straightforward.
+With error-free communication, the convergence rate of FL with non-convex loss functions was proved as $\mathcal{O}(1/\sqrt{T})$ [16], where $T=RE$ is the total number of local training steps.
+In this subsection, we investigate how BER affects the convergence rate of FL and determine the admissible BER for maintaining the convergence rate of $\mathcal{O}(1/\sqrt{T})$.
+Unlike FL with analog communication where model distortion is considered as additive noise[26, 27, 28, 29, 30, 31], quantifying the impact of BER on FL with digital communication is not straightforward.
 In particular, we need to consider different cases of bit errors and analyze the impact of different bit errors, e.g., most significant bits and least significant bits.
-To conduct the theoretical analysis, we take the following assumptions that are widely adopted in the literature.
+To conduct the theoretical analysis, we take the following assumptions that are widely adopted in literature.
 Assumption 1.
@@
 Assumption 3.
-(Bounded Local and Global Gradient Dissimilarity [39]) The dissimilarity between the local and global gradient is bounded, i.e., $||\nabla f_{k}(\mathbf{w})-\nabla f(\mathbf{w})||^{2}\leq\sigma_{G}^{2}$.
+(Bounded Local and Global Gradient Dissimilarity [32]) The dissimilarity between the local and global gradients is bounded, i.e., $||\nabla f_{k}(\mathbf{w})-\nabla f(\mathbf{w})||^{2}\leq\sigma_{G}^{2}$.
 Assumption 4.
-(Independent Bit Errors [40]) Each bit in the bitstream of the decoded model parameters encounters error independently.
+(Independent Bit Errors [33]) Each bit in the bitstream of the model parameters encounters error (i.e., bit flipping) independently.
 With the $N$-bit representation, each model parameter has $12^{N}-$ 1 possible distorted patterns.
@@
 In particular, Fig.
-4 compares the actual model error with the analytical model error expression in (7) assuming at most 1-bit error, which match closely with various BER.
-Therefore, we impose the following assumption to simplify our analysis.
+4 compares the actual model error with the analytical model error expression to be presented in (7) assuming at most 1-bit error, which match closely under various BER conditions.
+Therefore, we also impose the following assumption to simplify our analysis.
 Assumption 5.
-(At Most 1-bit Error of each Model Parameter) There is at most 1-bit error in the $N$-bit representation of each model parameter.
-Based on Assumptions 4 and 5, we are able to establish the relationship between the decoded model at a client and the broadcasted global model in the following lemma.
+There is at most 1-bit error in the $N$-bit representation of each model parameter.
+Based on Assumptions 4 and 5, we are able to establish the relationship between the decoded model at a client and the broadcast global model in the following lemma.
 Lemma 1.
-Denote $\tilde{\mathbf{w}}$ as the decoded model of $\mathbf{w}$. Based on Assumptions 4 and 5, the bias of each dimension in $\mathbf{w}$ is given as
-$\tilde{w}_{d}=w_{d}+\delta\left(w_{d}\right).$ (6)
-The mean square error caused by bit errors is given as
-$\mathbb{E}[||\mathbf{\tilde{w}}-\mathbf{w}||^{2}|\mathbf{w}]=\frac{d\left(4^{N}-1\right)}{3\left(2^{N}-1\right)^{2}}\cdot b(1-b)^{N-1}M^{2},$ (7)
-where $b$ is the BER.
+Denote $\tilde{\mathbf{w}}$ as the decoded model of $\mathbf{w}$. Based on Assumptions 4 and 5, the $d$-th dimension of $\tilde{\mathbf{w}}$ is given as
+$\tilde{w}_{d}=w_{d}+\delta\left(w_{d}\right),$ (6)
+where $w_{d}$ is the $d$-th dimension of $\mathbf{w}$ and $\delta\left(w_{d}\right)$ is the distortion term. The mean square model error caused by bit errors is given as
+$\mathbb{E}[||\mathbf{\tilde{w}}-\mathbf{w}||^{2}|\mathbf{w}]=\frac{D\left(4^{N}-1\right)}{3\left(2^{N}-1\right)^{2}}\cdot b(1-b)^{N-1}M_{\mathbf{w}}^{2},$ (7)
+where $b$ is the BER and $M_{\mathbf{w}}\triangleq\max(\mathbf{w})-\min(\mathbf{w})$.
 Proof.
-Please refer to Appendix V-A. ∎
-Figure 3: The difference between received noisy model and broadcasted perfect global model.
-(a) BER=$10^{-2}$
-(b) BER=$10^{-3}$
-(c) BER=$10^{-4}$
-Figure 4: Validation of the model error formula by the experiment of vanilla-CNN on Fashion-MNIST dataset.
-The experimental setting is in Section IV-A.
-In every communication round, we compute the actual model error between the noisy DNN model and perfect model, and compare with the value from the analytical function in (7).
-It is shown that the value from analytical function is very close to the actual value, for different BER levels and model range in different communication rounds.
-After understanding the effect of BER, we analyze the convergence property of wireless FL with communication errors.
+Please refer to Appendix A. ∎
+Figure 3: Distortion between the decoded and broadcast global model. Different error bits, e.g., the most significant bit (MSB) and least significant bit (LSB), lead to different model distortion.
+(a) $\text{BER}=10^{-2}$.
+(b) $\text{BER}=10^{-3}$.
+(c) $\text{BER}=10^{-4}$.
+Figure 4: Validation of the model error analysis in (7) by experiments on the Fashion-MNIST dataset.
+The experimental setting is detailed in Section IV-A.
+In each communication round, we compute the actual model error between the decoded and broadcast global model on a selected client, and compare with the value calculated by the RHS in (7) with $\mathbf{w}$ setting as $\mathbf{w}_{r}$.
+It is observed that the analytical and simulation results match closely for different BER conditions.
+The model errors also show a trend of boundness, which demonstrates the rationality of Assumption 6.
+Note that model distortion caused by bit errors $\mathbb{E}\left[\delta(w_{d})|\mathbf{w}\right]$ does not necessarily equal zero.
+For example, any bit error of an all-zero sequence always leads to positive model distortion.
+Such non-zero bias, which was overlooked in a related analysis [34], is carefully tackled in the convergence analysis in this paper.
+After understanding the effect of BER, we then analyze the convergence property of wireless FL with communication error, for which, the following boundness assumption on the global model parameters [14] is needed.
+The rationality of this assumption can be evident in Fig.
+4.
+From this figure, it can be observed that the model error is bounded under various BER conditions, and thus $M_{\mathbf{w}}$ is also bounded according to (7).
+Assumption 6.
+(Boundness of Model Parameters [14]) For every communication round $r=0,1,\cdots,R-1$, $M_{r}\triangleq\max(\mathbf{w}_{r})-\min(\mathbf{w}_{r})\leq M$.
 Lemma 2.
-With Assumptions 1-5 and the learning rate $\eta$ satisfies
-$1-\frac{n+L}{n}L\eta-2L^{2}\eta^{2}E(E-1)\geq 0,$ (8)
+With Assumptions 1-6 and if the learning rate $\eta$ satisfies
+$1-\frac{K+L}{K}L\eta-2L^{2}\eta^{2}E(E-1)\geq 0,$ (8)
 the average expected gradient norm of $\{\nabla f(\mathbf{\bar{w}}_{r,e})\}$ is bounded as follows:
-$\begin{split}&\frac{1}{RE}\sum_{r=0}^{R-1}\sum_{e=0}^{E-1}\mathbb{E}\|\nabla f(\mathbf{\bar{w}}_{r,e})\|^{2}\leq\frac{2(\mathbb{E}f(\mathbf{w}_{0})-f^{*})}{\eta RE}\\ &+\frac{L(L+2)d\hat{M}^{2}(4^{N}-1)}{3(2^{N}-1)^{2}\eta RE}\sum_{r=0}^{R-1}b_{r}(1-b_{r})^{N-1}\\ &+\frac{d\hat{M}^{2}}{\eta RE}\sum_{r=0}^{R-1}(b_{r}(1-b_{r})^{N-1})^{2}\\ &+\frac{L(L+1)\eta(\sigma_{L}^{2}+\sigma_{G}^{2})}{n}+\frac{L^{2}\eta^{2}\sigma_{L}^{2}(n+1)(E-1)}{n},\end{split}$ (9)
-where $f^{*}$ is the minimum global training loss, $\hat{M}$ is the maximum range of DNN weights over all communication rounds, $b_{r}$ is the BER in the $r$-th communication round, and $\mathbf{w}_{0}$ is the initial global model.
+$\begin{split}&\frac{1}{RE}\sum_{r=0}^{R-1}\sum_{e=0}^{E-1}\mathbb{E}\|\nabla f(\mathbf{\bar{w}}_{r,e})\|^{2}\leq\frac{2(\mathbb{E}f(\mathbf{w}_{0})-f^{*})}{\eta RE}\\ &+\frac{(L+1)^{2}DM^{2}(4^{N}-1)}{3(2^{N}-1)^{2}\eta RE}\sum_{r=0}^{R-1}b_{r}(1-b_{r})^{N-1}\\ &+\frac{L(L+1)\eta(\sigma_{L}^{2}+\sigma_{G}^{2})}{K}+\frac{L^{2}\eta^{2}\sigma_{L}^{2}(K+1)(E-1)}{K},\end{split}$ (9)
+where $f^{*}$ is the minimum global training loss, $M$ is defined in Assumption 6, $b_{r}$ is the BER in the $r$-th communication round, and $\mathbf{w}_{0}$ is the initial global model.
 Proof.
-Please refer to Appendix V-B. ∎
+Please refer to Appendix B. ∎
 Remark 1.
 It is worthy noting that the first term on the RHS of (9) captures the distance between the initial and minimum global training loss.
-The second and third terms indicates how bit errors in each communication round affect learning convergence.
+The second term indicates how bit errors in each communication round affect convergence.
 The last two terms stem from non-IID data among clients and the random mini-batch sampling.
-We next simplify (9) by applying $\eta=\frac{1}{L\sqrt{RE}}=\frac{1}{L\sqrt{T}}$, which satisfies (8) with a sufficient large value of $R$.
+We next simplify (9) by applying $\eta=\frac{1}{L\sqrt{RE}}=\frac{1}{L\sqrt{T}}$, which can satisfy (8) with a sufficiently large value of $R$.
 Theorem 1.
-With Assumptions 1-5 and a sufficient large value of $R$ such that $\eta=\frac{1}{L\sqrt{RE}}$ satisfies (8), (9) can be simplified as follows:
-$\begin{split}\frac{1}{T}&\sum_{r=0}^{R-1}\sum_{e=0}^{E-1}\mathbb{E}\|\nabla f(\mathbf{\bar{w}}_{r,e})\|^{2}\\ \leq&\frac{C_{0}}{\sqrt{T}}+\frac{C_{1}\sum_{r=0}^{R-1}b_{r}(1-b_{r})^{N-1}}{\sqrt{T}}\\ &+\frac{C_{2}\sum_{r=0}^{R-1}(b_{r}(1-b_{r})^{N-1})^{2}}{\sqrt{T}}+\frac{C_{3}}{\sqrt{T}}+\frac{C_{4}}{T},\end{split}$ (10)
-where $C_{0}\triangleq 2L(\mathbb{E}f(\mathbf{w}_{0})-f^{*})$, $C_{1}\triangleq\frac{L^{2}(L+2)dM^{2}(4^{N}-1)}{3(2^{N}-1)^{2}}$, $C_{2}\triangleq LdM^{2}$, $C_{3}\triangleq\frac{(L+1)(\sigma_{L}^{2}+\sigma_{G}^{2})}{n}$, $C_{4}\triangleq\frac{L^{2}\sigma_{L}^{2}(n+1)(E-1)}{n}$.
+With Assumptions 1-6 and a sufficiently large value of $R$ such that $\eta=\frac{1}{L\sqrt{RE}}$ satisfies (8), (9) can be simplified as follows:
+$\begin{split}\frac{1}{T}&\sum_{r=0}^{R-1}\sum_{e=0}^{E-1}\mathbb{E}\|\nabla f(\mathbf{\bar{w}}_{r,e})\|^{2}\\ \leq&\frac{C_{0}}{\sqrt{T}}+\frac{C_{1}\sum_{r=0}^{R-1}b_{r}(1-b_{r})^{N-1}}{\sqrt{T}}+\frac{C_{2}}{\sqrt{T}}+\frac{C_{3}}{T},\end{split}$ (10)
+where $C_{0}\triangleq 2L(\mathbb{E}f(\mathbf{w}_{0})-f^{*})$, $C_{1}\triangleq\frac{L(L+1)^{2}DM^{2}(4^{N}-1)}{3(2^{N}-1)^{2}}$, $C_{2}\triangleq\frac{(L+1)(\sigma_{L}^{2}+\sigma_{G}^{2})}{K}$, $C_{3}\triangleq\frac{L^{2}\sigma_{L}^{2}(K+1)(E-1)}{K}$.
 According to Theorem 1, in order to guarantee the convergence rate of $\mathcal{O}(1/\sqrt{T})$, it is crucial to control the BER in each communication round such that the RHS of (10) reduces to zero at a rate of $\mathcal{O}(1/\sqrt{T})$.
-Since the explicit relationship between $b_{r}$ and $r$ is difficult to obtain, we propose a simple BER regulation method in Corollary 1.
+Since the explicit relationship between $b_{r}$ and $r$ is difficult to obtain, we propose a simple yet practical BER regulation method in Corollary 1.
 As such, wireless FL systems can achieve the same convergence rate as those with error-free communication.
@@
 $b_{r}\propto\frac{1}{(r+1)^{2}},r=0,\cdots,R-1,$ (11)
-the RHS of (10) diminishes to zero at a rate of $\mathcal{O}(1/\sqrt{T})$, i.e., wireless FL converges to a local optimal solution of the global training loss funcation at a rate of $\mathcal{O}(1/\sqrt{T})$.
+the RHS of (10) diminishes to zero at a rate of $\mathcal{O}(1/\sqrt{T})$, i.e., wireless FL converges to a local optimal solution of the global training loss function at a rate of $\mathcal{O}(1/\sqrt{T})$.
 Proof.
-When $b_{r}\propto\frac{1}{(r+1)^{2}}\triangleq\frac{\vartheta}{(r+1)^{2}}$, to prove the RHS of (10) diminishes to zero at a rate of $\mathcal{O}(1/\sqrt{T})$, it is sufficient to verify
-$\lim_{T\rightarrow\infty}\frac{\sum_{r=0}^{R-1}b_{r}(1-b_{r})^{N-1}}{\sqrt{T}}=0$ (12)
-due to the fact that $0\leq\left(b_{r}(1-b_{r})^{N-1}\right)^{2}\leq b_{r}(1-b_{r})^{N-1}\leq 1$ for $0\leq b_{r}\leq 1$. Specifically, since
+Denote $b_{r}\propto\frac{1}{(r+1)^{2}}\triangleq\frac{\vartheta}{(r+1)^{2}}$. To prove the RHS of (10) diminishes to zero at a rate of $\mathcal{O}(1/\sqrt{T})$, it is sufficient to verify
+$\lim_{T\rightarrow\infty}\frac{\sum_{r=0}^{R-1}b_{r}(1-b_{r})^{N-1}}{\sqrt{T}}=0.$ (12)
+Specifically, since
 $\begin{split}0&\leq\frac{\sum_{r=0}^{R-1}b_{r}(1-b_{r})^{N-1}}{\sqrt{T}}\leq\frac{\sum_{r=0}^{R-1}b_{r}}{\sqrt{T}}=\vartheta\frac{\sum_{r=0}^{R-1}1/(r+1)^{2}}{\sqrt{T}}\\ &=\frac{\vartheta}{\sqrt{T}}\left(1+\sum_{r=1}^{R-1}\frac{1}{(r+1)^{2}}\right)\\ &<\frac{\vartheta}{\sqrt{T}}\left(1+\sum_{r=1}^{R-1}\frac{1}{r(r+1)}\right)=\frac{\vartheta}{\sqrt{T}}\left(2-\frac{1}{R}\right)\\ &=\frac{\vartheta}{\sqrt{T}}\left(2-\frac{E}{T}\right),\end{split}$ (13)
@@
 Corollary 1 indicates that gradually decreasing the BER at a rate of $\mathcal{O}(1/(r+1)^{2})$ can guarantee wireless FL converges at a rate of $\mathcal{O}(1/\sqrt{T})$.
-The proposed BER regulation rule achieves an improving communication quality with communication rounds.
-This is consistent with the solution developed in [36], which improves the communication quality by increasing the signal-to-noise ratio (SNR) with communication round.
-III-B BER vs. Number of Iterations
-The analysis in Section III-A indicates that the convergence rate of wireless FL with error-free communication can be achieved if the BER is properly controlled.
-In this subsection, we analyze the relationship between BER and the maximum number of LDPC decoding iterations, based on which, we develop an adaptive decoding scheme to maintain the same convergence rate.
-1 Input: The initial global model $\mathbf{w}_{0}$
-2 Output: The converged global model $\mathbf{w}_{R}$
+In other words, the proposed BER regulation rule achieves an improving communication quality over communication rounds.
+This is consistent with the solution developed in [29] for wireless FL on strongly convex loss functions, which improves the communication quality by increasing the signal-to-noise ratio (SNR) with communication rounds.
+III-B BER vs. Number of LDPC Decoding Iterations
+The analysis in Section III-A indicates that the convergence rate of conventional FL with error-free communication can still be achieved in wireless FL if the BER is properly controlled.
+In this subsection, we analyze the relationship between BER and the maximum number of LDPC decoding iterations, based on which, we develop an adaptive decoding scheme to maintain the desired convergence rate.
+1 Input: The initial global model $\mathbf{w}_{0}$, $R>1$, $b_{0}$, and $b_{R-1}$.
 2 for  Communication round $r=0,\cdots,R-1$ do
-4       for Client index $k=1,\cdots,n$ do
-5             $b_{r}\propto(1/(r+1)^{2})$ with $b_{0},b_{R-1}$
-6             $Q=f^{-1}(b_{r})$
-7             for $decoding.iter\leq Q$ do
-8                  one iteration of LDPC decoding
-9                   if Early termination is satisfied then
-10                        break
+3       The server broadcasts the $N$-bit representation of the global model $\mathbf{w}_{r}$ and scalars $\min(\mathbf{w}_{r}),\max(\mathbf{w}_{r})$.
+4       for Client $k=1,\cdots,K$ do
+5             Compute $b_{r}=\frac{(b_{0}-b_{R-1})R^{2}}{(R^{2}-1)(r+1)^{2}}+\frac{b_{R-1}R^{2}-b_{0}}{R^{2}-1}$.
+6             Find $Q_{r}$ by its mapping with $b_{r}$.
+7             Initialize $decoding.iter=0$.
+8             for $decoding.iter<Q_{r}$ do
+9                  Execute one iteration of the LDPC decoding algorithm for each frame that contains the global model parameters.
+10                   if Parity check passes then
+11                        Early termination and break.
 12                   end if
@@
 14             end for
-14            output: $\mathbf{\tilde{w}}_{r}$
+15            Reconstruct $\mathbf{\tilde{w}}_{r}^{k}$ according to the LDPC decoding results and $\max(\mathbf{w}_{r}),\min(\mathbf{w}_{r})$.
+16             Set $\mathbf{w}^{k}_{r,0}=\mathbf{\tilde{w}}_{r}^{k}$.
 17             for local training step $e=0,...,E-1$ do
-16                  $\mathbf{w}^{k}_{r,0}=\mathbf{\tilde{w}}_{r}$
 18                   $\mathbf{w}^{k}_{r,e+1}=\mathbf{w}^{k}_{r,e}-\eta\nabla f_{k}(\mathbf{w}^{k}_{r,e};\xi^{k}_{r,e})$
 19             end for
-19            $\Delta\mathbf{w}_{r}^{k}=\mathbf{w}^{k}_{r,E}-\mathbf{w}_{r,0}^{k}$
+20            Upload $\Delta\mathbf{w}_{r}^{k}=\mathbf{w}^{k}_{r,E}-\mathbf{w}_{r,0}^{k}$ to the server.
 21       end for
-21      $\mathbf{w}_{r+1}=\mathbf{w}_{r}+\frac{1}{n}\sum^{n}_{k=1}\Delta\mathbf{w}_{r}^{k}$
+22      Perform model aggregation at the server as $\mathbf{w}_{r+1}=\mathbf{w}_{r}+\frac{1}{K}\sum^{K}_{k=1}\Delta\mathbf{w}_{r}^{k}$.
 23 end for
-Algorithm 1 Regulating LDPC decoder in FL
-It has been shown in [41] that for message-passing-based channel decoders to achieve a fraction of 1-$\epsilon$ of the channel capacity over a binary erasure channel, the relationship between the decoding complexity (denoted as $\chi_{D}$) and BER (denoted as $b$) can be modeled as follows:
+Output: $\mathbf{\bar{w}}_{R,E}=\frac{1}{K}\sum_{k=1}^{K}\mathbf{w}_{R,E}^{k}$ as the learned model.
+Algorithm 1 Wireless FL with Adaptive LDPC Decoding
+It has been shown in [35] that for message-passing-based channel decoders, to achieve a fraction of 1-$\epsilon$ of the channel capacity over a binary erasure channel, the relationship between the decoding complexity (denoted as $\chi_{D}$) and BER (denoted as $b$) can be modeled as follows:
 $\chi_{D}=\mathcal{O}\left(\log\frac{1}{b}\right)+\mathcal{O}\left(\frac{1}{\epsilon}\log\frac{1}{\epsilon}\right).$ (14)
@@
 $Q=\frac{1}{\chi_{0}}\mathcal{O}\left(\log\frac{1}{b}\right)+\frac{1}{\chi_{0}}\mathcal{O}\left(\frac{1}{\epsilon}\log\frac{1}{\epsilon}\right).$ (16)
-In the next subsection, we develop an adaptive scheme to control the maximum number of decoding iterations for clients in a wireless FL system based on Corollary 1 and (16).
-III-C Decoding Iteration Control
+In the next subsection, we develop an adaptive scheme to control the maximum number of LDPC decoding iterations for clients in a wireless FL system based on Corollary 1 and (16).
+III-C LDPC Decoding Iteration Control
 In this subsection, we show how to adaptively control the maximum number of LDPC decoding iterations over communication rounds to achieve the desired BER.
 In principle, by combining (11) and (16), we obtain the following relationship between the maximum number of LDPC decoding iterations and communication rounds:
-$Q_{r}=\frac{1}{\chi_{0}}\mathcal{O}\left(2\log\ r\right)+\frac{1}{\chi_{0}}\mathcal{O}\left(\frac{1}{\epsilon}\log\frac{1}{\epsilon}\right),$ (17)
-which provides a guideline to select the maximum number of LDPC decoding iterations in the $r$-th communication round.
-Specifically, $Q_{r}$ can be set as $\alpha\log r+\beta$, where $\alpha$ and $\beta$ are constants.
+$Q_{r}=\frac{1}{\chi_{0}}\mathcal{O}\left(2\log\ (r+1)\right)+\frac{1}{\chi_{0}}\mathcal{O}\left(\frac{1}{\epsilon}\log\frac{1}{\epsilon}\right),$ (17)
+which provides a theoretical guideline to determine the maximum number of LDPC decoding iterations in the $r$-th communication round.
+Specifically, $Q_{r}$ can be set as $\alpha\log(r+1)+\beta$, where $\alpha$ and $\beta$ are constants.
 However, (17) cannot be directly utilized because $\alpha$ and $\beta$ are code-specific and depend on the receive SNR as well as channel decoding algorithms.
-Thus, we propose a practical approach by the mapping relationship which captures the relationship between BER and the maximum number of decoding iterations at different levels of receive SNR.
-Entries in the mapping relationship can be obtained through empirical measurements specific to the LDPC code used in the wireless FL system.
+Thus, we propose a practical approach via the mapping between BER and the maximum number of LDPC decoding iterations at different levels of receive SNR.
+Entries in the mapping can be obtained through empirical measurements specific to the LDPC code used in the wireless FL system.
 Fig.
-5 shows the exampled simulation results for a rate 1/2 LDPC code with a block length of 1008 bits and binary phase shift keying (BPSK) modulation over an additional white Gaussian noise (AWGN) channel with $\text{SNR}=1.5$ and 2.5 dB.
-Data in this figure can be stored in memory [43].
-From the mapping relationship, we can get the corresponding maximum number of decoding iteration for a target BER for clients with different receive SNR.
-Figure 5: The relation between BER and LDPC maximum decoding iterations under different SNRs.
+5 shows the sample simulation results for the rate 1/2 LDPC code with a block length of 1008 bits and binary phase shift keying (BPSK) modulation over an additive white Gaussian noise (AWGN) channel with SNR = 1.5 and 2.5 dB, which can be stored in memory of clients [37].
+From the mapping, we are able to obtain the corresponding maximum number of decoding iteration for a target BER for clients with different receive SNR in each communication round.
+Figure 5: Mappings between BER and the maximum number of LDPC decoding iterations at different receive SNRs.
 The wireless FL procedures with adaptive LDPC decoding are summarized in Algorithm 1.
-The algorithm starts with an initial target BER, denoted as $b_{0}$, which is set to a relatively high value, i.e., $10^{-1}$ in our experiments.
-The target BER at the last communication round, denoted as $b_{R-1}$, is set as a small value, i.e., $10^{-4}$ in our experiments.
-The concrete expression of $b_{r}$ in our experiment is $b_{r}=\frac{0.0999R^{2}}{(R^{2}-1)(r+1)^{2}}+\frac{10^{-4}R^{2}-0.1}{R^{2}-1}$.
-For each communication round, the BER is scaled according to (11), and the corresponding maximum number of LDPC decoding iterations in that round is obtained from the mapping relationship.
+The algorithm starts with an initial BER, denoted as $b_{0}$, and gradually decreases the BER with communication rounds at the rate of $\frac{1}{(r+1)^{2}}$.
+In practice, since the BER of LDPC decoding cannot be arbitrarily small (see Fig.
+5), we set a target BER at the last communication round, denoted as $b_{R-1}$.
+With $b_{0}$, $b_{R-1}$ ($b_{0}>b_{R-1}$), and the desired BER should scale at the rate of $\frac{1}{(r+1)^{2}}$, $b_{r}$ can be designed as
+$b_{r}=\frac{(b_{0}-b_{R-1})R^{2}}{(R^{2}-1)(r+1)^{2}}+\frac{b_{R-1}R^{2}-b_{0}}{R^{2}-1}$ (18)
+for $R>1$. The corresponding maximum number of LDPC decoding iterations in that round is obtained from the measured mapping relationship.
 Remark 2.
-Note that the number of entries in the mapping relationship depends on the possible discrete BER and SNR values.
-Since this number is not large, the memory overhead for storing the mapping relationship is expected to be small.
-Additionally, reading data from the memory takes constant time and does not introduce additional LDPC decoding latency.
+Note that the number of entries in the mappings depend on the granularity of discrete BER and SNR values.
+Since this number is not large, the overhead for storing them is expected to be small.
+Additionally, reading these entries from memory takes constant time and does not introduce additional LDPC decoding latency.
 IV Experimental Results
@@
 First, we introduce the experimental settings.
-Then, the performance of the proposed algorithm, including test accuracy and energy consumption, is compared with a baseline with a fixed number of LDPC decoding iterations.
-Figure 6: Fixed decoding schemes: On the client side, setting the LDPC maximum decoding iterations to 6, 12, 18, 24, and 52 can achieve a BER of around $10^{-1}$, $10^{-2}$, $10^{-3}$, $10^{-4}$, and $10^{-5}$, respectively.
+Then, performance of the proposed algorithm, including test accuracy and decoding energy consumption, is compared with a baseline by fixing the maximum number of LDPC decoding iterations in each communication round (hereinafter referred to as the “Fixed-$Q$” scheme).
+Figure 6: BER of a sample client achieved by the baseline scheme (“Fixed-$Q$”) with a fixed maximum number of LDPC decoding iterations: By setting the maximum number of LDPC decoding iterations $Q$ to 6, 12, 18, 24, and 52, clients achieve the BER at around $10^{-1}$, $10^{-2}$, $10^{-3}$, $10^{-4}$, and $10^{-5}$, respectively.
 IV-A Experimental Settings
-(a) Experiment of LeNet-300-100 on MNIST.
-(b) Experiment of vanilla-CNN on Fashion-MNIST.
-(c) Experiment of CNN on CIFAR-10.
-(d) Experiment of vanilla-CNN on non-IID Fashion-MNIST.
-Figure 7: Experiments on different datasets showing how fixed BER affects FL accuracy.
-(a) Consumed decoding iteration for each frame.
-(b) The distributed for different consumed iterations.
-Figure 8: Example of vanilla-CNN on Fashion-MNIST. The 8-bit quantized vanilla CNN model goes through AWGN channel with $E_{b}/N_{0}$ of 2.5dB, and is decoded by LDPC of block-length of 1008 bit.
-System Setups: We consider an FL system with 10 clients.
-All clients are assumed to be selected in every communication round and they communicate with the server using rate 1/2 min-sum LDPC code with a code length of 1008 bits[44].
-Besides, binary phase shift keying (BPSK) is adopted as the modulation scheme.
-In each communication round, the number of local SGD steps is set as 5 with the learning rate and batch size as 0.01 and 64, respectively.
-The receive SNR of all the wireless channels is assumed to be 2.5 dB.
-Datasets: We conduct experiments on three dataset, including MNIST[45], Fashion-MNIST[46], and CIFAR-10[47].
+(a) LeNet-300-100 on MNIST (IID).
+(b) Vanilla-CNN on Fashion-MNIST (IID).
+(c) 7-layers CNN on CIFAR-10 (IID).
+(d) Vanilla-CNN on Fashion-MNIST (non-IID).
+Figure 7: Test accuracy on different datasets under the baseline scheme “Fixed-$Q$” with a fixed maximum number of LDPC decoding iterations.
+(a) Actual number of LDPC decoding iterations.
+(b) Distribution of the actual number of LDPC decoding iterations.
+Figure 8: Actual numbers of LDPC decoding iterations and its distribution at a sample client in a selected communication round on Fashion-MNIST (IID).
+A 8-bit vanilla CNN model is encoded by the rate 1/2 min-sum LDPC code of block-length of 1008 bit, resulting in 4620 frames in a communication round.
+System Setup: We consider a wireless FL system with 10 clients.
+The clients communicate with the server via the rate 1/2 min-sum LDPC code using a code length of 1008 bits[38].
+BPSK is adopted as the modulation scheme.
+In each communication round, $N=8$ bits are used to digitalize each model parameter.
+The number of local training steps in each communication round is set to $E=5$ with the learning rate and batch size as 0.01 and 64, respectively.
+The receive SNR of all the downlink wireless channels is assumed to be 2.5 dB.
+Datasets: We conduct experiments on three classic datasets, including MNIST[39], Fashion-MNIST[40], and CIFAR-10[41].
 The training datasets are distributed among clients in an IID manner.
 We further consider one experiment where the Fashion-MNIST data are non-IID among clients.
-In particular, each client was assigned with two categories of training data, similar to the setup in [24].
-The test datasets are used for validation at the server.
+In particular, each client is assigned with two categories of training data similar to the setup in [42].
+The test datasets are used for model accuracy evaluation at the server.
 DNN Models: For the MNIST dataset, the LeNet-300-100 model[43] is employed, which is a fully connected neural network consisting of two hidden layers and one softmax layer, with approximately 0.27 million parameters.
-For the Fashion-MNIST dataset, a vanilla CNN [3] with two 5×5 convolution layers, one fully connected layer, and one softmax layer is employed, comprising around 0.58 million parameters.
+For the Fashion-MNIST dataset, a vanilla convolutional neural network (CNN) [4] with two 5×5 convolution layers, one fully connected layer, and one softmax layer is employed, comprising around 0.58 million parameters.
 For the CIFAR-10 dataset, a 7-layer CNN model with four 3×3 convolution layers, two fully connected layers, and one softmax layer is employed, which contains approximately 9.07 million parameters.
-IV-B Learning Performance of Conventional Schemes
-(a) Max.iter.
-(b) Learning curves.
-(c) Average iterations.
-(d) BER.
-(e) Final learning accuracy.
-(f) Total decoding energy.
-Figure 9: Experiment for LeNet-300-100 on MNIST dataset.
-Compared with the conventional fixed schemes, our method that adaptively increases the maximum decoding iteration to 24 can achieve the same final accuracy of 98%, and reduce the decoding energy by 6.7mJ.
-(a) Max.iter.
-(b) Learning curves.
-(c) Average iterations.
-(d) BER.
-(e) Final learning accuracy.
-(f) Total decoding energy.
-Figure 10: Experiment for vanilla CNN on IID Fashion-MNIST dataset.
-Compared with the conventional fixed schemes, our method that adaptively increases the maximum decoding iteration to 24 can achieve a similar final accuracy of 91%, and reduce the decoding energy by 14.1mJ.
-In this subsection, experimental results for the baseline scheme are presented. As shown in Fig. 6, it leads to approximately constant BERs over the training process.
-Next, we investigate the impact of a constant BER on the performance of wireless FL.
+IV-B Learning Performance of the Baseline Scheme
+(a) $Q_{r}$.
+(b) Test accuracy.
+(c) Avg. actual number of decoding iterations among clients.
+(d) BER at every communication round,
+i.e., ($b_{r}$).
+(e) Test accuracy at the last communication round.
+(f) Total decoding energy consumption.
+Figure 9: Experimental results on MNIST (IID).
+Compared with the baseline scheme, our proposed algorithm that adaptively increases the maximum number of LDPC decoding iterations to 24 after 50 communication rounds can achieve the same final test accuracy of 98% and reduce the decoding energy by 6.7 mJ.
+(a) $Q_{r}$.
+(b) Test accuracy.
+(c) Avg. actual number of decoding iterations among clients.
+(d) BER at every communication round,
+i.e., ($b_{r}$).
+(e) Test accuracy at the last communication round.
+(f) Total decoding energy consumption.
+Figure 10: Experimental results on Fashion-MNIST (IID).
+Compared with the baseline scheme, our proposed algorithm that adaptively increases the maximum number of LDPC decoding iterations to 24 after 50 communication rounds can achieve a similar final test accuracy of 91% and reduce the decoding energy by 14 mJ.
+In this subsection, experimental results for the baseline scheme “Fixed-$Q$” are presented.
+First, as shown in Fig.
+6, a fixed maximum number of LDPC decoding iterations leads to approximately constant BERs over the training process.
+Next, we investigate the impact of a roughly constant BER on the learning performance of wireless FL.
 Fig.
-7 summarizes the test accuracy with different maximum numbers of decoding iterations.
-It can be observed that for all the four FL tasks, when the BER is higher than $10^{-3}$, the learning performance was significantly impaired.
-Lowering the BER to $10^{-4}$ leads to improved accuracy and faster convergence towards wireless FL with error-free communication.
-However, further decreasing the BER to $10^{-5}$ does not yield any significant improvement in test accuracy.
-IV-C Learning Performance of the Proposed Scheme
-(a) Max.iter.
-(b) Learning curves.
-(c) Average iterations.
-(d) BER.
-(e) Final learning accuracy.
-(f) Total decoding energy.
-Figure 11: Experiment for 7-layer CNN on CIFAR-10 dataset.
-Compared with the conventional fixed schemes, our method that adaptively increases the maximum decoding iteration to 24 can achieve a similar final accuracy of 77.6%, and reduce the decoding energy by 190mJ.
-(a) Max.iter.
-(b) Learning curves.
-(c) Average iterations.
-(d) BER.
-(e) Final learning accuracy.
-(f) Total decoding energy.
-Figure 12: Experiment for vanilla CNN on non-IID Fashion-MNIST dataset.
-Compared with the conventional fixed schemes, our method that adaptively increases the maximum decoding iteration to 24 can achieve a similar final accuracy of 54.1%, and reduce the decoding energy by 13.9mJ.
-In this subsection, the proposed adaptive LDPC decoding scheme is compared to the baseline scheme.
-Specifically, the learning process of the proposed scheme is assigned with an initial BER of $10^{-1}$ and a final BER value of $10^{-4}$.
-The desired BER at every communication round is calculated via (11) and the maximum number of LDPC decoding iterations is determined according to the mapping with BER as stated in Section III.
-In the experiments, the initial and final maximum numbers of LDPC decoding iterations are thus set to 6 and 24, respectively.
-For the baseline scheme, a maximum of 24 LDPC decoding iterations are used in each communication round, which corresponds to a BER of $10^{-4}$.
-Fig.
-8(a) shows an example on the Fashion-MNIST (IID) datasets to illustrate the actual number of decoding iterations for each frame of data bits.
-It can observed that only a few frames exhaust all the decoding iterations which is set to a large value of 52.
-Note that the decoding energy consumption is proportional to the actual number of decoding iterations used for the FL tasks.
-From the distribution on the actual number of decoding iterations shown in Fig.
-8(b), it is observed that 18 decoding iterations are sufficient for most cases, implying the significant reduction of energy consumption.
-Fig.
-9 summarizes the experimental results for LeNet-300-100 on the MNIST (IID) dataset.
-Specifically, Fig.
-9(a) shows how the maximum number of LDPC decoding iterations varies with communication rounds in the proposed scheme.
-While the baseline scheme with a fixed number of maximum decoding iterations results in a constant BER, the proposed scheme achieves the desired BER by adopting a linearly increasing maximum number of decoding iterations.
-Fig.
-9(b) shows the test accuracy achieved by different schemes.
-It is observed that setting the maximum number of decoding iterations to 6 (which corresponds to a BER of $\sim 10^{-1}$) severely affects the learning accuracy.
-Similarly, when the maximum number of decoding iterations is set to 12 (which corresponds to a BER of $\sim 10^{-2}$), the training process still fails to converge.
+7 summarizes the test accuracy with different maximum numbers of LDPC decoding iterations.
+It can be observed that for all the four FL tasks, setting the maximum number of decoding iterations to 6 (which corresponds to a BER of $\sim 10^{-1}$) severely affects the learning accuracy.
+Similarly, when the maximum number of decoding iterations is set to 12 (which corresponds to a BER of $\sim 10^{-2}$), the training process fails to converge to satisfactory accuracy.
 By setting the maximum number of decoding iterations as 18 (which corresponds to a BER of $\sim 10^{-3}$), the training process converges, but with a noticeable loss in test accuracy compared to the case when the maximum number of decoding iterations is 24 (which corresponds to a BER of $\sim 10^{-4}$).
 However, further increasing the maximum number of decoding iterations to 52 does not lead to much accuracy improvement.
-The proposed scheme, which linearly increases the maximum number of decoding iterations from 6 to 24, achieves the same test accuracy as those with fixed 24 or 52 decoding iterations.
+Thus, we next focus on the most efficient configuration of the baseline scheme with $Q=24$, i.e., “Fixed-24”.
+IV-C Learning Performance of the Proposed Scheme
+(a) $Q_{r}$.
+(b) Test accuracy.
+(c) Avg. actual number of decoding iterations among clients.
+(d) BER at every communication round,
+i.e., ($b_{r}$).
+(e) Test accuracy at the last communication round.
+(f) Total decoding energy consumption.
+Figure 11: Experimental results on CIFAR-10 (IID).
+Compared with the baseline scheme, our proposed algorithm that adaptively increases the maximum number of LDPC decoding iterations to 24 after 40 communication rounds can achieve a similar final test accuracy of 77.6% and reduce the decoding energy by 191 mJ.
+(a) $Q_{r}$.
+(b) Test accuracy.
+(c) Avg. actual number of decoding iterations among clients.
+(d) BER at every communication round,
+i.e., ($b_{r}$).
+(e) Test accuracy at the last communication round.
+(f) Total decoding energy consumption.
+Figure 12: Experimental results on Fashion-MNIST (non-IID).
+Compared with the baseline scheme, our proposed algorithm that adaptively increases the maximum number of LDPC decoding iterations to 24 after 50 communication rounds can achieve a similar final test accuracy of 54.1% and reduce the decoding energy by 13.7 mJ.
+In this subsection, the proposed adaptive LDPC decoding scheme is compared to the baseline scheme.
+Specifically, the learning process of the proposed scheme is assigned with an initial BER of $b_{0}=10^{-1}$ and a final BER of $b_{R-1}=10^{-4}$.
+The maximum number of LDPC decoding iterations is determined according to its mapping with BER as stated in Section III.
+In the experiments, the initial maximum numbers of LDPC decoding iterations are thus set to 6 and gradually increased to 24 at the last communication round.
+For the baseline “Fixed-24” scheme, a maximum of 24 LDPC decoding iterations can be used in each communication round, which corresponds to a BER of $10^{-4}$.
 Fig.
-9(c) illustrates how the average number of decoding iterations over all DNN parameters varies during the training process.
-The proposed scheme can reduce the average number of decoding iterations and thus the overall energy consumption.
+8(a) shows an example on the Fashion-MNIST (IID) dataset to illustrate the actual number of decoding iterations of digitalized model parameters at a sample client in a selected communication round.
+It can observed that only a few frames exhaust all the admissible LDPC decoding iterations, which is set to 52.
+Note that the decoding energy consumption is proportional to the actual number of decoding iterations.
+From the distribution of the actual number of LDPC decoding iterations shown in Fig.
+8(b), it is observed that 18 decoding iterations are sufficient for 80% of the frames, implying the significant reduction of decoding energy consumption.
 Fig.
-9(d) depicts the BER of the proposed scheme over communication rounds, which scales as $\mathcal{O}(1/(r+1)^{2})$ from $10^{-1}$ to $10^{-4}$.
+9 summarizes the experimental results on the MNIST (IID) dataset.
+Specifically, Fig.
+9(a) shows how the maximum number of LDPC decoding iterations ($Q_{r}$) varies with communication rounds in different schemes.
+While the baseline scheme “Fixed-24” results in a constant BER, the proposed scheme achieves the desired BER by linearly increasing the maximum number of LDPC decoding iterations.
+We note that this behavior results from the empirical mapping between $Q_{r}$ and $b_{r}$ for the used LDPC decoder, which is different from the general theoretical guideline in (17).
 Fig.
-9(e) shows the final test accuracy of different schemes at the 50-th communication round.
-It can be observed that the proposed scheme achieves the same accuracy as the baseline schemes that always maintain a very low BER of $10^{-4}$ and $10^{-5}$ throughout the training process.
+9(b) shows the test accuracy achieved by the proposed scheme and the baseline “Fixed-24” scheme.
+Compared with the baseline scheme with a fixed number of 24 decoding iterations, the proposed scheme achieves the same test accuracy with 50 communication rounds despite with slower convergence.
 Fig.
-9(f) illustrates the decoding energy consumption on clients under different schemes.
-It is obtained by multiplying the total number of LDPC decoding iterations with the LDPC decoding energy efficiency reported in [29].
-Compared to the baseline scheme with a maximum of 24 decoding iterations, the proposed scheme reduces the overall decoding energy consumption by 19.4%.
-TABLE III: Experimental Results of Different Schemes. Our scheme can achieve similar accuracy while saving around 20% of the decoding energy compared to the best-fixed scheme.
-Final learning accuracy
+9(c) illustrates how the average actual number of decoding iterations at clients over all DNN parameters varies during the training process.
+The proposed scheme can reduce the average number of decoding iterations and thus the overall decoding energy consumption.
+Fig.
+9(d) depicts the BER (i.e., $b_{r}$) of the proposed scheme over communication rounds, which scales as $\mathcal{O}(1/(r+1)^{2})$ from $10^{-1}$ to $10^{-4}$.
+Fig.
+9(e) shows the final test accuracy of different schemes at the 49-th communication round.
+It can be observed that the proposed scheme achieves the same accuracy as the baseline schemes (“Fixed-24”) that always maintains a very low BER of $10^{-4}$ throughout the training process.
+Fig.
+9(f) illustrates the overall decoding energy consumption at clients under different schemes, which is obtained by multiplying the total number of decoding iterations with the LDPC decoding energy efficiency reported in [23].
+Compared to the baseline scheme, the proposed scheme significantly reduces the overall decoding energy consumption by 19.4%.
+TABLE III: Summary of Experimental Results.
+Our scheme can achieve similar accuracy while saving around 20% of the overall decoding energy consumption compared to the best baseline (“Fixed-$Q$”) scheme.
+Final Test Accuracy
 Fixed-18 Fixed-24 Fixed-52 Ours
-MNIST 96.74% 97.96% 98.06% 98.02%
-Fashion-MNIST 86.18% 90.63% 91.24% 91.16%
-CIFAR-10 76.72% 77.38% 77.57% 77.68%
-non-IID Fashion 51.3% 54.1% 54.3% 54.2%
-Total consumed decoding energy (mJ)
+MNIST (IID) 96.74% 97.96% 98.06% 98.02%
+Fashion-MNIST (IID) 86.18% 90.63% 91.24% 91.16%
+CIFAR-10 (IID) 76.72% 77.38% 77.57% 77.68%
+Fashion-MNIST (non-IID) 51.3% 54.1% 54.3% 54.2%
+Total Decoding Energy Consumption (mJ)
 Fixed-18 Fixed-24 Fixed-52 Ours
-MNIST 33.87 34.51 34.68 27.81
-Fashion-MNIST 74.23 76.10 76.29 62.17
-CIFAR-10 925.10 948.44 951.36 757.80
-non-IID Fashion 74.22 76.11 76.27 62.38
+MNIST (IID) 33.87 34.51 34.68 27.81
+Fashion-MNIST (IID) 74.23 76.10 76.29 62.17
+CIFAR-10 (IID) 925.10 948.44 951.36 757.80
+Fashion-MNIST (non-IID) 74.22 76.11 76.27 62.38
 Figs.
-10, 11, and 12 present the experimental results for Fashion-MNIST (IID), CIFAR-10 (IID), and Fashion-MNIST (non-IID) datasets, respectively.
+10, 11, and 12 present the experimental results on Fashion-MNIST (IID), CIFAR-10 (IID), and Fashion-MNIST (non-IID) datasets, respectively.
 The results are similar to those on the MNIST (IID) dataset in Fig.
-9, showing that adapting the maximum number of LDPC decoding iterations can achieve similar or even slightly better accuracy compared to the baseline scheme.
+9, showing that adapting the maximum number of LDPC decoding iterations can achieve comparable or even slightly better accuracy compared to the baseline scheme.
 This observation was also reported in [44, 45, 46], which suggests that adding annealed noise to the DNN model could enhance its generalization ability.
 Moreover, the proposed scheme reduces the LDPC decoding energy consumption on the Fashion-MNIST (IID), CIFAR-10 (IID), and Fashion-MNIST (non-IID) datasets by 18.3%, 20.1%, and 18.0%, respectively.
-Table III provides a summary of the experimental data.
+Table III provides a summary of the experimental results.
 V Conclusion
@@
 $\tilde{\mathbf{w}}=\mathbf{w}+\bm{\delta}\left(\mathbf{w}\right),$ (19)
-where $\bm{\delta}\left(\mathbf{w}\right)$ is a $d$-dimensional function of $\mathbf{w}$ denoting the model distortion caused by bit errors.
+where $\bm{\delta}\left(\mathbf{w}\right)$ is a $D$-dimensional random function of $\mathbf{w}$ denoting the model distortion caused by bit errors.
 Under Assumption 4, each entry of $\bm{\delta}\left(\mathbf{w}\right)$ depends only on $w_{d}$, i.e., the $d$-th dimension of $\mathbf{w}$.
@@
 $\mathbb{E}\left[\tilde{w}_{d}|w_{d}\right]=w_{d}+b\left(1-b\right)^{N-1}\Delta\sum_{i=0}^{N-1}\left(1-2w_{d}^{(i)}\right)2^{i},$ (21)
-The expected bias of $w_{d}$ is bounded as follows:
-$\begin{split}|\mathbb{E}\left[\tilde{w}_{d}|w_{d}\right]-w_{d}|&\leq b\left(1-b\right)^{N-1}\left(2^{N}-1\right)\Delta\\ &=b\left(1-b\right)^{N-1}M.\end{split}$ (21)
-The mean square error of $w_{d}$ is given as follows:
-$\begin{split}\mathbb{E}\left[||\tilde{w}_{d}-w_{d}||^{2}|w_{d}\right]&=b\left(1-b\right)^{N-1}\sum_{i=0}^{N-1}\left(\Delta 2^{i}\right)^{2}\\ &=b\left(1-b\right)^{N-1}\frac{4^{N}-1}{3(2^{N}-1)^{2}}M^{2}.\end{split}$ (22)
+where $\Delta=\frac{M_{\mathbf{w}}}{2^{N-1}}$.
+Note that the second term on the RHS of (21) is the non-zero expected bias caused by random bit errors.
+Thus, the mean square error of $w_{d}$ is given as follows:
+$\begin{split}\mathbb{E}\left[||\tilde{w}_{d}-w_{d}||^{2}|w_{d}\right]&=b\left(1-b\right)^{N-1}\sum_{i=0}^{N-1}\left(\Delta 2^{i}\right)^{2}\\ &=b\left(1-b\right)^{N-1}\frac{4^{N}-1}{3(2^{N}-1)^{2}}M_{\mathbf{w}}^{2}.\end{split}$ (22)
 Therefore,
-$\begin{split}\mathbb{E}\left[||\mathbf{\tilde{w}}-\mathbf{w}||^{2}|\mathbf{w}\right]&=\sum_{j=1}^{d}\mathbb{E}\left[||\tilde{w}_{j}-w_{j}||^{2}|\mathbf{w}\right]\\ &=\frac{d(4^{N}-1)}{3(2^{N}-1)^{2}}\cdot b\left(1-b\right)^{N}M^{2}.\end{split}$ (23)
+$\begin{split}\mathbb{E}\left[||\bm{\delta}(\mathbf{w})||^{2}|\mathbf{w}\right]&=\mathbb{E}\left[||\mathbf{\tilde{w}}-\mathbf{w}||^{2}|\mathbf{w}\right]\\ &=\sum_{d=1}^{D}\mathbb{E}\left[||\tilde{w}_{d}-w_{d}||^{2}|\mathbf{w}\right]\\ &=\frac{D(4^{N}-1)}{3(2^{N}-1)^{2}}\cdot b\left(1-b\right)^{N}M_{\mathbf{w}}^{2}.\end{split}$ (23)
 V-B Proof of Lemma 2
-We define
-$\mathbf{w}_{r+1}$ $=\mathbf{w}_{r}+\frac{1}{n}\sum_{k=1}^{n}(\mathbf{w}_{r,E}^{k}-\mathbf{\tilde{w}}_{r}),$ (24)
-$\mathbf{\bar{w}}_{r,e}$ $=\frac{1}{n}\sum_{k=1}^{n}\mathbf{w}_{r,e}^{k},$ (25)
-for each communication round $r=0,...,R-1$ and each local iteration $e=0,\cdots,E-1$, where $\mathbf{w}_{r+1}$ denotes the updated global model and $\mathbf{\bar{w}}_{r,e}$ is the average of client models after the $e$-th local iteration.
+We rewrite (5) by substituting $\mathbf{w}_{r,0}^{k}=\tilde{\mathbf{w}}_{r}$ as
+$\mathbf{w}_{r+1}$ $=\mathbf{w}_{r}+\frac{1}{K}\sum_{k=1}^{K}(\mathbf{w}_{r,E}^{k}-\mathbf{\tilde{w}}_{r}).$ (24)
 To prove the Lemma 2, we first present Lemmas 25$\sim$5.
 Lemma 3.
-For $\mathbf{\bar{w}}_{r,E}$ and $\mathbf{w}_{r+1}$ defined in (24) and (25), we have
+Based on the definitions of $\mathbf{w}_{r+1}$ and $\mathbf{\bar{w}}_{r,E}$, we have
 $\mathbf{w}_{r+1}-\mathbf{\bar{w}}_{r,E}=\mathbf{w}_{r}-\mathbf{\tilde{w}}_{r},\ r=0,\cdots,R-1.$ (25)
 Proof.
-The proof can be obtained by substituting the definition of $\mathbf{\bar{w}}_{r,E}$ to the LHS of (26). ∎
+The proof is straightforward by substituting the definition of $\mathbf{\bar{w}}_{r,E}$ in (24) to the LHS of (25). ∎
 Lemma 25 translates the difference between $\mathbf{w}_{r+1}$ and $\mathbf{\bar{w}}_{r,E}$ to the model distortion incurred by downlink bit errors.
@@
 Proof.
-The proof can be obtained using the $L$-smoothness of $f(\cdot)$, which is omitted for brevity. ∎
+The proof can be obtained using the $L$-smoothness property of $f(\cdot)$, which is omitted for brevity. ∎
 Lemma 26 provides an upper bound on the expected value of the global loss function at $\mathbf{w}_{r+1}$.
@@
 With Assumptions 1 - 3, for $r=0,\cdots,R-1$,
-$\begin{split}&\mathbb{E}f(\mathbf{\bar{w}}_{r,E})\leq\mathbb{E}f(\mathbf{w}_{r})+\frac{L}{2}\mathbb{E}\|\mathbf{\tilde{w}}_{r}-\mathbf{w}_{r}\|^{2}+\mathbb{E}\left\langle\nabla f(\mathbf{w}_{r}),\bm{\delta}(\mathbf{w}_{r})\right\rangle\\ &-\frac{\eta}{2n}\left[1-L\eta-2L^{2}\eta^{2}E(E-1)\right]\sum_{k=1}^{n}\sum_{e=0}^{E-1}\mathbb{E}\|\nabla f(\mathbf{w}_{r,e}^{k})\|^{2}\\ &-\frac{\eta}{2}\sum_{e=0}^{E-1}\mathbb{E}\|\nabla f(\mathbf{\bar{w}}_{r,e})\|^{2}+\frac{LE\eta^{2}(\sigma_{L}^{2}+\sigma_{G}^{2})}{2n}\\ &+\frac{L^{2}\eta^{3}\sigma_{L}^{2}(n+1)E(E-1)}{2n}.\end{split}$ (28)
+$\begin{split}&\mathbb{E}f(\mathbf{\bar{w}}_{r,E})\leq\mathbb{E}f(\mathbf{w}_{r})+\frac{L}{2}\mathbb{E}\|\bm{\delta}(\mathbf{w}_{r})\|^{2}+\mathbb{E}\left\langle\nabla f(\mathbf{w}_{r}),\bm{\delta}(\mathbf{w}_{r})\right\rangle\\ &-\frac{\eta}{2K}\left[1-L\eta-2L^{2}\eta^{2}E(E-1)\right]\sum_{k=1}^{K}\sum_{e=0}^{E-1}\mathbb{E}\|\nabla f(\mathbf{w}_{r,e}^{k})\|^{2}\\ &-\frac{\eta}{2}\sum_{e=0}^{E-1}\mathbb{E}\|\nabla f(\mathbf{\bar{w}}_{r,e})\|^{2}+\frac{LE\eta^{2}(\sigma_{L}^{2}+\sigma_{G}^{2})}{2K}\\ &+\frac{L^{2}\eta^{3}\sigma_{L}^{2}(K+1)E(E-1)}{2K}.\end{split}$ (27)
 Proof.
 The proof can be obtained following similar lines for Lemma 7 in [16] by further considering the non-IID data distribution among clients (i.e., Assumption 3), which is omitted for brevity. ∎
-Lemma 28 provides an upper bound on the expected value of the global loss function at $\bar{\mathbf{w}}_{r,E}$ in terms of the expected model distortion caused by downlink bit errors, i.e., $\mathbb{E}\|\mathbf{\tilde{w}}_{r}-\mathbf{w}_{r}\|^{2}$.
-By applying Lemmas 26 and 28 to Lemma 27, we obtain the following inequality:
-$\begin{split}&\mathbb{E}f(\mathbf{w}_{r+1})\leq\mathbb{E}f(\mathbf{w}_{r})+L\mathbb{E}\|\mathbf{\tilde{w}}_{r}-\mathbf{w}_{r}\|^{2}\\ &+\mathbb{E}\left\langle\nabla f(\mathbf{w}_{r})-\nabla f(\mathbf{\bar{w}}_{r,E}),\bm{\delta}(\mathbf{w}_{r})\right\rangle-\frac{\eta}{2}\sum_{e=0}^{E-1}\mathbb{E}\|\nabla f(\mathbf{\bar{w}}_{r,e})\|^{2}\\ &-\frac{\eta}{2n}\left[1-L\eta-2L^{2}\eta^{2}E(E-1)\right]\sum_{k=1}^{n}\sum_{e=0}^{E-1}\mathbb{E}\|\nabla f(\mathbf{w}_{r,e}^{k})\|^{2}\\ &+\frac{LE\eta^{2}(\sigma_{L}^{2}+\sigma_{G}^{2})}{2n}+\frac{L^{2}\eta^{3}\sigma_{L}^{2}(n+1)E(E-1)}{2n}.\end{split}$ (29)
-By leveraging $2\left\langle\mathbf{a},\mathbf{b}\right\rangle=\|\mathbf{a}\|^{2}+\|\mathbf{b}\|^{2}-\|\mathbf{a}-\mathbf{b}\|^{2}\leq\|\mathbf{a}\|^{2}+\|\mathbf{b}\|^{2}$ holds for any $\mathbf{a}$, $\mathbf{b}$, an upper bound for the third term on the RHS of (29) is derived as follows:
-$\begin{split}&\mathbb{E}\left\langle\nabla f(\mathbf{w}_{r})-\nabla f(\mathbf{\bar{w}}_{r,E}),\bm{\delta}(\mathbf{w}_{r})\right\rangle\\ &\leq\frac{\mathbb{E}\|\nabla f(\mathbf{w}_{r})-\nabla f(\mathbf{\bar{w}}_{r,E})\|^{2}}{2}+\frac{\mathbb{E}\|\bm{\delta}(\mathbf{w}_{r})\|^{2}}{2}\\ &\overset{(\text{a})}{\leq}\frac{L^{2}}{2}\mathbb{E}\|\mathbf{\bar{w}}_{r,E}-\mathbf{w}_{r}\|^{2}+\frac{\mathbb{E}\|\bm{\delta}(\mathbf{w}_{r})\|^{2}}{2}\\ &=\frac{L^{2}}{2}\mathbb{E}\bigg{\|}\frac{1}{n}\sum_{k=1}^{n}(\mathbf{w}_{r,E}^{k}-\mathbf{w}_{r})\bigg{\|}^{2}+\frac{\mathbb{E}\|\bm{\delta}(\mathbf{w}_{r})\|^{2}}{2}\\ &=\frac{L^{2}}{2}\mathbb{E}\bigg{\|}\frac{1}{n}\sum_{k=1}^{n}(\mathbf{w}_{r,E}^{k}-\tilde{\mathbf{w}}_{r})+\tilde{\mathbf{w}}_{r}-\mathbf{w}_{r}\bigg{\|}^{2}+\frac{\mathbb{E}\|\bm{\delta}(\mathbf{w}_{r})\|^{2}}{2}\\ &\overset{(\text{b})}{=}\frac{L^{2}}{2}\mathbb{E}\bigg{\|}\frac{\eta}{n}\sum_{k=1}^{n}\sum_{e=0}^{E-1}\nabla f_{k}(\mathbf{w}_{e}^{k};\xi_{r,e}^{k})+\tilde{\mathbf{w}}_{r}-\mathbf{w}_{r}\bigg{\|}^{2}\\ &\ \ \ +\frac{\mathbb{E}\|\bm{\delta}(\mathbf{w}_{r})\|^{2}}{2}\end{split}$ (30)
-where (a) is due to the $L$-smoothness of $f(\cdot)$ and (b) is due to the local updating rules. Based on the Assumptions 2 and 3, we have
+Lemma 27 provides an upper bound on the expected value of the global loss function at $\bar{\mathbf{w}}_{r,E}$ in terms of the expected model distortion caused by downlink bit errors, i.e., $\mathbb{E}\|\bm{\delta}(\mathbf{w}_{r})\|^{2}$.
+By applying Lemmas 25 and 27 to Lemma 26, we obtain the following inequality:
+$\begin{split}&\mathbb{E}f(\mathbf{w}_{r+1})\leq\mathbb{E}f(\mathbf{w}_{r})+L\mathbb{E}\|\bm{\delta}(\mathbf{w}_{r})\|^{2}\\ &+\mathbb{E}\left\langle\nabla f(\mathbf{w}_{r})-\nabla f(\mathbf{\bar{w}}_{r,E}),\bm{\delta}(\mathbf{w}_{r})\right\rangle-\frac{\eta}{2}\sum_{e=0}^{E-1}\mathbb{E}\|\nabla f(\mathbf{\bar{w}}_{r,e})\|^{2}\\ &-\frac{\eta}{2K}\left[1-L\eta-2L^{2}\eta^{2}E(E-1)\right]\sum_{k=1}^{K}\sum_{e=0}^{E-1}\mathbb{E}\|\nabla f(\mathbf{w}_{r,e}^{k})\|^{2}\\ &+\frac{LE\eta^{2}(\sigma_{L}^{2}+\sigma_{G}^{2})}{2K}+\frac{L^{2}\eta^{3}\sigma_{L}^{2}(K+1)E(E-1)}{2K}.\end{split}$ (28)
+By leveraging $2\left\langle\mathbf{a},\mathbf{b}\right\rangle=\|\mathbf{a}\|^{2}+\|\mathbf{b}\|^{2}-\|\mathbf{a}-\mathbf{b}\|^{2}\leq\|\mathbf{a}\|^{2}+\|\mathbf{b}\|^{2}$ for any $\mathbf{a}$, $\mathbf{b}$, an upper bound for the third term on the RHS of (28) is derived as follows:
+$\begin{split}&\mathbb{E}\left\langle\nabla f(\mathbf{w}_{r})-\nabla f(\mathbf{\bar{w}}_{r,E}),\bm{\delta}(\mathbf{w}_{r})\right\rangle\\ &\leq\frac{\mathbb{E}\|\nabla f(\mathbf{w}_{r})-\nabla f(\mathbf{\bar{w}}_{r,E})\|^{2}}{2}+\frac{\mathbb{E}\|\bm{\delta}(\mathbf{w}_{r})\|^{2}}{2}\\ &\overset{(\text{a})}{\leq}\frac{L^{2}}{2}\mathbb{E}\|\mathbf{\bar{w}}_{r,E}-\mathbf{w}_{r}\|^{2}+\frac{\mathbb{E}\|\bm{\delta}(\mathbf{w}_{r})\|^{2}}{2}\\ &=\frac{L^{2}}{2}\mathbb{E}\bigg{\|}\frac{1}{K}\sum_{k=1}^{K}(\mathbf{w}_{r,E}^{k}-\mathbf{w}_{r})\bigg{\|}^{2}+\frac{\mathbb{E}\|\bm{\delta}(\mathbf{w}_{r})\|^{2}}{2}\\ &=\frac{L^{2}}{2}\mathbb{E}\bigg{\|}\frac{1}{K}\sum_{k=1}^{K}(\mathbf{w}_{r,E}^{k}-\tilde{\mathbf{w}}_{r})+\tilde{\mathbf{w}}_{r}-\mathbf{w}_{r}\bigg{\|}^{2}+\frac{\mathbb{E}\|\bm{\delta}(\mathbf{w}_{r})\|^{2}}{2}\\ &\overset{(\text{b})}{=}\frac{L^{2}}{2}\mathbb{E}\bigg{\|}\frac{\eta}{K}\sum_{k=1}^{K}\sum_{e=0}^{E-1}\nabla f_{k}(\mathbf{w}_{r,e}^{k};\xi_{r,e}^{k})+\tilde{\mathbf{w}}_{r}-\mathbf{w}_{r}\bigg{\|}^{2}\\ &\ \ \ +\frac{\mathbb{E}\|\bm{\delta}(\mathbf{w}_{r})\|^{2}}{2}\end{split}$ (29)
+where (a) is due to the $L$-smoothness property of $f(\cdot)$ and (b) applies the local updating rules in (3). Based on the Assumptions 2 and 3, we have
 $\begin{split}\mathbb{E}\|\nabla&f_{k}(\mathbf{w}_{r,e}^{k};\xi_{r,e}^{k})\|^{2}=\mathbb{E}\|\nabla f_{k}(\mathbf{w}_{r,e}^{k};\xi_{r,e}^{k})\\ &-\nabla f_{k}(\mathbf{w})+\nabla f_{k}(\mathbf{w})-\nabla f(\mathbf{w})+\nabla f(\mathbf{w})\|^{2}\\ &\leq\sigma_{L}^{2}+\sigma_{G}^{2}+\mathbb{E}\|\nabla f(\mathbf{w})\|^{2}.\end{split}$ (30)
 By combing (30) and (29), the bound in (29) can be further updated as
-$\begin{split}&\mathbb{E}\left\langle\nabla f(\mathbf{w}_{r})-\nabla f(\mathbf{\bar{w}}_{r,E}),\bm{\delta}(\mathbf{w}_{r})\right\rangle\\ &\leq\frac{L^{2}\eta^{2}}{2n^{2}}\sum_{k=1}^{n}\sum_{e=0}^{E-1}\mathbb{E}\|\nabla f_{k}(\mathbf{w}_{r,e}^{k};\xi_{r,e}^{k})\|^{2}+\frac{L^{2}}{2}\mathbb{E}\|\tilde{\mathbf{w}}_{r}-\mathbf{w}_{r}\|^{2}\\ &\ \ \ +\frac{\mathbb{E}\|\bm{\delta}(\mathbf{w}_{r})\|^{2}}{2}\\ &\overset{(\text{b})}{\leq}\frac{L^{2}\eta^{2}}{2n^{2}}\sum_{k=1}^{n}\sum_{e=0}^{E-1}\mathbb{E}\|\nabla f(\mathbf{w}_{r,e}^{k})\|^{2}+\frac{L^{2}\eta^{2}E(\sigma_{G}^{2}+\sigma_{L}^{2})}{2n}\\ &\ \ \ +\frac{L^{2}}{2}\mathbb{E}\|\tilde{\mathbf{w}}_{r}-\mathbf{w}_{r}\|^{2}+\frac{\mathbb{E}\|\bm{\delta}(\mathbf{w}_{r})\|^{2}}{2},\end{split}$ (32)
+$\begin{split}&\mathbb{E}\left\langle\nabla f(\mathbf{w}_{r})-\nabla f(\mathbf{\bar{w}}_{r,E}),\bm{\delta}(\mathbf{w}_{r})\right\rangle\\ &\leq\frac{L^{2}\eta^{2}}{2K^{2}}\sum_{k=1}^{K}\sum_{e=0}^{E-1}\mathbb{E}\|\nabla f_{k}(\mathbf{w}_{r,e}^{k};\xi_{r,e}^{k})\|^{2}+\frac{L^{2}}{2}\mathbb{E}\|\tilde{\mathbf{w}}_{r}-\mathbf{w}_{r}\|^{2}\\ &\ \ \ +\frac{\mathbb{E}\|\bm{\delta}(\mathbf{w}_{r})\|^{2}}{2}\\ &\leq\frac{L^{2}\eta^{2}}{2K^{2}}\sum_{k=1}^{K}\sum_{e=0}^{E-1}\mathbb{E}\|\nabla f(\mathbf{w}_{r,e}^{k})\|^{2}+\frac{L^{2}\eta^{2}E(\sigma_{G}^{2}+\sigma_{L}^{2})}{2K}\\ &\ \ \ +\frac{(L^{2}+1)\mathbb{E}\|\bm{\delta}(\mathbf{w}_{r})\|^{2}}{2}.\end{split}$ (31)
 Substituting (31) in (28), we have
-$\begin{split}&\mathbb{E}f(\mathbf{w}_{r+1})\leq\mathbb{E}f(\mathbf{w}_{r})+\frac{L(L+2)}{2}\mathbb{E}\|\mathbf{w}_{r}-\mathbf{\tilde{w}}_{r}\|^{2}+\frac{\mathbb{E}\|\bm{\delta}(\mathbf{w}_{r})\|^{2}}{2}\\ &-\frac{\eta}{2n}\left[1-\frac{n+L}{n}L\eta-2L^{2}\eta^{2}E(E-1)\right]\\ &\cdot\sum_{k=1}^{n}\sum_{e=0}^{E-1}\mathbb{E}\|\nabla f(\mathbf{w}_{r,e}^{k})\|^{2}-\frac{\eta}{2}\sum_{e=0}^{E-1}\mathbb{E}\|\nabla f(\mathbf{\bar{w}}_{r,e})\|^{2}\\ &+\frac{L(L+1)E\eta^{2}(\sigma_{L}^{2}+\sigma_{G}^{2})}{2n}+\frac{L^{2}\eta^{3}\sigma_{L}^{2}(n+1)E(E-1)}{2n}.\end{split}$ (33)
+$\begin{split}&\mathbb{E}f(\mathbf{w}_{r+1})\leq\mathbb{E}f(\mathbf{w}_{r})+\frac{(L+1)^{2}\mathbb{E}\|\bm{\delta}(\mathbf{w}_{r})\|^{2}}{2}\\ &-\frac{\eta}{2K}\left[1-\frac{K+L}{K}L\eta-2L^{2}\eta^{2}E(E-1)\right]\\ &\cdot\sum_{k=1}^{K}\sum_{e=0}^{E-1}\mathbb{E}\|\nabla f(\mathbf{w}_{r,e}^{k})\|^{2}-\frac{\eta}{2}\sum_{e=0}^{E-1}\mathbb{E}\|\nabla f(\mathbf{\bar{w}}_{r,e})\|^{2}\\ &+\frac{L(L+1)E\eta^{2}(\sigma_{L}^{2}+\sigma_{G}^{2})}{2K}+\frac{L^{2}\eta^{3}\sigma_{L}^{2}(K+1)E(E-1)}{2K}.\end{split}$ (32)
 For any $\eta$ that satisfies (8), we have
-$\begin{split}&\mathbb{E}f(\mathbf{w}_{r+1})\leq\mathbb{E}f(\mathbf{w}_{r})+\frac{L(L+2)}{2}\mathbb{E}\|\mathbf{w}_{r}-\mathbf{\tilde{w}}_{r}\|^{2}\\ &+\frac{\mathbb{E}\|\bm{\delta}(\mathbf{w}_{r})\|^{2}}{2}-\frac{\eta}{2}\sum_{e=0}^{E-1}\mathbb{E}\|\nabla f(\mathbf{\bar{w}}_{r,e})\|^{2}\\ &+\frac{L(L+1)E\eta^{2}(\sigma_{L}^{2}+\sigma_{G}^{2})}{2n}+\frac{L^{2}\eta^{3}\sigma_{L}^{2}(n+1)E(E-1)}{2n}.\end{split}$ (34)
-Summing (34) over $r=0,\cdots,R-1$, rearranging the terms, and multiplying $\frac{2}{\eta RE}$ on both sides, we obtain the following expression:
-$\begin{split}&\frac{1}{RE}\sum_{r=0}^{R-1}\sum_{e=0}^{E-1}\mathbb{E}\|\nabla f(\mathbf{\bar{w}}_{r,e})\|^{2}\leq\frac{2(\mathbb{E}f(\mathbf{w}_{0})-\mathbb{E}f(\mathbf{w}_{R}))}{\eta RE}\\ &+\frac{L(L+2)}{\eta RE}\sum_{r=0}^{R-1}\mathbb{E}\|\mathbf{\tilde{w}}_{r}-\mathbf{w}_{r}\|^{2}+\frac{1}{\eta RE}\sum_{r=0}^{R-1}\mathbb{E}\|\bm{\delta}(\mathbf{w}_{r})\|^{2}\\ &+\frac{L(L+1)\eta(\sigma_{L}^{2}+\sigma_{G}^{2})}{n}+\frac{L^{2}\eta^{2}\sigma_{L}^{2}(n+1)(E-1)}{n}.\end{split}$ (35)
-Based on (23) and (21) respectively, we have
-$\begin{split}\mathbb{E}\left[||\mathbf{\tilde{w}}_{r}-\mathbf{w}_{r}||^{2}\right]&=\mathbb{E}\left[\mathbb{E}\left[||\mathbf{\tilde{w}}_{r}-\mathbf{w}_{r}||^{2}|\mathbf{w}_{r}\right]\right]\\ &=\frac{d(4^{N}-1)}{3(2^{N}-1)^{2}}\cdot b_{r}\left(1-b_{r}\right)^{N-1}\cdot M_{r}^{2},\end{split}$ (36)
-$\begin{split}\mathbb{E}\|\bm{\delta}(\mathbf{w}_{r})\|^{2}&=\sum_{j=1}^{d}\mathbb{E}\left[|\delta(w_{r,d})|^{2}|{w}_{r,d}\right]\\ &\leq\sum_{j=1}^{d}\mathbb{E}\left[(b_{r}(1-b_{r})^{N-1}M)^{2}|{w}_{r,d}\right]\\ &\leq d(b_{r}(1-b_{r})^{N-1})^{2}M_{r}^{2}.\end{split}$ (37)
-By substituting (36) and (37) into (35), and denote $f^{*}$ as the minimum global training loss, we have
+$\begin{split}&\mathbb{E}f(\mathbf{w}_{r+1})\leq\mathbb{E}f(\mathbf{w}_{r})+\frac{(L+1)^{2}\mathbb{E}\|\bm{\delta}(\mathbf{w}_{r})\|^{2}}{2}\\ &-\frac{\eta}{2}\sum_{e=0}^{E-1}\mathbb{E}\|\nabla f(\mathbf{\bar{w}}_{r,e})\|^{2}\\ &+\frac{L(L+1)E\eta^{2}(\sigma_{L}^{2}+\sigma_{G}^{2})}{2K}+\frac{L^{2}\eta^{3}\sigma_{L}^{2}(K+1)E(E-1)}{2K}.\end{split}$ (33)
+Summing (33) over $r=0,\cdots,R-1$, rearranging the terms, and multiplying $\frac{2}{\eta RE}$ on both sides, we have
+$\begin{split}&\frac{1}{RE}\sum_{r=0}^{R-1}\sum_{e=0}^{E-1}\mathbb{E}\|\nabla f(\mathbf{\bar{w}}_{r,e})\|^{2}\leq\frac{2(\mathbb{E}f(\mathbf{w}_{0})-\mathbb{E}f(\mathbf{w}_{R}))}{\eta RE}\\ &+\frac{(L+1)^{2}}{\eta RE}\sum_{r=0}^{R-1}\mathbb{E}\|\bm{\delta}(\mathbf{w}_{r})\|^{2}\\ &+\frac{L(L+1)\eta(\sigma_{L}^{2}+\sigma_{G}^{2})}{K}+\frac{L^{2}\eta^{2}\sigma_{L}^{2}(K+1)(E-1)}{K}.\end{split}$ (34)
+By substituting (23) into (34), and denote $f^{*}$ as the minimum global training loss, we have
 $\frac{1}{RE}\sum_{r=0}^{R-1}\sum_{e=0}^{E-1}\mathbb{E}\|\nabla f(\mathbf{\bar{w}}_{r,e})\|^{2}\leq\frac{2(\mathbb{E}f(\mathbf{w}_{0})-f^{*})}{\eta RE}$
-$+\frac{L(L+2)d(4^{N}-1)}{3(2^{N}-1)^{2}\eta RE}\sum_{r=0}^{R-1}b_{r}(1-b_{r})^{N-1}M_{r}^{2}$
-$+\frac{d}{\eta RE}\sum_{r=0}^{R-1}(b_{r}(1-b_{r})^{N-1})^{2}M_{r}^{2}$
-$+\frac{L(L+1)\eta(\sigma_{L}^{2}+\sigma_{G}^{2})}{n}+\frac{L^{2}\eta^{2}\sigma_{L}^{2}(n+1)(E-1)}{n}.$ (38)
-Denoting $\hat{M}=max(M_{0},M_{1},\cdots,M_{R-1})$, we have
-$\frac{1}{RE}\sum_{r=0}^{R-1}\sum_{e=0}^{E-1}\mathbb{E}\|\nabla f(\mathbf{\bar{w}}_{r,e})\|^{2}\leq\frac{2(\mathbb{E}f(\mathbf{w}_{0})-f^{*})}{\eta RE}$
-$+\frac{L(L+2)d\hat{M}^{2}(4^{N}-1)}{3(2^{N}-1)^{2}\eta RE}\sum_{r=0}^{R-1}b_{r}(1-b_{r})^{N-1}$
-$+\frac{d\hat{M}^{2}}{\eta RE}\sum_{r=0}^{R-1}(b_{r}(1-b_{r})^{N-1})^{2}$
-$+\frac{L(L+1)\eta(\sigma_{L}^{2}+\sigma_{G}^{2})}{n}+\frac{L^{2}\eta^{2}\sigma_{L}^{2}(n+1)(E-1)}{n}.$ (39)
+$+\frac{(L+1)^{2}D(4^{N}-1)}{3(2^{N}-1)^{2}\eta RE}\sum_{r=0}^{R-1}b_{r}(1-b_{r})^{N-1}M_{r}^{2}$
+$+\frac{L(L+1)\eta(\sigma_{L}^{2}+\sigma_{G}^{2})}{K}+\frac{L^{2}\eta^{2}\sigma_{L}^{2}(K+1)(E-1)}{K}.$ (35)
+Finally, based on Assumption 6, we are able to show (9) in Lemma 2.